### v1

# Markov Chain Variance Estimation: A Stochastic Approximation Approach

Abstract
We consider the problem of estimating the asymptotic variance of a function defined on a Markov chain, an important step for statistical inference of the stationary mean. We design the first recursive estimator that requires $O(1)$ computation at each step, does not require storing any historical samples or any prior knowledge of run-length, and has optimal $O(\frac{1}{n})$ rate of convergence for the mean-squared error (MSE) with provable finite sample guarantees. Here, $n$ refers to the total number of samples generated. The previously best-known rate of convergence in MSE was $O(\frac{\log n}{n})$, achieved by jackknifed estimators, which also do not enjoy these other desirable properties. Our estimator is based on linear stochastic approximation of an equivalent formulation of the asymptotic variance in terms of the solution of the Poisson equation.
We generalize our estimator in several directions, including estimating the covariance matrix for vector-valued functions, estimating the stationary variance of a Markov chain, and approximately estimating the asymptotic variance in settings where the state space of the underlying Markov chain is large. We also show applications of our estimator in average reward reinforcement learning (RL), where we work with asymptotic variance as a risk measure to model safety-critical applications. We design a temporal-difference type algorithm tailored for policy evaluation in this context. We consider both the tabular and linear function approximation settings. Our work paves the way for developing actor-critic style algorithms for variance-constrained RL.
1 Introduction
Consider an irreducible and aperiodic discrete time Markov chain (DTMC) on a finite state space. Given a function $f$, consider the partial sum $S_{n}(f):=\sum\nolimits_{k=0}^{n-1}f(X_{k}).$ Under appropriate assumptions, it is well known that $\frac{S_{n}}{n}$ converges almost surely to $\bar{f}$, the stationary expectation of $f$. Moreover, Central Limit Theorem (CLT) holds (Douc et al., 2018, Chapter 21), and implies that for a positive constant $\kappa(f)$,
$\frac{S_{n}(f)-n\bar{f}}{\sqrt{n}}\Rightarrow N(0,\kappa(f))\quad\text{ as }\quad n\rightarrow\infty,$
where the above convergence is in distributions, and $N(0,\kappa(f))$ denotes the Gaussian distribution with mean 00 and variance $\kappa(f)$. Since $\kappa(f)$ corresponds to the variance in CLT, it is commonly referred to as asymptotic variance of the Markov chain.
The focus of the paper is to estimate $\kappa(f)$, which has been a long-standing area of research. Firstly, estimates of $\kappa(f)$ play an important role in the statistical analysis of simulation outputs obtained from a Markov chain Monte Carlo experiment, for example, in statistical inference for the stationary mean, $\bar{f}$ (see, Schmeiser (1982); Schruben (1983); Carlstein (1986); Kunsch (1989); Glynn and Whitt (1992); Chien et al. (1997); Flegal and Jones (2010); Wu (2009); Nordman (2009); Atchadé (2011); Dingeç et al. (2015)).
A second motivation comes from reinforcement learning (RL). In RL framework, the transition dynamics of the underlying MDP are unknown. The algorithm can obtain a sample of the MDP under any given policy, which specifies how actions are chosen in a given state. The traditional goal in an average reward RL problem is to find a policy that maximizes $\frac{S_{n}}{n}$. While the need to optimize over the average reward is well motivated, for applications in safety-critical domains, for example, healthcare or finance, it is also crucial to control adverse outcomes. As an example, one can consider the portfolio optimization problem, where the objective is to find an investment strategy that maximizes the return (expected value), while keeping the variability under control. As we will see later in Section 5.2, a natural measure of variability in this setup is the asymptotic variance $\kappa(f)$. Estimating $\kappa(f)$ is a vital sub-problem in mean-variance policy optimization, for instance, as a critic in an actor-critic framework, cf. Prashanth and Ghavamzadeh (2016).
Estimating $\kappa(f)$, while practically relevant, is particularly challenging compared to estimating variance using independent and identically-distributed (i.i.d.) samples because of temporal correlations introduced due to the underlying Markov chain structure. The problem of estimating $\kappa(f)$ has been extensively studied in literature. However, to the best of our knowledge, all the existing estimators have limitations of being either non-recursive in that the estimator at step $n$ cannot be expressed as a function of the various statistics computed at step $n-1$, necessitating re-computing the estimator from scratch at each step, or requiring to store a large number of historical observations, or having a sub-optimal rate of convergence of mean-squared estimation error to 00, or some combination of these. We discuss the related literature in details in Section 1.2. In this work, our goal is to design an estimator for $\kappa(\cdot)$ that utilizes samples from the given Markov chain along a single trajectory, is recursive and memory-efficient, easy to compute, and has an optimal rate of convergence in mean-squared sense.
1.1 Contributions
We now briefly describe the main contributions of this work.
  1. 1.
    We propose the first recursive estimator (or algorithm) for estimating $\kappa(f)$ that requires $O(1)$ computation at each step, does not require storing any historical samples, and enjoys optimal $O(\frac{1}{n})$ rate of convergence for the mean-squared error (MSE). Here, $n$ refers to the total number of samples generated. The best previously known rate-of-convergence in MSE for any estimator (irrespective of their memory and computation requirements or being recursive) was $O(\frac{\log n}{n})$ (cf., jackknifed estimators in Dingeç et al. (2015)).
    The proposed estimator does not require any prior knowledge of the run length. In addition, we establish finite sample bounds on MSE of the proposed estimator that hold for any run length. In an earlier version of this paper Agrawal et al. (2024), we proposed an algorithm that required generating samples along two independent trajectories of the underlying Markov chain. But this may not always be feasible. In this version, we develop a novel single-trajectory algorithm.
  2. 2.
    Our algorithm can be specialized to estimate the stationary variance of a Markov chain. It has similar convergence guarantees, as well as all the other desirable properties, even in this setting. Additionally, it can be used to design a recursive estimator for the variance in i.i.d. setting, which requires $O(1)$ memory and computation at each step, and has an optimal $O(\frac{1}{n})$ rate of convergence for MSE.
  3. 3.
    In contrast to prior work on DTMC variance estimation, a key idea behind our approach is to use stochastic approximation (SA) for statistical estimation problems to develop estimators with provably-good finite sample guarantees. In particular, we work with an expression for the DTMC asymptotic variance in terms of the solution of the so-called Poisson equation (these solutions are called value functions). We use linear SA on this expression to develop an algorithm for variance estimation. Our algorithm uses Temporal Difference (TD) learning from prior work for estimating the value function Zhang et al. (2021). We also obtain finite sample bounds for TD learning, improving the MSE bound of Zhang et al. (2021) from $O(\frac{\log n}{n})$ to $O(\frac{1}{n})$, which is of independent interest.
    The mean square guarantees are obtained by first establishing that the underlying matrix of our linear SA is Hurwitz, and then builds upon recent work Zhang et al. (2021) on finite sample mean square error bounds for linear SA. A novel step in our proof is the use of Poisson equation (again) to handle Markov noise, which enables us to shave off the multiplicative $\log n$ factor in the error.
  4. 4.
    We generalize our estimator in several directions.
    1. (a)
      First, we develop extensions of it to estimate the covariance matrix of a vector-valued function $f$.
    2. (b)
      Second, we extend our estimator to the settings where the underlying Markov chain has a large state space. Here, we consider linear approximations of the value function using a small number of feature vectors. We then use this approximate value function to estimate the asymptotic variance.
    3. (c)
      Finally, we consider the average reward RL framework in the risk-sensitive setting. We propose asymptotic variance as a natural measure of risk in contrast to the stationary variance studied in prior work. We then focus on evaluating the asymptotic variance of a given policy. We adapt our variance estimator to this setting, and obtain a novel TD like algorithm and establish its finite sample error bounds.
1.2 Related Literature
We now discuss the existing estimators for $\kappa(f)$ and related literature in details.
Table 1: Summary of the results on estimators for $\kappa(f)$
Estimator Representative references (non exhaustive) Mean squared error Recursive[a] Memory requirement[b]
Batched Means (BM) Schmeiser (1982); Chien et al. (1997) $O\left(\frac{1}{n^{2/3}}\right)$ $O(n)$
Overlapping BM (OBM) Flegal and Jones (2010)
Meketon and Schmeiser (1984)
Block bootstrap Lahiri (1999); Nordman (2009) $O\left(\frac{1}{n^{2/3}}\right)$ $O(n)$
Standardized Goldsman et al. (1990) $O\left(\frac{1}{n^{4/5}}\right)$ $O(n)$
time series (STS) Aktaran-Kalayci et al. (2011)
Recursive BM Wu (2009); Yau and Chan (2016) $O\left(\frac{1}{n^{2/3}}\right)$ $O(1)$
Jackknifed-(O)BM/STS Dingeç et al. (2015) $O\left(\frac{\log n}{n}\right)$ $O(n)$
Linear SA This work $O\left(\frac{1}{n}\right)$ $O(1)$
  • [a] The estimator at step $n$ can be expressed as a function of the various estimates at step $n-1$, and can be computed in a constant time (independent of $n$). A non recursive estimator requires re-computing the estimator from scratch at each step.
    [b] The memory required for computing the estimator at step $n$.
Batch Means (BM)
A popular estimator for $\kappa(f)$ is the BM estimator Conway et al. (1959); Schmeiser (1982), which we informally describe next. Consider $n$ observations of the function $f$ evaluated on the $n$ consecutive states of the underlying Markov chain, and divide these observations into $b$ non-overlapping batches, each of size $m$ (for simplicity, assume that $n=mb$). For $i\in[b]$, let $m_{i}$ denote the sample mean of the observations from batch $i$. Then, the BM estimator corresponds to the sample variance of $m_{i}$ scaled by the batch size $m$. BM and its variants have been extensively studied (see, Schmeiser (1982); Meketon and Schmeiser (1984); Carlstein (1986); Welch (1987); Kunsch (1989); Alexopoulos et al. (2007); Meketon and Schmeiser (2007); Alexopoulos et al. (2011)). See Politis et al. (1999) for an introduction to the different estimators and their relations. Roughly speaking, here, the batch size $m$ governs the bias, while the number of batches $b$ governs the variance of the estimator.
Glynn and Whitt (1991) establish limitations of BM in the settings where the number of batches $b$ does not increase with the total number of samples $n$. Glynn and Whitt (1992) establish conditions that ensure the asymptotic validity of BM. Chien et al. (1997) show that BM with both $b$ and $m$ increasing with $n$, is asymptotically unbiased (as $n$ increases), and converges to $\kappa(f)$ in mean square. In addition, they also show an asymptotic equality for MSE of BM, establishing it to be $O(\frac{1}{n^{2/3}})$.
Overlapping BM
While BM may look wasteful, overlapping BM Meketon and Schmeiser (1984, 2007) uses $b$ overlapping batches, each of size $m$, where the consecutive batches use just $1$ additional observation, along with $m-1$ previous (consecutive) ones. Here, total number of samples satisfy $n=m+b-1$. When $n$ is known ahead of time, Flegal and Jones (2010) establish $O(\frac{1}{n^{2/3}})$ MSE for overlapping BM estimator. Notably, overlapping BM has the same MSE as BM (order wise) since the bias of the two estimators is the same, and the variance of the overlapping BM is $\frac{2}{3}$ times that of BM (improvement only by a constant factor). Thus, both BM and overlapping BM have a sub-optimal rate of convergence for MSE. See Alexopoulos et al. (2011) for a survey of results on BM and overlapping BM estimator.
In addition to being sub-optimal, another drawback of the classical and overlapping BM is that either their error guarantees are only satisfied at time $n$ and are not anytime, or they are not recursive (sequential), i.e., one needs to re-compute the estimator at each time using computation that is linear in the number of observations. This is because the batch sizes are predetermined, and depend on $n$. Moreover, one requires storing all the previous observations in order to re-compute the estimator at each step, hence, $O(n)$ computation at time $n$. Wu (2009) develop a recursive BM estimator ($O(1)$ computation at each time) with increasing batch sizes that are independent of $n$. Their estimator does not require storing the past observations, improving the memory requirement to $O(1)$, and is asymptotically unbiased. However, Wu (2009) only give asymptotic guarantees, and establish a sub-optimal rate of $O(\frac{1}{n^{2/3}})$ for MSE. Yau and Chan (2016) also propose a different recursive BM estimator with similar guarantees.
Other estimators
Several bootstrap-based estimators have also been proposed in literature Kunsch (1989); Liu et al. (1992); Carlstein (1986); Romano (1992); Politis and Romano (1994). However, these are computationally intensive, and hence, we do not discuss them in details in this work. Nordman (2009); Lahiri (1999) study the asymptotic rates for bias and variance of the different bootstrap-based estimators; results summarized in Table 1. We refer the reader to Lahiri (2013) for an introduction to this approach. Schruben (1983) develop standardized time series based estimators, extensively studied in Goldsman and Schruben (1984); Glynn and Iglehart (1988); Goldsman et al. (1990); Goldsman and Schruben (1990). Trevezas and Limnios (2009) propose maximum likelihood estimator for $\kappa(f)$, and prove that it is strongly consistent and asymptotically normal.
Jackknifed versions of estimators
Since a significant contribution to MSE of BM and its variants is due to bias of the estimator, Dingeç et al. (2015) propose using jackknife Tukey (1958) to reduce it at a cost of modest increase in variance of the estimator. They develop jackknifed versions of BM and overlapping BM estimators with a significantly reduced MSE of $O(\frac{\log n}{n})$. However, these are not recursive, and require $O(n)$ computation and $O(n)$ memory at time step $n$, rendering them less practical.
Poisson Equations
Finally, Douc et al. (2022) resort to an alternative formulation for $\kappa(f)$ involving solution to a certain Poisson equation, propose a coupled Markov chains-based estimator for $\kappa(f)$, and provide its asymptotic convergence guarantees. The work that is closest to ours is Benton (2022) which proposed a SA-based estimator for $\kappa(f)$, and established its asymptotic convergence. While their estimator is recursive, and requires $O(1)$ memory and computation at each step, they do not provide either the rates of convergence, nor finite-sample bounds on MSE of the estimator. In the current work, we propose a different SA-based estimator, along with finite-sample guarantees on its MSE.
Note that the focus of this work is to estimate the asymptotic variance $\kappa(f)$. To this end, as we will see later, we design a TD-like linear SA algorithm. This problem of estimating $\kappa(f)$ is very different from studying the variance of iterates of any particular algorithm like TD, or designing algorithms with minimum variance of the iterates themselves. These latter objectives too are well-studied in literature; cf. Devraj and Meyn (2017); Yin and Wang (2020); Chen et al. (2020a, b); Hao et al. (2021); Zhu et al. (2023). In contrast to these orthogonal prior works, which focus on studying the variance of TD learning and its variants, we do not study the variance of a linear SA algorithm itself. We use linear SA simply as a tool for solving the variance estimation problem. To further elaborate on this distinction, for concreteness, consider the setting of Markov Decision Processes (MDPs), where the transition probabilities are known. Here, for evaluating a given policy, one can use value iteration (VI). Since the MDP is known, there is no variance associated with the iterates of VI. However, the asymptotic variance that we consider in this work is still well-defined, and exists due to inherent randomness associated with the probabilistic transitions of the MDP. This work focuses on estimating this asymptotic variance.
Later, in Section 5, we present an application of the proposed estimator in the average reward RL setting. We postpone the related literature from RL to Section 5.3.
1.3 Setup and Asymptotic Variance
In this subsection, we formally introduce the problem. To this end, we first define some notation. Let $\mathbb{N}$ denote the set of Natural numbers, $\mathbb{R}$ the set of Real numbers, and let $d\geq 1$. For a column vector $v\in\mathbb{R}^{d}$, let $v^{T}$ denote its transpose and let $\|v\|_{2}$ denote the $\ell_{2}$-norm of $v$. Let ${\bf 1}$ (or ${\bf e}$) and ${\bf 0}$ denote the all ones and all zeros vectors in appropriate dimensions, respectively.
We now describe the underlying dynamics, followed by formally defining the asymptotic variance. For $S\in\mathbb{N}$, consider an irreducible and aperiodic DTMC, ${\mathcal{M}}$, on a finite state space ${\mathcal{S}}=\left\{1,\dots,S\right\}$ with transition matrix $P$ and a unique stationary distribution, denoted by $\pi$ (Levin and Peres, 2017, Chapter 1). For $n\in\mathbb{N}$, let $X_{n}\in{\mathcal{S}}$ denote the state at time $n$. Consider $f:{\mathcal{S}}\rightarrow\mathbb{R}$, a function defined on state space of ${\mathcal{M}}$, and let $\bar{f}$ denote its stationary expectation, i.e., $\bar{f}:=\mathbb{E}_{\pi\!}\!\left[f(X)\right]$. Since $f(X)-\bar{f}$ has stationary mean zero, there exists $V:{\mathcal{S}}\rightarrow\mathbb{R}$ that solves the following Poisson equation (Douc et al., 2018, Section 21.2):
$f-\bar{f}{\bf 1}=V-PV.$
Since $V$ solves (1.3), observe that $V+c{\bf 1}$ for any constant $c\in\mathbb{R}$ is also a solution. Let $V^{*}$ denote the solution normalized so that it is orthogonal to ${\bf 1}$ in $\ell_{2}$ norm, i.e., ${\bf 1}^{T}V^{*}=0$. It is well-known that the set of solutions of (1.3) takes the form $S_{f}:=\left\{V^{*}+c{\bf 1}|c\in\mathbb{R}\right\}$ (Douc et al., 2018, Section 21.2).
For a Markov chain starting in state $x_{0}$, the asymptotic variance is given by
$\lim\limits_{n\rightarrow\infty}~{}\operatorname{Var}\left[\frac{1}{\sqrt{n}}\sum\limits_{k=0}^{n-1}f(X_{k})\bigg{|}X_{0}=x_{0}\right],$ (2)
and is independent of the starting state (Proposition 1.1 below). This follows from geometric mixing of $\mathcal{M}$. We denote this constant by $\kappa(f)$ (or $\kappa$, since $f$ is fixed). In this paper, we assume that $|f|$ is bounded. Henceforth, without loss of generality, this bound is assumed to be $1$. Below, we first present equivalent formulations for $\kappa(f)$ that will be useful in designing efficient algorithms for estimating it.
Proposition 1.1.
Given a irreducible and aperiodic DTMC, asymptotic variance defined in (2) is a constant independent of the starting state, and is given by
$\kappa(f)=\lim\limits_{n\rightarrow\infty}~{}\operatorname{Var}\left[\frac{1}{\sqrt{n}}\sum\limits_{k=0}^{n-1}f(X_{k})\bigg{|}X_{0}\sim\pi\right].$
Furthermore, given any solution $V\in S_{f}$ to (1.3),
$\kappa(f)$ $=~{}~{}\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})^{2}\right]+2\lim\limits_{n\rightarrow\infty}\sum\limits_{j=1}^{n-1}\mathbb{E}_{\pi\!}\!\left[\left(f(X_{0})-\bar{f}\right)\left(f(X_{j})-\bar{f}\right)\right]$ (3)
$=~{}~{}2\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})V(X)\right]-\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})^{2}\right]$ (4)
$=~{}~{}\mathbb{E}_{\pi\!}\!\left[V^{2}(X)\right]-\mathbb{E}_{\pi\!}\!\left[(PV)^{2}(X)\right].$ (5)
It is worth noting that the first term in (3) accounts for the per-step stationary variance, while the second term encompasses the temporal correlation introduced due to the Markov chain structure. These different representations for asymptotic variance are well known in literature. For example, see (Douc et al., 2018, Theorem 21.2.6) for a similar formulation. For completeness, we give a proof of the Proposition in Appendix B.1.
Continuing, the second representation in equation (4) formulates the variance of $f(\cdot)$ in terms of the corresponding solution of the Poisson equation (1.3). Note that this representation isn’t affected by choice of $V$ (any constant shift in $V$ doesn’t affect $\kappa$). We will use a modification of the form in (4) to design a linear SA algorithm to estimate $\kappa$ for a given function $f$.
Finally, the third equality follows by using the Poisson equation to replace $(f(X)-\bar{f})$ terms in (4) by $V(X)-PV(X)$.
1.4 Stochastic Approximation: An Overview
Before proceeding to design an SA-based estimator for $\kappa(f)$, in this subsection, we give a brief overview of the SA updates we will repeatedly use throughout the rest of the paper. The earliest SA algorithms, which are iterative updates that can be used to solve for roots of certain functions, date back at least to Robbins and Monro (1951). While the asymptotic convergence of SA algorithms has long been known Borkar (2008), finite sample guarantees for linear SA algorithm have only recently been established Srikant and Ying (2019); Zhang et al. (2021). In this work, we will use SA for finding averages as well as roots of linear functions, which we discuss next.
Mean Estimation with Step Sizes
An SA algorithm to estimate $\Gamma:=\mathbb{E}_{\pi\!}\!\left[g(X)\right]$ for a function $g$ using samples $g(X_{0}),\dots,g(X_{n})$, corresponds to the following recursive update with an appropriately chosen step size $\alpha_{n}>0$, and an initial iterate $\Gamma_{0}$:
$\Gamma_{n+1}=(1-\alpha_{n})\Gamma_{n}+\alpha_{n}g(X_{n}).$
Observe that for $\Gamma_{0}=0$ and $\alpha_{n}=\frac{1}{n+1}$, this reduces to the classical empirical average estimator for mean. Moreover, if $X_{0},\dots,X_{n}$ are i.i.d. according to $\pi$, then empirical mean is known to converge almost surely to $\Gamma$ (from Law of Large Numbers) under mild regularity conditions. Similar convergence results hold for averages in (1.4) under certain regularity conditions on $\alpha_{n}$ and data-generating distributions.
Linear SA
[…]
As we will see in the following sections, these update methods play a crucial role in arriving at a simple recursive estimator that satisfies all the desirable properties listed earlier in Table 1.
2 Variance Estimation: Tabular Setting
With the expressions for asymptotic variance derived in Proposition 1.1 above, in this section, we design an algorithm for estimating it. A natural first approach is to use the Monte Carlo technique for estimating $\kappa(f)$, for example, the Batch Means estimator and its variants, which proceed by appropriately truncating the infinite summation in (1.1). As discussed earlier in Section 1.2, these approaches suffer from either being non-recursive, computationally demanding, having sub-optimal rates of convergence, or a combination of these. We refer the reader to Section 1.2 and Appendix B.3 for a discussion.
2.1 The Algorithm
Since the equivalent formulation for $\kappa$ in (3) involves an infinite summation, algorithms approximating it using a finite summation could possibly suffer from being non-adaptive as the vanilla Monte Carlo approaches. Instead, we design a fully-adaptive TD-type algorithm for estimating $\kappa$, using SA for a modification of the expression in (4) (Equation (8) below). This alternative representation helps us to design a linear SA algorithm with finite-sample guarantees on the mean (squared) estimation error. For a solution $V$ of (1.3), let $\bar{V}=\mathbb{E}_{\pi\!}\!\left[V(X)\right]$. Then,
$\kappa(f)=\mathbb{E}_{\pi\!}\!\left[2f(X)V(X)-2f(X)\bar{V}-f^{2}(X)+f(X)\bar{f}\right].$ (8)
This immediately follows from rearranging terms in (4).
Input: Time horizon $n>0$, constants $c_{1}>0$, $c_{2}>0$, $c_{3}>0$ and step-size sequence $\left\{\alpha_{k}\right\}$.
Initialization: $\bar{f}_{0}=0$, $V_{0}=\vec{\bf 0}$, $\bar{V}_{0}=0$, and $\kappa_{0}=0$.
for $k\leftarrow 1$ to $n$ do
      Observe ($X_{k},f(X_{k}),X_{k+1}$).
      // Average reward $\bar{f}$ estimation
$\bar{f}_{k+1}=\bar{f}_{k}+c_{1}\alpha_{k}(f(X_{k})-\bar{f}_{k}).$ (9)
      // $V$ estimation
$\delta_{k}:=f(X_{k})-\bar{f}_{k}+V_{k}(X_{k+1})-V_{k}(X_{k}),$ (10)
$V_{k+1}(x)=\begin{cases}V_{k}(x)+\alpha_{k}\delta_{k}\left(1-{1}/{S}\right),&\text{ for }x=X_{k},\\ V_{k}(x)-{\alpha_{k}\delta_{k}}/{S},&\text{ otherwise}.\end{cases}$ (11)
       // $\bar{V}$ estimation
$\bar{V}_{k+1}:=\bar{V}_{k}+c_{2}\alpha_{k}(V_{k}(X_{k})-\bar{V}_{k}),$ (12)
      // Asymptotic Variance $\kappa(f)$ estimation
$\kappa_{k+1}=(1-c_{3}\alpha_{k})\kappa_{k}+c_{3}\alpha_{k}\left(2f(X_{k})V_{k}(X_{k})-2f(X_{k})\bar{V}_{k}-f^{2}(X_{k})+f(X_{k})\bar{f}_{k}\right).$ (13)
end for
Output : Estimates $\kappa_{n}$, $\bar{V}_{n}$, $V_{n}$, and $\bar{f}_{n}$
Algorithm 1 Estimating $\kappa(f)$: Tabular Setting
Estimating $\kappa(f)$
For $n\geq 1$, and time steps $k=1,2,\dots,n$, let $X_{k}$ represent the state of the underlying Markov chain. To estimate $\kappa(f)$, our algorithm uses SA for the expression in (8), which corresponds to the mean estimation with step sizes, discussed in (1.4). To get observations of the function being averaged, it first estimates $\bar{f}$, $V$, and the corresponding $\bar{V}$. Denote these estimates at time $k$ by $\bar{f}_{k}$, $V_{k}$, and $\bar{V}_{k}$. Equation (13) in Algorithm 1 corresponds to SA update for estimating $\kappa(f)$ using these estimates in (8).
Estimating $\bar{f}$
This corresponds to estimating $\mathbb{E}_{\pi\!}\!\left[f(X)\right]$ using the latest estimate $\bar{f}_{k}$, as well as the new observed value $f(X_{k})$. We again use (1.4) for this step, which corresponds to update (9) in Algorithm 1.
Estimating $V$
Recall that $V$ is a solution for (1.3), which is a linear equation in $V$. Hence, we will use the SA algorithm in (1.4) to estimate it. Term $\delta_{k}$ in (10) in Algorithm 1 represents the SA adjustment term for estimating $V$ using the estimates $\bar{f}_{k}$, $V_{k}$ computed at the previous step and the observed function value $f(X_{k})$. Notice that since (1.3) may not have a unique solution, we particularly consider the one that is orthogonal to the all ones vector ${\bf 1}\in\mathbb{R}^{S}$ in $\ell_{2}$-norm. Thus, in each iteration of Algorithm 1, we project $\delta_{k}$ to the subspace orthogonal to ${\bf 1}$. This corresponds to the adjustments made in the $V$ update in (11), where $S$ denotes the size of state space $\left|{{\mathcal{S}}}\right|$ of the Markov chain ${\mathcal{M}}$.
Estimating $\bar{V}$
This corresponds to the mean estimation update in (12) using the estimates $\bar{V}_{k}$ and $V_{k}$ computed at the previous step (using (1.4)).
Remark 2.1 (Justification for projection in Algorithm 1).
Observe that $\kappa(f)$ is independent of the choice of $V$ (choice of solution of (1.3) does not affect $\kappa$), since $\bar{V}$ in (8) corresponds to the stationary expectation of the particular solution $V$ used in that formulation. Moreover, estimating a solution $V$ of (1.3) does not require the projection step, since the SA for (1.3) converges to some solution; cf. (Zhang et al., 2021, Algorithm 1). To further elaborate, if in each iteration $k$ of Algorithm 1, $V_{k}$ corresponds to an estimate for a different solution of (1.3), then as $k\rightarrow\infty$, $\{V_{k}\}$ can be still be shown to converge to some solution of (1.3), say $V^{\dagger}$. However, to estimate $\kappa(f)$ using (8), we further need to estimate the stationary mean of $V^{\dagger}$, say $\bar{V}^{\dagger}$. However, since without projection, the estimates $V_{k}$ in Algorithm 1 may correspond to different solutions of (1.3), and do not necessarily all estimate $V^{\dagger}$, the iterates $\bar{V}_{k}$ are not guaranteed to converge to $\bar{V}^{\dagger}$. Hence, we introduce the projection step to ensure that for each $k$, $V_{k}$ correspond to the same estimate for a solution of (1.3) (which is the one orthogonal to $\bf 1$ in $\ell_{2}$ norm), and hence, $\bar{V}_{k}$ converge to its stationary expectation.
Remark 2.2.
Notice that all the updates in Algorithm 1 in (9), (11), (12), and (13) at time $k$ use estimates from the previous time step $k-1$. Hence, the order in which these updates are performed is irrelevant.
2.2 Convergence Rates
We now bound the estimation error of the proposed algorithm. For each $k\in\mathbb{N}$, let $Y_{k}:=(X_{k},X_{k+1}).$ Observe that $\mathcal{M}_{2}:=\left\{Y_{k}\right\}_{k\geq 1}$ is a Markov chain. Let its state space be denoted by ${\mathcal{Y}}$, which is a finite set. As for $\mathcal{M}$, $\mathcal{M}_{2}$ mixes geometrically fast (Bhatnagar and Lakshmanan, 2016, Proposition 1) and has a unique stationary distribution.
Next, let $\mathcal{F}\in\mathbb{R}^{S}$ be the vector of function values, i.e., for $i\in{\mathcal{S}}$, the $i^{th}$ coordinate of $\mathcal{F}$ equals $f(i)$, and $f_{\max}:=\|\mathcal{F}\|_{\infty}$. Recall, without loss of generality, we assume that $f\in[-1,1]$, and hence, $f_{\max}\leq 1$. Moreover, $V^{*}$ was defined after (1.3), and let $D_{\pi}$ denote the diagonal matrix of the stationary distribution $\pi$ of ${\mathcal{M}}$, i.e., $D_{\pi}(i,i)=\pi(i)$, for all $i\in{\mathcal{S}}$. Define
$\Delta_{1}:=\min\!\left\{\!{\bf v}^{T}\!D_{\pi}(I\!-\!P){\bf v}~{}|~{}{\bf v}\!\in\!\mathbb{R}^{S},\|{\bf v}\|_{2}=1,{\bf v}^{T}\!{\bf 1}=0\right\}.$
Clearly, feasible vectors $\bf v$ in $\Delta_{1}$ are non-constant vectors. Then, from (Tsitsiklis and Van Roy, 1997, Lemma 7), we have that ${\bf v}^{T}D_{\pi}(I-P){\bf v}>0.$ Since the feasible region in $\Delta_{1}$ is non-empty and compact, we get $\Delta_{1}>0$.
Next, for the step-size constants $c_{1}$, $c_{2}$, and $c_{3}$ (inputs to Algorithm 1), define
$\eta:=\left(c^{2}_{1}+5+2c^{2}_{2}+10c^{2}_{3}\right)^{\frac{1}{2}},$ (14)
and let $\Theta_{\pi}:=[\bar{f}~{}V^{*}~{}\bar{V}^{*}~{}\kappa(f)]$. The following theorem bounds the estimation error of Algorithm 1, which immediately implies ${O}(\frac{1}{n})$ convergence rate for the mean-squared estimation error of $\kappa(f)$ to 00.
Theorem 2.1.
Consider Algorithm 1 with $c_{1}$, $c_{2}$, and $c_{3}$ satisfying:
$c_{1}\geq\frac{1}{2\Delta_{1}}+\frac{\Delta_{1}}{2},\qquad\frac{5}{249}(5-2\sqrt{2})\Delta_{1}\leq c_{3}\leq\frac{5}{249}(5+2\sqrt{2})\Delta_{1},$
and
$c_{3}-\frac{498c^{2}_{3}}{7\Delta_{1}}+\frac{7\Delta_{1}}{498}\leq c_{2}\leq-3c_{3}+\frac{5}{83}\sqrt{498c_{3}\Delta_{1}-17\Delta^{2}_{1}}.$
Recall that $\eta$ and $\Theta_{\pi}$ were defined around (14). For a constant $B>1$, let $\xi_{1}:=3\left(1+\|\Theta_{\pi}\|_{2}\right)^{2}$, $\xi_{2}:=112B(1+\|\Theta_{\pi}\|_{2})^{2}$, and let $\Theta_{n}:=[\bar{f}_{n}~{}V^{T}_{n}~{}\bar{V}_{n}~{}\kappa_{n}]$ denote the vector of estimates at step $n$.
  1. (a)
    Let $\alpha_{i}=\alpha$ for all $i$, such that
    $\alpha<\min\left\{\frac{40}{\Delta_{1}},\frac{1}{28B\left(1+\frac{20\eta^{2}}{\Delta_{1}}\right)}\right\}.$
    Then, for all $n\geq 1$,
    $\mathbb{E}\!\left[\|\Theta_{n}-\Theta_{\pi}\|^{2}_{2}\right]\leq\xi_{1}\left(1-\frac{\Delta_{1}\alpha}{40}\right)^{n}+\frac{20\xi_{2}\alpha\eta^{2}}{\Delta_{1}}+\xi_{2}\alpha.$
  2. (b)
    Let $\alpha_{i}=\frac{\alpha}{i+h}$ for all $i\geq 0$, with $\alpha$ and $h$ chosen so that
    $h\geq\max\left\{1+\frac{\alpha\Delta_{1}}{40},1+28B\left(\frac{20\eta^{2}\alpha}{\Delta_{1}}+\alpha+\frac{20}{\Delta_{1}}\right)\right\},\quad\text{and}\quad\alpha>\frac{40}{\Delta_{1}}.$
    Then, for all $n\geq 0$,
    $\mathbb{E}\!\left[\|\Theta_{n}-\Theta_{\pi}\|^{2}_{2}\right]$ $\leq\xi_{1}\left(\frac{h}{n+h}\right)^{\frac{\alpha\Delta_{1}}{40}}+\frac{5\xi_{2}e^{2}\eta^{2}(20+\Delta_{1})\alpha^{2}}{(n+h)(\alpha\Delta_{1}-40)}+\frac{\xi_{2}\alpha}{n+h}.$
The above theorem follows from a more general result (Theorem 3.1) presented in the next section. We refer the reader to Appendix C.1 for a formal reduction from the setting of Theorem 3.1 to that in the above theorem. We now present some remarks about the bound in Theorem 2.1 above.
First, observe that since $\mathbb{E}\!\left[(\kappa_{n}-\kappa(f))^{2}\right]\leq\mathbb{E}\!\left[\|\Theta_{n}-\Theta_{\pi}\|^{2}_{2}\right]$ and $\mathbb{E}\!\left[\|V_{n}-V^{*}\|^{2}_{2}\right]\leq\mathbb{E}\!\left[\|\Theta_{n}-\Theta_{\pi}\|^{2}_{2}\right]$, the bound in Theorem 2.1 also holds for MSE in estimation of $\kappa(f)$ as well as $V^{*}$.
Theorem 2.1$(a)$ bounds the mean-squared estimation error of the proposed algorithm in case of constant step size. Though the first term in this bound decays exponentially fast as the number of iterations $n$ increase, the other two terms are constant and become a bottleneck after sufficiently large $n$. This is a well-known behavior of SA algorithms with constant step sizes. In fact, this algorithm suffers similar drawback of being non-recursive as the vanilla Monte Carlo BM estimator (see, Table 1). In order to ensure the mean-squared error smaller than $\epsilon^{2}$ for $\epsilon>0$ (or mean absolute error smaller than $\epsilon$), one needs to pick the step size $\alpha$ as a function of $\epsilon$ (also see Corollary 2.1.1 below). This drawback of being non-recursive is overcome by choosing the step size to be diminishing (as in Theorem 2.1$(b)$).
Second, it has been shown recently that linear SA with constant step size, in presence of Markov noise, suffers from an asymptotic bias (as $n\rightarrow\infty$); see Huo et al. (2023); Nagaraj et al. (2020). This is in contrast to the i.i.d. noise setting, where the asymptotic bias is shown to be 00 in Mou et al. (2020); Lakshminarayanan and Szepesvari (2018). This motivates us to study SA with diminishing step-size, considered in Theorem 2.1$(b)$.
Remark 2.3.
Recall that previously best-known rate of convergence for MSE for estimating $\kappa(f)$ for a Markov chain was $O(\frac{\log n}{n})$ (see Table 1). For diminishing step size, Algorithm 1 achieves ${O}(\frac{1}{n})$ rate of convergence (Theorem 2.1$(b)$). As in part $(a)$, the first term in the bound in part $(b)$ decays faster, and the other two terms are rate-determining $O(\frac{1}{n})$ terms. We believe that the rate of convergence of $\Omega(\frac{1}{n})$ for MSE is tight. In the special setting of i.i.d. noise with parametric distributions, this follows from the Cramér-Rao lower bound Nielsen (2013). We discuss a simple estimator that achieves this rate in i.i.d. setting in Appendix C.4, and also propose a SA-based recursive estimator achieving the same rate in i.i.d. setting in the following section (Section 2.3).
Using bounds in Theorem 2.1, the corollary below presents the number of iterations of Algorithm 1 needed to have the mean estimation error bounded by $\epsilon$. It follows from setting the mean-squared estimation error bounds in Theorem 2.1 to at most $\epsilon^{2}$. We refer the reader to Section C.2 for a proof of Corollary 2.1.1.
Corollary 2.1.1.
To estimate $\kappa(f)$ using the iterates $\kappa_{n}$ generated by Algorithm 1 with diminishing step size, up to mean estimation error $\mathbb{E}\!\left[\left|{\kappa_{n}-\kappa(f)}\right|\right]\leq\epsilon$, we require
$n=O\left(\frac{1}{\epsilon^{2}}\right)O\left(\frac{\|V^{*}\|^{2}_{2}}{\Delta^{4}_{1}}\right).$
The sample complexity in Corollary 2.1.1 depends on $|S|$ via $\Delta_{1}$, and $\|V^{*}\|_{2}$. Here, $V^{*}\in\mathbb{R}^{|S|}$, and hence, $\|V^{*}\|^{2}_{2}=O(|S|^{2})$. The dependence of $\Delta_{1}$ on $|S|$ is more implicit, and relates to the mixing properties of the Markov chain induced by the policy. We make this dependence explicit in two specific examples. If the underlying Markov chain is a random walk on a complete graph (with $|S|$ vertices), then $\Delta_{1}=O(1)$, and hence we get ${O}(|S|^{2})$ dependence. On the other hand, if the underlying Markov chain is a random walk on a cycle graph (Levin and Peres, 2017, Section 12.3.1), then $\Delta_{1}=O({1}/{|S|^{2}})$, and hence, we get a dependence of ${O}(|S|^{10})$ in the sample complexity.
Remark 2.4.
Using Jensen’s inequality, we have $\mathbb{E}\!\left[\left|{\kappa_{n}-\kappa(f)}\right|\right]\leq\sqrt{\mathbb{E}\!\left[(\kappa_{n}-\kappa(f))^{2}\right]},$ which then gives a bound on the mean estimation error. In particular, for the diminishing step-sizes of the form $\alpha_{i}=\frac{\alpha}{i+h}$ for all $i\geq 1$, Theorem 2.1(b) gives
$\mathbb{E}\!\left[\left|{\kappa_{n}-\kappa(f)}\right|\right]\leq{O}\left(\frac{1}{\sqrt{n}}\right).$
Remark 2.5.
Consider the problem of estimating the standard deviation, $\sqrt{\kappa(f)}$. Mean-squared error for the estimator $\sqrt{\kappa_{n}}$ satisfies
$\mathbb{E}\!\left[\left(\sqrt{\kappa_{n}}-\sqrt{\kappa(f)}\right)^{2}\right]=\mathbb{E}\!\left[\frac{({\kappa_{n}}-{\kappa(f)})^{2}}{\left(\sqrt{\kappa_{n}}+\sqrt{\kappa(f)}\right)^{2}}\right]\leq\frac{1}{\kappa(f)}\mathbb{E}\!\left[\left({\kappa_{n}}-{\kappa(f)}\right)^{2}\right],$
which is at most ${O}(\frac{1}{n})$ for $\alpha_{k}=O(\frac{1}{n})$.
2.3 Estimating Stationary Variance
In the previous section, we saw an estimator for the asymptotic variance $\kappa(f)$, which captures the temporal correlations introduced due to the underlying Markov chain. Suppose you want to estimate the stationary variance, instead. Formally, this corresponds to estimating $v(f):=\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})^{2}\right].$ In Algorithm 2, we present a sequential (recursive) estimator for $v(f)$, that also satisfies all the properties listed in Table 1. Our estimator corresponds to SA update for the following equation:
$v(f)=\mathbb{E}_{\pi\!}\!\left[f^{2}(X)-f(X)\bar{f}\right].$
Input: Time horizon $n>0$, constant $c>0$, and step-size sequence $\left\{\alpha_{k}\right\}$.
Initialization: $\bar{f}_{0}=0$, and $v_{0}=0$.
for $k\leftarrow 1$ to $n$ do
      Observe $(X_{k},f(X_{k}))$.
      // Average reward $\bar{f}$ estimation
$\bar{f}_{k+1}=(1-\alpha_{k})\bar{f}_{k}+\alpha_{k}f(X_{k}).$
      // Stationary Variance $v(f)$ estimation
$v_{k+1}=(1-c\alpha_{k})v_{k}+c\alpha_{k}(f^{2}(X_{k})-f(X_{k})\bar{f}_{k})$
end for
Output : Estimates $v_{n}$, $\bar{f}_{n}$
Algorithm 2 Estimating Stationary Variance $v(f)$
Theorem 2.2.
For $\gamma>1$, consider Algorithm 2 with $c$ satisfying:
$c\bar{f}^{2}\leq 2\left(-(\gamma-1)+\sqrt{(\gamma-1)\left(\gamma-1+\gamma{\bar{f}}^{2}\right)}\right).$
Let $(\Theta^{S})^{T}:=[\bar{f}~{}~{}v(f)]$, and let $\eta=\sqrt{1+c^{2}\left(1+\bar{f}^{2}\right)}$. For a constant $B>1$, let $\xi_{1}:=3\left(1+\|\Theta^{S}\|_{2}\right)^{2}$, $\xi_{2}:=112B(1+\|\Theta^{S}\|_{2})^{2}$, and let $\Theta^{S}_{n}:=[\bar{f}_{n}~{}v_{n}]$ denote the vector of estimates at step $n$.
  1. (a)
    Let $\alpha_{i}=\alpha$ for all $i$, such that
    $\alpha<\min\left\{\frac{2}{\gamma},\frac{1}{28B\left(1+\frac{\eta^{2}}{\gamma}\right)}\right\}.$
    Then, for all $n\geq 1$,
    $\mathbb{E}\!\left[\|\Theta^{S}_{n}-\Theta^{S}\|^{2}_{2}\right]\leq\xi_{1}\left(1-\frac{\gamma\alpha}{2}\right)^{n}+\frac{\xi_{2}\alpha\eta^{2}}{\gamma}+\xi_{2}\alpha.$
  2. (b)
    Let $\alpha_{i}=\frac{\alpha}{i+h}$ for all $i\geq 1$, with $\alpha$ and $h$ chosen so that
    $h\geq\max\left\{1+\frac{\alpha\gamma}{2},1+28B\left(\frac{\eta^{2}\alpha}{\gamma}+\alpha+\frac{1}{\gamma}\right)\right\},\quad\text{and}\quad\alpha>\frac{2}{\gamma}.$
    Then, for all $n\geq 0$,
    $\mathbb{E}\!\left[\|\Theta^{S}_{n}-\Theta^{S}\|^{2}_{2}\right]$ $\leq\xi_{1}\left(\frac{h}{n+h}\right)^{\frac{\alpha\gamma}{2}}+\frac{5\xi_{2}e^{2}\eta^{2}(1+\gamma)\alpha^{2}}{(n+h)(\alpha\gamma-2)}+\frac{\xi_{2}\alpha}{n+h}.$
We refer the reader to Appendix C.3 for a complete proof of the Theorem above. It follows a 2-step approach. In the first step, we view Algorithm 2 as a linear SA update, and establish a contraction property for the average of the update matrices (Lemma C.1 in Appendix C.3.1). In the second step, we use a general bound on MSE for iterates of a linear SA satisfying the proven contraction property, that we prove in Theorem A.1. In fact, we use a similar 2-step approach to prove other results in this work. We describe this approach in details in Section 3.4, where we prove a MSE bound for estimating $\kappa(f)$ in a more general setting, presented in the following section.
Remark 2.6.
In Step 2 of the proof of Theorem 2.2, we could use an existing finite sample bound for MSE of a linear SA update in presence of Markov noise Srikant and Ying (2019). This would result in $O(\frac{\log n}{n})$ MSE bound in Theorem 2.2${}(b)$. However our tighter bound in Theorem A.1, that uses Poisson equation for handling the bias introduced due to presence of Markov dependence, results in $O(\frac{1}{n})$ MSE in Theorem 2.2${}(b)$ above.
Remark 2.7.
Notice that there is no dependence on the size of the state space $S$ in the bound of Theorem 2.2.
I.i.d. Variance Estimation
The corollary below shows that Algorithm 2 can also be used to estimate the variance in i.i.d. setting, achieving the optimal rate of convergence of $O(\frac{1}{n})$ in MSE. Moreover, the estimator satisfies all the desirable properties stated in Table 1 – it is recursive, does not require storing historical samples, and does not require knowing the total run length apriori.
Corollary 2.2.1.
For $n\in\mathbb{N}$, consider a sequence of i.i.d. observations $\left\{X_{i}\right\}_{i=1}^{n}$ from an unknown distribution $P$. Let $\bar{f}:=\mathbb{E}_{P\!}\!\left[f(X)\right]$. Define
$v(f):=\mathbb{E}_{P\!}\!\left[\left(f(X)-\mathbb{E}_{P\!}\!\left[f(X)\right]\right)^{2}\right]=\mathbb{E}_{P\!}\!\left[f^{2}(X)-f(X)\bar{f}\right].$
Then, the output $v_{n}$ of Algorithm 2 satisfies
$\mathbb{E}\!\left[\left(v_{n}-v(f)\right)^{2}\right]=O\left(\frac{1}{n}\right).$
The rate of convergence of $\Omega(\frac{1}{n})$ for MSE is tight for i.i.d. setting. In the special setting of i.i.d. noise with parametric distributions, this follows from the Cramér-Rao lower bound Nielsen (2013).
3 Variance Estimation: Linear Function Approximation
When the underlying state and action spaces are large, estimating $V$ function for each state-action pair may be computationally intractable. To address this, we consider evaluating an approximation of $V$ by its projection on a linear subspace spanned by a given fixed set of $d$ vectors $\{\tilde{\phi}_{1},\dots,\tilde{\phi}_{d}\}$, where $\tilde{\phi}_{i}\in\mathbb{R}^{S}$ for each $i\in[d]$. More concretely, for $s\in{\mathcal{S}}$, and $\theta\in\mathbb{R}^{d}$, we consider the following linear function approximation (LFA) of $V(s)$:
$V_{\theta}(s):=\phi(s)^{T}\theta,\quad\text{where}\quad\phi^{T}(s):=[\tilde{\phi}_{1}(s)\dots\tilde{\phi}_{d}(s)],$
$\phi^{T}(s)$ is the feature vector for state $s$, and $\phi(s)\in\mathbb{R}^{d}$. Recall that $S=|{\mathcal{S}}|$. Let $\Phi$ be a $S\times d$ matrix with $\tilde{\phi}_{i}$ being the $i^{\text{th}}$ column, and let
$W_{\Phi}=\{\Phi\theta:\theta\in\mathbb{R}^{d}\}$
denote the column space of $\Phi$. Then, $V_{\theta}=\Phi\theta$, where $V_{\theta}\in\mathbb{R}^{S\times 1}$ is an approximation for $V$ using $\theta$.
Assumption 1.
The matrix $\Phi$ is full rank, i.e., the set of feature vectors $\{\tilde{\phi}_{1},\dots,\tilde{\phi}_{d}\}$ are linearly independent. Additionally, $\|\phi(s)\|_{2}\leq 1$, for each $s\in{\mathcal{S}}$.
This is a standard assumption, and can be achieved by feature normalization, see Tsitsiklis and Van Roy (1999); Bertsekas and Tsitsiklis (1996). We now introduce some notation that will be used in this section. Let $\theta_{e}\in\mathbb{R}^{d}$ be the unique vector (if it exists) such that $\Phi\theta_{e}={\bf e}$. Define subspace $S_{\Phi,e}$ of $\mathbb{R}^{d}$ as
$S_{\Phi,e}:=\text{span}\left(\left\{\theta|\Phi\theta={\bf e}\right\}\right)=\begin{cases}\left\{c\theta_{e}|c\in\mathbb{R}\right\},\quad&\text{ if }{\bf e}\in W_{\Phi},\\ \{0\},\quad&\text{ otherwise. }\end{cases}$
Let $E$ be the subspace of $\mathbb{R}^{d}$ that is orthogonal complement (in $\ell_{2}$-norm) of $S_{\Phi,e}$, i.e.,
$E=\begin{cases}\{\theta\in\mathbb{R}^{d}:~{}\theta^{T}\theta_{e}=0\},\quad&\text{ if }{\bf e}\in W_{\Phi},\\ \mathbb{R}^{d},\quad&\text{ otherwise.}\end{cases}$
Observe that for $\theta\in E$, $\Phi\theta$ is a non-constant vector. This is because there does not exist $\theta\in E$ such that $\Phi\theta={\bf e}$. This follows since if ${\bf e}\in W_{\Phi}$, then the unique vector $\theta_{e}$ such that $\Phi\theta_{e}={\bf e}$, does not belong to $E$. On the other hand, if ${\bf e}\notin W_{\Phi}$, then there does not exist a vector $\theta\in\mathbb{R}^{d}$, hence in $E$, such that $\Phi\theta={\bf e}$.
Finally, let $\Pi_{2,E}$ denote the orthogonal projection of vectors in $\mathbb{R}^{d}$ (in $\ell_{2}$-norm) on the subspace $E$, i.e.,
$\Pi_{2,E}:=\begin{cases}I-\theta_{e}(\theta^{T}_{e}\theta_{e})^{-1}\theta^{T}_{e},\quad&\text{ if }{\bf e}\in W_{\Phi},\\ I,\quad&\text{ otherwise.}\end{cases}$
3.1 The Algorithm
We now propose an algorithm for estimating the asymptotic variance with LFA that, as in the previous section, uses the formulation of $\kappa$ in (8), with $V$ replaced by its approximation. In the setting considered here, the estimation of $V$ corresponds to an estimate for $\theta$ at each step; call it $\theta_{k}$. The corresponding estimate for $V$ is then $\Phi\theta_{k}$.
The algorithm in this setting is a modification of Algorithm 1, with  (10), (11), (12), and (13) replaced by (15), (16), (17), and (18) below, respectively.
$\delta_{k}:=f(X_{k})-\bar{f}_{k}+(\phi(X_{k+1})-\phi(X_{k}))^{T}\theta_{k},$ (15)
$\theta_{k+1}=\theta_{k}+\alpha_{k}\Pi_{2,E}\phi(X_{k})\delta_{k},$ (16)
$\widetilde{V}_{k+1}=(1-c_{2}\alpha_{k})\widetilde{V}_{k}+c_{2}\alpha_{k}\phi^{T}(X_{k})\theta_{k}$ (17)
$\kappa_{k+1}=(1-c_{3}\alpha_{k})\kappa_{k}+c_{3}\alpha_{k}\left(2f(X_{k})\phi^{T}(X_{k})\theta_{k}-2f(X_{k})\widetilde{V}_{k}-f^{2}(X_{k})+f(X_{k})\bar{f}_{k}\right)$ (18)
Here, updates in (15) and (16) correspond to SA for estimating $\theta$, adjusted with the projection on subspace $E$. As in Section 2, this projection step is to ensure convergence of the iterates $\widetilde{V}_{k}$, and is not required otherwise (Remark 2.1). Note that Algorithm 1 for the tabular setting is a special case with $d=S$ and $\Phi=I$, the identity matrix in $S\times S$ dimensions.
3.2 Convergence Rates
In this section, we present a finite-sample bound on the estimation error of the proposed algorithm with LFA. Recall the Markov chain $\mathcal{M}_{2}$ introduced in Section 2.2. Define
$\Delta_{2}:=\min\limits_{\|\theta\|_{2}=1,\theta\in E}~{}\theta^{T}\Phi^{T}D_{\pi}(I-P)\Phi\theta.$
As earlier, this can be shown to be strictly positive since for $\theta\in E$, $\Phi\theta$ is a non-constant vector. Finally, let $\Theta^{*T}:=[\bar{f}~{}\theta^{*T}~{}\widetilde{V}~{}\kappa^{*}]$, where
$\kappa^{*}=\mathbb{E}_{\pi\!}\!\left[2f(X)[\Phi\theta^{*}](X)-2f(X)\widetilde{V}-f^{2}(X)+f(X)\bar{f}\right],$ (19)
where $\Phi\theta^{*}\in\mathbb{R}^{S}$, and $[\Phi\theta^{*}](i)$ represents its $i^{th}$ component, $\widetilde{V}=\mathbb{E}_{\pi\!}\!\left[[\Phi\theta^{*}](X)\right]$, and $\theta^{*}$ is the unique vector in $E$ that is also a solution for
$\Phi\theta=\Pi_{D_{\pi},W_{\Phi}}T\Phi\theta,$
where $\Pi_{D_{\pi},W_{\Phi}}$ is the projection matrix onto $W_{\Phi}:=\left\{\Phi\theta|\theta\in\mathbb{R}^{d}\right\}$ with respect to $D_{\pi}$ norm, and $D_{\pi}$ is defined above (14). Specifically,
$\Pi_{D_{\pi},W_{\Phi}}=\Phi(\Phi^{T}D_{\pi}\Phi)^{-1}\Phi^{T}D_{\pi}.$
Further, $T$ is an operator that for a vector $V\in\mathbb{R}^{S}$, satisfies $TV=\mathcal{F}-\bar{f}{\bf e}+PV.$
Observe that $\kappa^{*}$ differs from $\kappa(f)$ in that $V$ and $\bar{V}$ in formulation (8) are replaced by the estimate in the subspace spanned by $\Phi$ and $\widetilde{V}$, respectively. The following theorem bounds the mean-squared distance between the estimate at time $n$ and the limit point $\kappa^{*}$ under constant as well as diminishing step sizes.
Theorem 3.1.
Consider the algorithm with $c_{1}$, $c_{2}$, and $c_{3}$ satisfying:
$c_{1}\geq\frac{1}{2\Delta_{2}}+\frac{\Delta_{2}}{2},\qquad\frac{5}{249}(5-2\sqrt{2})\Delta_{2}\leq c_{3}\leq\frac{5}{249}(5+2\sqrt{2})\Delta_{2},$
and
$c_{3}-\frac{498c^{2}_{3}}{7\Delta_{2}}+\frac{7\Delta_{2}}{498}\leq c_{2}\leq-3c_{3}+\frac{5}{83}\sqrt{498c_{3}\Delta_{2}-17\Delta^{2}_{2}}.$
Recall that $\eta$ was defined in (14), and $\Theta^{*}$ before (19). For a constant $B>1$, let $\xi_{1}:=3\left(1+\|\Theta^{*}\|_{2}\right)^{2}$, $\xi_{2}:=112B(1+\|\Theta^{*}\|_{2})^{2}$, and let $\Theta_{n}:=[\bar{f}_{n}~{}\theta^{T}_{n}~{}\widetilde{V}_{n}~{}\kappa_{n}]$ denote the vector of estimates at step $n$.
  1. (a)
    Let $\alpha_{i}=\alpha$ for all $i$, such that
    $\alpha<\min\left\{\frac{40}{\Delta_{2}},\frac{1}{28B\left(1+\frac{20\eta^{2}}{\Delta_{2}}\right)}\right\}.$
    Then, for all $n\geq 1$,
    $\mathbb{E}\!\left[\|\Theta_{n}-\Theta^{*}\|^{2}_{2}\right]\leq\xi_{1}\left(1-\frac{\Delta_{2}\alpha}{40}\right)^{n}+\frac{20\xi_{2}\alpha\eta^{2}}{\Delta_{2}}+\xi_{2}\alpha.$
  2. (b)
    Let $\alpha_{i}=\frac{\alpha}{i+h}$ for all $i\geq 0$, with $\alpha$ and $h$ chosen so that
    $h\geq\max\left\{1+\frac{\alpha\Delta_{2}}{40},1+28B\left(\frac{20\eta^{2}\alpha}{\Delta_{2}}+\alpha+\frac{20}{\Delta_{2}}\right)\right\},\quad\text{and}\quad\alpha>\frac{40}{\Delta_{2}}.$
    Then, for all $n\geq 0$,
    $\mathbb{E}\!\left[\|\Theta_{n}-\Theta^{*}\|^{2}_{2}\right]$ $\leq\xi_{1}\left(\frac{h}{n+h}\right)^{\frac{\alpha\Delta_{2}}{40}}+\frac{5\xi_{2}e^{2}\eta^{2}(20+\Delta_{2})\alpha^{2}}{(n+h)(\alpha\Delta_{2}-40)}+\frac{\xi_{2}\alpha}{n+h}.$
The above theorem bounds the MSE between the iterates of the algorithm, $\Theta_{n}$, and its limit point $\Theta^{*}$. In particular, since $\mathbb{E}\!\left[(\kappa_{n}-\kappa^{*})^{2}\right]\leq\mathbb{E}\!\left[\|\Theta_{n}-\Theta^{*}\|^{2}_{2}\right]$, the same bounds also hold for the MSE for $\kappa$ estimation. To prove Theorem 3.1, we view the proposed algorithm as a linear SA update, and prove appropriate contraction properties for the associated matrices (Lemma 3.3). We refer the reader to Section 3.4 for a proof of the theorem.
3.3 Approximation Error
Since we approximate the elements of the set $S_{f}$ of $V$ functions (defined after (1.3)) by a linear combination of basis vectors, we incur an approximation error in estimating $\kappa(f)$ using (8). In particular, $\kappa^{*}$ may differ from $\kappa(f)$ depending on the chosen approximation architecture.
In this section, we bound the squared approximation error: $(\kappa(f)-\kappa^{*})^{2}$. To this end, we first define the approximation error associated with approximating $V$ using LFA. Since each element of $S_{f}$ is a valid $V$ function, for any $\theta\in\mathbb{R}^{d}$, we define the error of approximating a point in $S_{f}$ by $\Phi\theta$ as the minimum weighted distance (in $D_{\pi}$-norm) of $\Phi\theta$ from $S_{f}$ Tsitsiklis and Van Roy (1999), i.e., $\inf\nolimits_{V\in S_{f}}\|\Phi\theta-V\|_{D_{\pi}}$. The minimum possible approximation error in $V$, due to the chosen architecture, is given by
$\mathcal{E}:=\inf\limits_{\theta\in\mathbb{R}^{d}}\inf\limits_{V\in S_{f}}\|\Phi\theta-V\|_{D_{\pi}}.$
This essentially captures the distance between the two sets, $S_{f}$ and the column space of $\Phi$. The following proposition bounds the squared difference between $\kappa^{*}$ and the asymptotic variance $\kappa(f)$, in terms of the approximation error for $V$.
Proposition 3.2.
Given $d$ basis vectors represented as columns of $\Phi$, there exists $\lambda\in(0,1)$ such that
$(\kappa^{*}-\kappa(f))^{2}\leq\frac{16~{}{\mathcal{E}}^{2}}{1-\lambda^{2}}.$
Notice that $\mathcal{E}$, and hence the RHS above, equals 00 if the chosen basis functions are such that the span of these basis functions intersects with the set $S_{f}$ of $V$ functions, i.e., if the approximation architecture is capable of representing some value function from $S_{f}$. This is particularly true for the tabular setting, where $d=S$ and the basis vectors are the standard basis in this dimension, spanning the entire space.
To prove Proposition 3.2, we first bound the approximation error in estimation of $V$ using the proposed Algorithm from Section 3.1. This is analogous to that for the value function in (Tsitsiklis and Van Roy, 1999, Theorem 3). Next, using this bound, we arrive at the bound in (3.2). We refer the reader to Section E for a proof of the proposition.
3.4 Proof of Theorem 3.1
The proof of Theorem 3.1 proceeds in two steps. In the first step, we formulate the proposed algorithm as a linear SA update, and establish an appropriate contraction property of the update matrices averaged under the stationary distribution $\pi$ of the underlying Markov chain. In the second step, we use a Lyapunov drift argument to establish a MSE bound for a linear SA update satisfying the contraction property from step one. To this end, we use Poisson equation again (recall its use in the estimator design), to bound the errors associated with the noisy updates (since the algorithm does not know $\pi$) at each step. We detail these below.
3.4.1 Step 1: Algorithm as a Linear SA
For $X_{k}\in{\mathcal{S}}$, $X_{k+1}\in{\mathcal{S}}$, and $k\in\mathbb{N}$, define
$f_{k}:=f(X_{k}),~{}~{}\phi_{k}=\phi(X_{k}).$
Recall that for each $x\in{\mathcal{S}}$, $\phi(x)\in\mathbb{R}^{d}$. Further, recall that $Y_{k}:=(X_{k},X_{k+1})$, and let $\vec{\bf 0}$ be the 00 vector in $\mathbb{R}^{d}$. Define $A(Y_{k})\in\mathbb{R}^{d+3}\times\mathbb{R}^{d+3}$,
$A(Y_{k}):=\begin{bmatrix}-c_{1}&\vec{\bf 0}^{T}&0&0\\ -\Pi_{2,E}\phi_{k}&\Pi_{2,E}\phi_{k}(\phi^{T}_{k+1}-\phi^{T}_{k})&\vec{\bf 0}&\vec{\bf 0}\\ 0&c_{2}\phi^{T}_{k}&-c_{2}&0\\ c_{3}f_{k}&2c_{3}f_{k}\phi^{T}_{k}&-2c_{3}f_{k}&-c_{3}\end{bmatrix},$
and $b(Y_{k})\in\mathbb{R}^{d+3}$ as
$b(Y_{k})^{T}:=\begin{bmatrix}c_{1}f_{k}~{}~{}~{}~{}f_{k}\left[\Pi_{2,E}\phi_{k}\right]^{T}~{}~{}~{}0~{}~{}~{}-{c_{3}}f^{2}_{k}\end{bmatrix}.$
Let $A=\mathbb{E}_{\pi\!}\!\left[A(Y_{k})\right]$ and $b=\mathbb{E}_{\pi\!}\!\left[b(Y_{k})\right]$ denote the stationary averages of $A(Y_{k})$ and $b(Y_{k})$, respectively (see Appendix D.1.1 for exact form of $A$ and $b$), and let $\Theta^{T}:=[\theta_{1}~{}\theta^{T}_{2}~{}\theta_{3}~{}\theta_{4}]$ with $\theta_{2}\in E$. Observe that $\Theta=\Theta^{*}$ (defined before (19)) is the unique solution to $A\Theta+b=0$ with $\theta\in E$ (see Appendix D.1.2 for a detailed justification). However, the algorithm doesn’t have access to matrices $A$ and $b$. Instead, we use SA to solve for $\Theta^{*}$, which corresponds to the following update rule at step $k+1$ with step size $\alpha_{k}$ (also, recall update from (1.4)):
${\Theta}_{k+1}={\Theta}_{k}+\alpha_{k}\left(A(Y_{k}){\Theta}_{k}+b(Y_{k})\right).$ (23)
Observe that the above equation with $\Theta^{T}_{k}:=[\bar{f}_{k}~{}\theta^{T}_{k}~{}\widetilde{V}_{k}~{}\kappa_{k}]$ coincides with the updates in the proposed algorithm (Algorithm in Section 3.1).
Next, once we have the linear SA form in (23), to establish the convergence of iterates, in Lemma 3.3, we prove that $A$, the average matrix, is contractive (under a semi-norm) for suitably chosen step-size constants $c_{1}$, $c_{2}$, and $c_{3}$. This result is crucial in establishing the finite sample bounds in Theorem 3.1.
Lemma 3.3.
Under the conditions on $c_{1}$, $c_{2}$, and $c_{3}$ from Theorem 3.1, the matrix $A$ satisfies
$\min\limits_{\begin{subarray}{c}\Theta\in\mathbb{R}\times E\times\mathbb{R}\times\mathbb{R},\|\Theta\|^{2}_{2}=1\end{subarray}}~{}-\Theta^{T}A\Theta~{}>~{}\frac{\Delta_{2}}{20}>0.$
We refer the reader to Appendix D.1.3 for a proof of Lemma 3.3. In the special case of $E=\mathbb{R}$, the above property can be shown to imply that $A$ is Hurwitz.
Remark 3.1.
The constant $20$ in the above Lemma is not special, and is a result of certain simplifying bounds used in the proof of the Lemma, which helped us in arriving at explicit bounds on $c_{1}$, $c_{2}$, and $c_{3}$ in Theorem 3.1. Modifying this constant would only change the constraints on the step size constants. We present the most general conditions (but slightly opaque) on $c_{1}$, $c_{2}$, and $c_{3}$ in Equations (68), (69), and (70) in Appendix D.1.3.
3.4.2 Step 2: Bounding MSE for Linear SA with Markov Noise
In Step 1, we formulated the updates in our estimator as a linear SA. One could use existing finite sample bounds on MSE for linear SA Srikant and Ying (2019), which require $A$ to be Hurwitz. However, Lemma 3.3 only guarantees this when $E=\mathbb{R}$, i.e., when $\bf{1}$ (the all $1$s vector) does not belong to the column space of $\Phi$. Recall that we do not make any such assumption.
Recently, Zhang et al. (2021) extend the MSE bound of Srikant and Ying (2019) for the setting when $E\neq\mathbb{R}$. Using (Zhang et al., 2021, Theorem 1), we could get MSE bound of $O(\frac{\log n}{n})$ at step $n$ for our estimator with $\alpha_{n}=O(\frac{1}{n})$. Instead, we conduct a tighter analysis of linear SA, in general. We use Poisson equation to get a tight control on the error terms in our analysis in Appendix A (Theorem A.1), which result in $O(\frac{1}{n})$ MSE bound, instead. Our bounds from Theorem A.1 immediately imply those in Theorem 3.1. We refer the reader to Section D.2 for details.
4 Vector Valued Rewards
Our algorithm and analysis easily extend to the setting where the function ${\bf f}$ is ${\mathfrak{d}}>1$ dimensional, i.e., ${\bf f}:{\mathcal{S}}\rightarrow\mathbb{R}^{\mathfrak{d}}$. We illustrate this with a discussion on the extension of Algorithm 1 for estimating the (per-step) asymptotic covariance matrix (defined next) and a brief proof sketch.
The (per-step) asymptotic covariance matrix for a function defined on states of a Markov chain starting from state $s$ is defined as
$\lim\limits_{n\rightarrow\infty}\operatorname{Cov}\left(\frac{1}{\sqrt{n}}\sum\limits_{l=0}^{n}{\bf f}(X_{l})~{}\bigg{|}~{}X_{0}=s\right),$
where, for a vector-valued function ${\bf r}:{\mathcal{S}}\rightarrow\mathbb{R}^{\mathfrak{d}}$ with mean $\bar{\bf r}:=\mathbb{E}_{\pi\!}\!\left[{\bf r}(X)\right]$, where ${\bar{\bf r}}\in\mathbb{R}^{\mathfrak{d}}$ (represented as column vectors),
$\operatorname{Cov}\left(\bf r\right):=\mathbb{E}\!\left[(\bf r-\bar{\bf r})(\bf r-\bar{\bf r})^{T}\right].$
For $i\in[{\mathfrak{d}}]:=\{1,\dots,{\mathfrak{d}}\}$ and $s\in{\mathcal{S}}$, let $f^{(i)}(s)$ denote the $i^{th}$ component of the (column) vector ${\bf f}(s)\in\mathbb{R}^{{\mathfrak{d}}}$ of the function value in state $s$, and let
$\bar{\bf f}:=\mathbb{E}_{\pi\!}\!\left[{\bf f}(X)\right]$
be the mean vector under $\pi$, the stationary distribution of $\mathcal{M}$. Then, $\bar{\bf f}\in\mathbb{R}^{\mathfrak{d}}$. For $i\in[{\mathfrak{d}}]$, let $\bar{f}^{(i)}\in\mathbb{R}$ denote its $i^{th}$ coordinate. Then, there exists a solution $V^{(i)}:{\mathcal{S}}\rightarrow\mathbb{R}$ ($V^{(i)}\in\mathbb{R}^{S}$) of the Poisson equation for $f^{(i)}-\bar{f}^{(i)}$ for each $i\in[{\mathfrak{d}}]$, i.e., for each $s\in{\mathcal{S}}$, and $i\in[{\mathfrak{d}}]$, $V^{(i)}(s)$ solves
$f^{(i)}(s)-\bar{f}^{(i)}=V^{(i)}(s)-\sum\limits_{s^{\prime}\in{\mathcal{S}}}P(s,s^{\prime})V^{(i)}(s^{\prime}).$
In fact, as in the settings of previous sections (${\mathfrak{d}}=1$ case), one can show that for each coordinate $i\in[{\mathfrak{d}}]$, if $V^{(i)}$ is a solution, then $V^{(i)}+c{\bf e}$ is also a solution. We denote by $V^{*(i)}$ the solution normalized so that $\mathbb{E}_{\pi}[V^{*(i)}]=0$.
As in Proposition 1.1, one can argue that the asymptotic covariance matrix defined above is independent of the starting state, and is given by
${\mathfrak{C}}({\bf f})=\lim\limits_{n\rightarrow\infty}~{}\operatorname{Cov}\left(\frac{1}{\sqrt{n}}\sum\limits_{l=0}^{n-1}{\bf f}(X_{l})~{}\bigg{|}~{}X_{0}\sim\pi\right).$
For $s\in{\mathcal{S}}$, let ${\bf V}(s)\in\mathbb{R}^{\mathfrak{d}}$ be
${\bf V}^{T}(s)=[V^{(1)}(s)~{}\dots~{}V^{({\mathfrak{d}})}(s)].$
Moreover, given (column) vectors ${\bf u}$ and ${\bf v}$ in $\mathbb{R}^{\mathfrak{d}}$, let ${\bf u}\otimes{\bf v}^{T}$ denote their outer product, resulting in a matrix in $\mathbb{R}^{{\mathfrak{d}}\times{\mathfrak{d}}}$. Then, ${\mathfrak{C}}(\bf f)$, which is a matrix in $\mathbb{R}^{{\mathfrak{d}}\times{\mathfrak{d}}}$, has the following equivalent representation.
${\mathfrak{C}}({\bf f})=\mathbb{E}_{\pi\!}\!\left[\left({\bf f}(X)-\bar{\bf f}\right)\otimes{\bf V}^{T}(X)\right]$ $+\mathbb{E}_{\pi\!}\!\left[{\bf V}(X)\otimes\left({\bf f}(X)-\bar{\bf f}\right)^{T}\right]$
$\quad-\mathbb{E}_{\pi\!}\!\left[\left({\bf f}(X)-\bar{\bf f}\right)\otimes\left({\bf f}(X)-\bar{\bf f}\right)^{T}\right].$ (24)
This can be proven using arguments similar to those in the proof for the reformulation for $\kappa(f)$ in (4) in Proposition 1.1. Notice that the expression for the diagonal entries of ${\mathfrak{C}}({\bf f})$ in the formulation in (24) is exactly same as that for $\kappa(f)$ in (4).
Next, let $\bar{\bf V}\in\mathbb{R}^{\mathfrak{d}}$ be $\bar{\bf V}:=\mathbb{E}_{\pi\!}\!\left[{\bf V}(X)\right]$. As in (8), we now re-express (24).
${\mathfrak{C}}({\bf f})$ $=\mathbb{E}_{\pi\!}\!\left[{\bf f}(X)\otimes{\bf V}^{T}(X)\right]+\mathbb{E}_{\pi\!}\!\left[{\bf V}(X)\otimes{\bf f}^{T}(X)\right]-\mathbb{E}_{\pi\!}\!\left[{\bf f}(X)\otimes\bar{\bf V}^{T}\right]$
$\qquad\qquad-\mathbb{E}_{\pi\!}\!\left[\bar{\bf V}\otimes{\bf f}^{T}(X)\right]+\mathbb{E}_{\pi\!}\!\left[{\bf f}(X)\otimes{\bf f}^{T}(X)\right]+\mathbb{E}_{\pi\!}\!\left[{\bf f}(X)\otimes\bar{\bf f}^{T}\right].$ (25)
Again, notice that the expression for the diagonal entries of ${\mathfrak{C}}({\bf f})$ in the above formulation is same as that for $\kappa(f)$ in (8). Our SA algorithm (Algorithm 1) can be extended to estimate ${\mathfrak{C}}({\bf f})$ using the formulation in (25). In particular, the updates in (9), (10), (11), and (13) in the tabular setting can be replaced by the following update equations. Here, for $k\in\mathbb{N}$ and $s\in{\mathcal{S}}$, $\bar{\bf f}_{k}$, ${\bf\delta}_{k}$, ${\bf V}_{k}(s)$, and $\bar{\bf V}_{k}$ are vectors in $\mathbb{R}^{\mathfrak{d}}$, and ${\mathfrak{C}}_{k}$ are matrices in $\mathbb{R}^{{\mathfrak{d}}\times{\mathfrak{d}}}$.
$\bar{\bf f}_{k+1}=\bar{\bf f}_{k}+c_{1}\alpha_{k}({\bf f}(X_{k})-\bar{\bf f}_{k}),$
${\bf\delta}_{k}:={\bf f}(X_{k})-\bar{\bf f}_{k}+({\bf V}_{k}(X_{k+1})-{\bf V}_{k}(X_{k})),$
${\bf V}_{k+1}(x)=\begin{cases}{\bf V}_{k}(x)+{\alpha_{k}}{\delta}_{k}\left(1-1/{S}\right),\quad&\text{ for }x=X_{k},\\ {\bf V}_{k}(x)-{\alpha_{k}}{\delta}_{k}/{S},\quad&\text{ otherwise},\end{cases}$
$\bar{\bf V}_{k+1}=\bar{\bf V}_{k}+c_{2}\alpha_{k}({\bf V}_{k}(X_{k})-\bar{\bf V}_{k}),$
${\mathfrak{C}}_{k+1}$ $=(1-c_{3}\alpha_{k}){\mathfrak{C}}_{k}+c_{3}\alpha_{k}\bigg{(}{\bf f}(X_{k})\otimes{\bf V}^{T}_{k}(X_{k})+{\bf V}_{k}(X_{k})\otimes{\bf f}^{T}(X_{k})-{\bf f}(X_{k})\otimes\bar{\bf V}^{T}_{k}$
$\qquad\qquad\qquad\qquad\qquad\qquad-\bar{\bf V}_{k}\otimes{\bf f}^{T}(X_{k})+{\bf f}(X_{k})\otimes{\bf f}^{T}(X_{k})+{\bf f}(X_{k})\otimes\bar{\bf f}^{T}_{k}\bigg{)}.$
The linear function approximation setting of Section 3 can also be extended similarly to the current setup, with a set of $d$ feature vectors for each component in $[{\mathfrak{d}}]$. Let’s denote the feature vectors for approximating $V^{(i)}$ by $\{\tilde{\phi}^{(i)}_{1},\dots,\tilde{\phi}^{(i)}_{d}\}$, and let $\Phi^{(i)}$ denote the $S\times d$ matrix with these $d$ vectors as columns. Algorithm from Section 3.1 can be appropriately modified to estimate ${\mathfrak{C}}({\bf f})$ using SA for (25). The resulting algorithm can again be analyzed by formulating it as a linear SA for each pair $(i,j)$ of ${\mathfrak{C}}({\bf f})$, for $i\in[{\mathfrak{d}}]$, $j\in[{\mathfrak{d}}]$, and $i\leq j$, giving similar guarantees on mean-squared estimation error for each component of ${\mathfrak{C}}({\bf f})$ as in Theorems 2.1 and 3.1.
5 Applications in Reinforcement Learning
Consider an average reward reinforcement learning (RL) framework, where the transition dynamics of the underlying Markov decision process (MDP) are unknown Bertsekas and Tsitsiklis (1996); Sutton and Barto (2018). An RL algorithm can obtain a sample of the MDP under a given policy, which specifies how actions are chosen in a given state. The traditional goal in an average reward RL problem is to find a policy that maximizes the long run average reward. While the need to optimize over average reward is well motivated, for applications in safety-critical domains like healthcare and finance, it is also crucial to control adverse outcomes. To address this, within the domain of risk-sensitive MDPs, several risk measures have been considered, namely variance Sobel (1982); Filar et al. (1989), conditional value-at-risk Chow and Ghavamzadeh (2014), exponential utility Whittle (1990), the general class of coherent risk measures Tamar et al. (2015), and cumulative prospect theory Prashanth et al. (2016), which is not coherent. The choice of the risk measure largely depends on the application at hand.
We consider variance as a risk measure. Several previous works incorporate variance in a constrained setting, where the goal is to maximize the average reward (which is an expectation), while ensuring a certain bound on the variance. This is the so-called “mean-variance tradeoff”, considered in the seminal work of Markowitz Markowitz (1952), and later in MDP contexts, cf. Mandl (1971); Sobel (1982); Filar et al. (1989). An alternative to such a formulation is to consider the exponential utility formulation, see Arrow (1971); Howard and Matheson (1972), where one optimizes the exponential. The constrained formulation is preferred over the exponential utility for two reasons. First, the mean-variance tradeoff can be controlled directly through a parameter that is a bound on the variance, while this trade-off is implicit in an exponential utility formulation. Second, the algorithms for the latter formulation do not extend easily when one considers feature-based representations and function approximation, see (L.A. and Fu, 2022, Section 7.2) for a detailed discussion.
Broadly, for a given policy, two different notions of variance are suggested for an average reward MDP in Filar et al. (1989). The first notion is the asymptotic variance, while the second one is the per-period stationary variance. The latter has been studied in an RL setting in Prashanth and Ghavamzadeh (2016), while the former has hardly been investigated in the literature, to the best of our knowledge. We use the asymptotic variance as a measure of the risk associated with a given policy, which is the variance of the random variable whose mean is typically optimized. This asymptotic variance can be decomposed into two additive terms, where the first term coincides with per-step variance, while the other term involves correlations between states across time periods. In a setting where the state sequence is independent, the second term is zero, while it is not in any non-trivial MDP. A mean-variance optimization formulation would consider maximizing the expectation of a random variable which represents the long-run average reward, and it is natural to consider the variance of this random variable; also see Section 5.2.
In this section, we consider the problem of policy evaluation for asymptotic variance, where the goal is to estimate the asymptotic variance associated with the Markov chain induced by a given stationary policy. This problem of estimating the asymptotic variance is a vital sub-problem in mean-variance policy optimization, for instance, as a critic in an actor-critic framework, cf. Prashanth and Ghavamzadeh (2016). For the discounted RL setting, a TD type algorithm for estimating variance has been proposed/analyzed in Tamar et al. (2013), while a TD algorithm with provable finite sample guarantees, which caters to the variance risk measure in an average reward RL setting, is not present in the literature to the best of our knowledge.
The rest of this section is organized as follows. We begin by formally introducing the average reward RL setting and the problem setup in Section 5.1. In Section 5.2, we motivate the choice of asymptotic variance as a natural measure of risk associated with a given policy in this framework and discuss the related literature in Section 5.3. We design a novel TD-like linear SA algorithm for policy evaluation in a tabular setting in Section 5.4. Finally, we conclude in Section 5.5 with the finite sample error bounds for the proposed algorithm, proving ${O}(\frac{1}{n})$ rate of convergence for the mean-squared error, where $n$ is the time step. We defer the algorithm and its finite sample error bounds for the setting of linear function approximation in RL to Appendix F.
5.1 Setup
Consider an infinite-horizon, average-reward MDP specified by $({\mathcal{S}},{\mathcal{A}},r,p)$, where ${\mathcal{S}}=\left\{1,\dots,\left|{S}\right|\right\}$ denotes the finite state-space, and ${\mathcal{A}}=\left\{1,\dots,\left|{A}\right|\right\}$ denotes the action space. At each time $k$, the agent is in state $S_{k}\in{\mathcal{S}}$, takes an action $A_{k}\in{\mathcal{A}}$, receives a reward $r(S_{k},A_{k})$, and transitions to state $S_{k+1}$. Here, $r:{\mathcal{S}}\times{\mathcal{A}}\rightarrow\mathbb{R}$, and the next state $S_{k+1}$ is sampled according to $p(S_{k},\cdot,A_{k})$, where, $p:{\mathcal{S}}\times{\mathcal{S}}\times{\mathcal{A}}\rightarrow[0,1]$, is the map that for states $s,s^{\prime}$ and action $a$ associates probability $p(s,s^{\prime},a)$ with the transition from state $s$ to $s^{\prime}$, when action $a$ is taken.
A stationary policy $\mu:{\mathcal{S}}\rightarrow\Sigma_{\left|{{\mathcal{A}}}\right|}$ is a map from state space to a probability-simplex in $\mathbb{R}^{\left|{{\mathcal{A}}}\right|}$, i.e., it maps each state to a distribution over the actions. Given a stationary policy $\mu$ and a state $x$, the associated probability of transitioning from state $x$ to $x^{\prime}$ is given by
$P_{\mu}(x,x^{\prime})=\sum\nolimits_{a\in{\mathcal{A}}}\mu(a|x)p(x,x^{\prime},a).$
Assumption 2.
The sets ${\mathcal{S}}$ and ${\mathcal{A}}$ are finite. Under the stationary policy $\mu$, the induced Markov chain ${\mathcal{M}}_{1}$ with state space in ${\mathcal{S}}$ and transition probabilities given by $P_{\mu}$ is irreducible and aperiodic.
This is a standard assumption in literature (see Tsitsiklis and Van Roy (1999); Bertsekas (2012)), and guarantees that each state is visited infinitely often. As a consequence of Assumption 2, we have a unique stationary distribution associated with ${\mathcal{M}}_{1}$, and the Markov chain starting from any initial distribution, converges to the stationary distribution geometrically-fast (Levin and Peres, 2017, Section 4.3). In particular, let $\pi_{\mu}$ denote the unique stationary distribution on ${\mathcal{S}}$ that satisfies $\pi^{T}_{\mu}P_{\mu}=\pi^{T}_{\mu}$.
Next, let $X_{k}:=(S_{k},A_{k})$. Observe that the process ${\mathcal{M}}_{2}:=\left\{X_{k}\right\}$ is also a Markov chain with finite states. Let its state space be denoted by ${\mathcal{X}}\subset{\mathcal{S}}\times{\mathcal{A}}$. For any two states $x_{1}:=(s_{1},a_{1})$ and $x_{2}:=(s_{2},a_{2})$ in ${\mathcal{X}}$, the probability of transitioning from $x_{1}$ to $x_{2}$ under ${\mathcal{M}}_{2}$ is given by
$P_{2}(x_{1},x_{2}):=p(s_{1},s_{2},a_{1})\mu(a_{2}|s_{2}).$ (26)
For $s\in{\mathcal{S}}$ and $a\in{\mathcal{A}}$, define $d_{\mu}(s,a):=\pi_{\mu}(s)\mu(a|s).$
Remark 5.1.
Under Assumption 2, ${\mathcal{M}}_{2}$ has a unique stationary distribution $d_{\mu}$ defined above, and it mixes geometrically-fast. This follows from (Bhatnagar and Lakshmanan, 2016, Proposition 1).
Notation
Let $D_{\mu}\in\mathbb{R}^{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}}\right|}\times\mathbb{R}^{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}}\right|}$ be the diagonal array of stationary distribution $d_{\mu}$, i.e.,
$D_{\mu}((s,a),(s,a)):=\begin{cases}d_{\mu}(s,a),&\text{ for }(s,a)\!\in\!{\mathcal{S}}\times{\mathcal{A}},\\ 0,&\text{ otherwise}.\end{cases}$
For vectors $x$ and $y$ in $\mathbb{R}^{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}}\right|}$, let $<x,y>_{D_{\mu}}:=x^{T}D_{\mu}y,$ and $\|x\|_{D_{\mu}}:=x^{T}D_{\mu}x$ denote the $D_{\mu}$ weighted inner product and the induced norm, respectively.
Average Reward
The long-term per-step expected reward accumulated by a stationary policy $\mu$ starting from state $s$ is given by
$\lim\limits_{n\rightarrow\infty}~{}\frac{1}{n}\mathbb{E}\!\left[\sum\limits_{k=0}^{n-1}r(S_{k},A_{k})|S_{0}=s\right].$ (27)
Here, the expectation is with respect to both the transition dynamics of the environment, as well as the randomness in the policy $\mu$. It is well known that under Assumption 2, this per-step expected reward in (27) is a constant that is independent of the starting state Bertsekas (2012). We denote this quantity by $J_{\mu}$. It can be shown that
$J_{\mu}=\mathbb{E}_{d_{\mu}\!}\!\left[r(S,A)\right]=\sum\limits_{x\in{\mathcal{S}},a\in{\mathcal{A}}}\pi_{\mu}(x)\mu(a|x)r(x,a).$ (28)
Here, $\mathbb{E}_{d_{\mu}\!}\!\left[\cdot\right]$ denotes the expectation when $(S,A)$ is sampled according to the distribution $d_{\mu}$, which is defined after (26).
Asymptotic Variance
Observe that $r(\cdot,\cdot)$ is a function defined on the state space of $\mathcal{M}_{2}$, which is fixed by the given policy $\mu$. In this work, we consider the problem of estimating the associated asymptotic variance,
$\lim\limits_{T\rightarrow\infty}~{}\frac{1}{T}\operatorname{Var}\left[\sum\limits_{k=0}^{T-1}r(S_{k},A_{k})\bigg{|}S_{0}=s\right].$ (29)
It can be shown to be a constant, independent of the starting state. We denote it by $\kappa_{\mu}$. Moreover (Proposition 1.1),
$\kappa_{\mu}=\lim\limits_{n\rightarrow\infty}\operatorname{Var}\left[\frac{1}{\sqrt{n}}\sum\limits_{k=0}^{n-1}r(S_{k},A_{k})\bigg{|}(S_{0},A_{0})\sim d_{\mu}\right].$
Q Function
Since $J_{\mu}=\mathbb{E}_{d_{\mu}\!}\!\left[r(S,A)\right]$, there exists a solution of the following Poisson equation or Bellman equation (Douc et al., 2018, Section 21.2) $Q_{\mu}:{\mathcal{X}}\rightarrow\mathbb{R}$ such that for each $(s,a)\in{\mathcal{X}}$, and for $P_{2}$ given in (26),
$r(s,\!a)\!-\!J_{\mu}=~{}$ $Q_{\mu}(s,a)-\!\sum\limits_{\!(\!s^{\prime}\!,a^{\prime}\!)\in{\mathcal{X}}\!}\!P_{2}((s,a)),\!(s^{\prime},a^{\prime}))Q_{\mu}(s^{\prime},a^{\prime}).$ (30)
Let $Q_{\mu}$ be a solution to (30). Then, $Q_{\mu}+c{\bf e}$ for any constant $c\in\mathbb{R}$ is also a solution. Let $Q^{*}_{\mu}$ denote the solution normalized so that it is orthogonal to ${\bf e}$, i.e., $Q^{*T}_{\mu}{\bf e}=0$.
Remark 5.2.
Note that the equivalent formulations for $\kappa_{\mu}$ from Proposition 1.1 also hold with $r$, $J_{\mu}$, $Q_{\mu}$, and $d_{\mu}$ replacing $f$, $\bar{f}$, $V$, and $\pi$, respectively.
In this section, our goal is to estimate the asymptotic variance associated with the Markov chain induced by a given stationary policy, i.e., estimate $\kappa_{\mu}$. Note that this goal is orthogonal to that of estimating the variance of iterates of any particular learning algorithm like temporal-difference (TD), or designing algorithms with minimum variance of the iterates of the algorithm, studied extensively in Devraj and Meyn (2017); Yin and Wang (2020); Chen et al. (2020a); Hao et al. (2021); see Section 1.2 for a discussion and an example highlighting this distinction.
5.2 Motivation for Asymptotic Variance
As discussed in the beginning of Section 5, different notions of variance have been considered in the risk-sensitive RL literature. We now motivate the choice of asymptotic variance in a risk-sensitive setting
First, in the average reward setting, a classical goal is to optimize the long-term expected cumulative reward, which corresponds to expectation of the random variable $\sum_{k}r(S_{k},A_{k})$. The asymptotic variance considered in this work, corresponds to the variance of the same random variable. Since both the mean and variance of $\sum_{k}r(S_{k},A_{k})$ summed up to $n$ terms are $O(n)$, we equivalently consider $\frac{1}{n}$ scaling of both the mean and variance to arrive at $J_{\mu}$ and $\kappa_{\mu}$ in (27) and (29), respectively.
Second, consider an investor receiving return $r(S_{k},A_{k})$ on an investment at each time $k$. The total return of the investor in $n$ steps is $\sum_{k=1}^{n}r(S_{k},A_{k})$. A risk-averse investor would aim to maximize the average cumulative return subject to its variance being small. The variance on the return at each time is $\mathbb{E}_{d_{\mu}\!}\!\left[(r(S,A)-J_{\mu})^{2}\right]$. If the sequence $\left\{r(S_{k},A_{k})\right\}_{k\geq 1}$ were i.i.d. according to $d_{\mu}$, the variance of total reward over $n$ steps is $n\mathbb{E}_{d_{\mu}\!}\!\left[(r(S,A)-J_{\mu})^{2}\right]$. This corresponds to the first term in (3). However, since the returns are Markovian, and thus are correlated across time steps, this term captures only the variance over individual times, and does not capture the correlations across time steps. The second term in (3) captures the covariance of returns across time, and so the variance of the total reward $\sum_{k=1}^{n}r(S_{k},A_{k})$ is approximately $n\kappa_{\mu}$. In other words, the per step variance can be modeled by $\lim\limits_{n\rightarrow\infty}\frac{1}{n}\operatorname{Var}[\sum_{k=1}^{n}r(S_{k},A_{k})]=\kappa_{\mu}$.
Finally, it is important to emphasize that prior research on variance-constrained average reward RL, cf. Prashanth and Ghavamzadeh (2016), treats the first term in (3) as a surrogate for variance. In a setting where the state sequence is independent, the second term is zero, while in a Markovian context, the correlation between time steps is non-negligible, and should not be overlooked.
5.3 Related Literature
In this subsection, we briefly discuss the RL literature that is relevant to our work.
MDPs: Average Reward and Risk-sensitivity
MDPs have a long history. We refer the reader to classical books Puterman (1994); Bertsekas (2012) for a textbook introduction to MDPs in general, and to Howard (1960); Blackwell (1962); Brown (1965); Veinott (1966); Arapostathis et al. (1993) for an introduction to average-reward MDPs, in particular. Risk-sensitive objectives have also been well-studied in the MDP setting. For instance, see Sobel (1982); Filar et al. (1995); Mannor and Tsitsiklis (2013) for variance in discounted and average reward MDPs, Borkar and Meyn (2002); Borkar and Jain (2010); Whittle (1990) for the exponential utility formulation, Ruszczyński (2010) for Markov risk measures, Chow and Ghavamzadeh (2014) for conditional value-at-risk (CVaR). However, in a MDP setting, algorithms require complete knowledge of the underlying model, which may not be feasible in many practical applications.
Risk-neutral RL
In the risk-neutral RL setting, expected value is the sole objective. TD type algorithms have been proposed for policy evaluation in discounted as well as average reward settings, and their asymptotic convergence is shown in Tsitsiklis and Van Roy (1996, 1999), respectively. TD algorithms have also been used in actor-critic style algorithms for solving the problem of control, cf. Konda and Tsitsiklis (2003); Bhatnagar et al. (2009). Asymptotic convergence of the classical Q-learning algorithm was established in Borkar and Meyn (2000); Tsitsiklis (1994). In the non-asymptotic regime, finite-sample mean-square convergence bounds for classical discounted setting algorithms such as TD, TD($\lambda$), n-step TD, and Q-learning, have been developed in Chen et al. (2021). On the other hand, in the average reward setting, finite-sample bounds for TD are derived in Zhang et al. (2021).
Risk-sensitivity in RL
In a risk-sensitive optimization setting, the goal is either to optimize the usual expected value objective, while factoring a risk measure in a constraint, or optimize a risk measure in the objective. Tail-based risk measures such as variance, CVaR are meaningful to consider as a constraint, while risk measures such as exponential utility and prospect theory can be considered as the optimization objective, since they consider the entire distribution. In the context of RL, variance as a risk measure has been studied earlier in a discounted reward MDP setting in Mihatsch and Neuneier (2002); La and Ghavamzadeh (2013), in a stochastic shortest path and discounted settings Tamar and Mannor (2013); Tamar et al. (2013, 2012, 2016), and in an average reward MDP setting in Prashanth and Ghavamzadeh (2016). Exponential utility has been explored in an average reward RL setting earlier, see Borkar (2010) for a survey and Moharrami et al. (2022) for a recent contribution. Huang et al. (2021); Ding and Feinberg (2022); Köse and Ruszczyński (2021) study CVaR as a risk measure in the discounted setting. Other risk measures such as more general coherent risk measures and cumulative prospect theory have been explored in an RL setting in the literature, and some representative works include Prashanth (2014); Prashanth et al. (2016); Markowitz et al. (2023); Ruszczynski and Yang (2023).
5.4 The Algorithm
We now present a TD-type algorithm for estimating $\kappa_{\mu}$. Recall that $r(\cdot,\cdot)$ is a function defined on ${\mathcal{X}}$, states of the discrete time Markov chain $\mathcal{M}_{2}$ (introduced above (26)), with stationary expectation $J_{\mu}$. Further, we have from (30) that $Q_{\mu}$ is a solution for the corresponding Poisson equation. Let $\bar{Q}_{\mu}:=\mathbb{E}_{d_{\mu}\!}\!\left[Q_{\mu}(S,A)\right]$. Then,
$\kappa_{\mu}=\mathbb{E}_{d_{\mu}\!}\!\left[2r(S,A)Q_{\mu}(S,A)-2r(S,A)\bar{Q}_{\mu}-r^{2}(S,A)+r(S,A)J_{\mu}\right].$ (31)
This immediately follows from the formulation in (8) with $r(\cdot,\cdot)$, $J_{\mu}$, $Q_{\mu}(\cdot,\cdot)$, $\bar{Q}_{\mu}$, and $d_{\mu}$ replacing $f(\cdot)$, $\bar{f}$, $V(\cdot)$, $\bar{V}$, and $\pi$, respectively. Algorithm 3, presented next, corresponds to a linear SA algorithm for estimating $\kappa_{\mu}$ using the above formulation.
Input: Time horizon $n>0$, constants $c_{1}>0$, $c_{2}>0$, $c_{3}>0$ and step-size sequence $\left\{\alpha_{k}\right\}$.
Initialization: $\bar{r}_{0}=0$, $Q_{0}=\vec{\bf 0}$, $\bar{Q}_{0}=0$, and $\kappa_{0}=0$.
for $k\leftarrow 1$ to $n$ do
      Observe ($S_{k},A_{k},r(S_{k},A_{k}),S_{k+1},A_{k+1}$).
      // Average reward $J_{\mu}$ estimation
$\bar{r}_{k+1}=\bar{r}_{k}+c_{1}\alpha_{k}(r(S_{k},A_{k})-\bar{r}_{k}).$ (32)
      // $Q_{\mu}$ estimation
$\delta_{k}:=r(S_{k},A_{k})-\bar{r}_{k}+Q_{k}(S_{k+1},A_{k+1})-Q_{k}(S_{k},A_{k}),$ (33)
$Q_{k+1}(s,a)=\begin{cases}Q_{k}(s,a)+\alpha_{k}\delta_{k}\left(1-{1}/{|{\mathcal{S}}||{\mathcal{A}}|}\right),&\text{ for }(s,a)=(S_{k},A_{k}),\\ Q_{k}(s,a)-{\alpha_{k}\delta_{k}}/{|{\mathcal{S}}||{\mathcal{A}}|},&\text{ otherwise}.\end{cases}$ (34)
       // $\bar{Q}_{\mu}$ estimation
$\bar{Q}_{k+1}:=\bar{Q}_{k}+c_{2}\alpha_{k}(Q_{k}(S_{k},A_{k})-\bar{Q}_{k}),$ (35)
      // Asymptotic Variance $\kappa_{\mu}$ estimation
$\kappa_{k+1}=(1-c_{3}\alpha_{k})\kappa_{k}+c_{3}\alpha_{k}\bigg{(}2r(S_{k},A_{k})Q_{k}(S_{k},A_{k})-2r(S_{k},A_{k})\bar{Q}_{k}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-r^{2}(S_{k},A_{k})+r(S_{k},A_{k})\bar{r}_{k}\bigg{)}.$ (36)
end for
Output : Estimates $\kappa_{n}$, $\bar{Q}_{n}$, $Q_{n}$, $\bar{r}_{n}$
Algorithm 3 Estimating $\kappa_{\mu}$: Tabular Setting
5.5 Convergence Rates
We now bound the estimation error of Algorithm 3. Let $Y_{k}:=(S_{k},A_{k},S_{k+1},A_{k+1}).$ Observe that $\mathcal{M}_{3}:=\left\{Y_{k}\right\}_{k\geq 1}$ is a Markov chain that mixes geometrically fast (Bhatnagar and Lakshmanan, 2016, Proposition 1), and has a unique stationary distribution. Let $\mathcal{R}\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$ be the vector of rewards, and $r_{\max}:=\|\mathcal{R}\|_{\infty}$. Without loss of generality, for simplicity of presentation, we assume that $r_{\max}=1$. Recall that $D_{\mu}$ denotes the diagonal matrix of the stationary distribution $d_{\mu}$ of ${\mathcal{M}}_{2}$. Define
$\tilde{\Delta}_{1}:=\min\!\left\{\!{\bf v}^{T}\!D_{\mu}(I\!-\!P_{2}){\bf v}~{}|~{}{\bf v}\!\in\!\mathbb{R}^{S},\|{\bf v}\|_{2}=1,{\bf v}^{T}\!{\bf 1}=0\right\},$
where $P_{2}$ is the transition matrix for $\mathcal{M}_{2}$. It can be argued that $\tilde{\Delta}_{1}>0$ (Tsitsiklis and Van Roy, 1997, Lemma 7).
For the step-size constants $c_{1}$, $c_{2}$, and $c_{3}$ (inputs to Algorithm 3), define
$\eta:=\left(c^{2}_{1}+5+2c^{2}_{2}+10c^{2}_{3}\right)^{\frac{1}{2}}.$
Let $Q^{*}_{\mu}$ be the Q Function orthogonal to ${\bf 1}$ in $\ell_{2}$-norm, $\bar{Q}^{*}_{\mu}:=\mathbb{E}_{d_{\mu}\!}\!\left[Q^{*}_{\mu}(S,A)\right]$, and define $\Theta_{\mu}:=[J_{\mu}~{}Q^{*T}_{\mu}~{}\bar{Q}^{*}_{\mu}~{}\kappa_{\mu}].$ The following theorem bounds the estimation error of Algorithm 3, which immediately implies ${O}(\frac{1}{n})$ convergence rate for the mean-squared estimation error.
Theorem 5.1.
Consider Algorithm 3 with $c_{1}$, $c_{2}$, and $c_{3}$ satisfying:
$c_{1}\geq\frac{1}{2\tilde{\Delta}_{1}}+\frac{\tilde{\Delta}_{1}}{2},\qquad\frac{5}{249}(5-2\sqrt{2})\tilde{\Delta}_{1}\leq c_{3}\leq\frac{5}{249}(5+2\sqrt{2})\tilde{\Delta}_{1},$
and
$c_{3}-\frac{498c^{2}_{3}}{7\tilde{\Delta}_{1}}+\frac{7\tilde{\Delta}_{1}}{498}\leq c_{2}\leq-3c_{3}+\frac{5}{83}\sqrt{498c_{3}\tilde{\Delta}_{1}-17\tilde{\Delta}^{2}_{1}}.$
Recall that $\eta$ and $\Theta_{\mu}$ were defined below (5.5). For a constant $B>1$, let $\xi_{1}:=3(1+\|\Theta_{\mu}\|_{2})^{2}$, $\xi_{2}:=112B(1+\|\Theta_{\mu}\|_{2})^{2}$, and let $\Theta_{n}:=[\bar{r}_{n}~{}Q^{T}_{n}~{}\bar{Q}_{n}~{}\kappa_{n}]$ denote the vector of iterates of Algorithm 3 at step $n$.
  1. (a)
    Let $\alpha_{i}=\alpha$ for all $i$, such that
    $\alpha<\min\left\{\frac{40}{\tilde{\Delta}_{1}},\frac{1}{28B\left(1+\frac{20\eta^{2}}{\tilde{\Delta}_{1}}\right)}\right\}.$
    Then, for all $n\geq 1$,
    $\mathbb{E}\!\left[\|\Theta_{n}-\Theta_{\mu}\|^{2}_{2}\right]\leq\xi_{1}\left(1-\frac{\tilde{\Delta}_{1}\alpha}{40}\right)^{n}+\frac{20\xi_{2}\alpha\eta^{2}}{\tilde{\Delta}_{1}}+\xi_{2}\alpha.$
  2. (b)
    Let $\alpha_{i}=\frac{\alpha}{i+h}$ for all $i\geq 0$, with $\alpha$ and $h$ chosen so that
    $h\geq\max\left\{1+\frac{\alpha\tilde{\Delta}_{1}}{40},1+28B\left(\frac{20\eta^{2}\alpha}{\tilde{\Delta}_{1}}+\alpha+\frac{20}{\tilde{\Delta}_{1}}\right)\right\},\quad\text{and}\quad\alpha>\frac{40}{\tilde{\Delta}_{1}}.$
    Then, for all $n\geq 0$,
    $\mathbb{E}\!\left[\|\Theta_{n}-\Theta_{\mu}\|^{2}_{2}\right]$ $\leq\xi_{1}\left(\frac{h}{n+h}\right)^{\frac{\alpha\tilde{\Delta}_{1}}{40}}+\frac{5\xi_{2}e^{2}\eta^{2}(20+\tilde{\Delta}_{1})\alpha^{2}}{(n+h)(\alpha\tilde{\Delta}_{1}-40)}+\frac{\xi_{2}\alpha}{n+h}.$
As earlier, since $\mathbb{E}\!\left[(\kappa_{n}-\kappa_{\mu})^{2}\right]\leq\mathbb{E}\!\left[\|\Theta_{n}-\Theta_{\mu}\|^{2}_{2}\right]$, the bound in the Theorem above also hold for MSE in estimating $\kappa_{\mu}$. Theorem 5.1 immediately follows from Theorem 2.1 by recalling that estimating $\Theta_{\mu}$ corresponds to the setup of Section 2, with $r(\cdot,\cdot)$, $Q_{\mu}(\cdot,\cdot)$, $d_{\mu}$ replacing $f(\cdot)$, $V(\cdot)$, and $\pi$. Hence, we omit its proof.
Corollary 5.1.1.
Estimates from Algorithm 3 with step size as in Theorem 5.1 $(b)$ satisfy $\mathbb{E}\!\left[\|Q_{n}-Q_{\mu}\|^{2}_{2}\right]=O(\frac{1}{n})$ MSE bound for policy evaluation problem in the average reward RL setting.
The above Corollary follows since $\mathbb{E}\!\left[\|Q_{n}-Q_{\mu}\|^{2}_{2}\right]\leq\mathbb{E}\!\left[\|\Theta_{n}-\Theta_{\mu}\|^{2}_{2}\right]$. Hence, the bounds in Theorem 5.1 also hold for the MSE in estimating $Q_{\mu}$. Recall that the previously best-known rate for MSE for policy evaluation in average reward RL is $O(\frac{\log n}{n})$ Zhang et al. (2021). A tighter analysis for bounding the MSE of linear SA in presence of Markov noise, that uses Poisson Equations (again) in the proof, leads to this improvement by a multiplicative factor of $\log n$ in Corollary 5.1.1. We refer the reader to Appendix A for a proof using this approach.
Next, to estimate $\kappa_{\mu}$ or $Q_{\mu}$ up to MSE $\epsilon^{2}$ using Algorithm 3 with diminishing step size, requires $n=O(\frac{1}{\epsilon^{2}})O(\frac{\|Q_{\mu}\|^{2}_{2}}{\tilde{\Delta}^{4}_{1}})$. This follows from inverting the above bound at $\epsilon^{2}$, as in Corollary 2.1.1. This bound depends implicitly on $|S||A|$ via $\|Q_{\mu}\|_{2}$ and $\tilde{\Delta}_{1}$, which depends on the mixing properties of the underlying Markov chain. Observe that compared to (Zhang et al., 2021, Corollary 1), here, in addition to improving the log dependence of $\epsilon$ in the sample complexity bound, we improve the dependence on the mixing-time parameters. In particular, the bound in Zhang et al. (2021) has an additional multiplicative term $K$ that relates to the mixing properties.
Remark 5.3.
Unlike in the discounted RL setting, we point out that the operators in the average reward RL framework are not contractive under any norm. In fact, as in Zhang et al. (2021), we establish semi-norm contraction (Lemma 3.3) and convergence in an appropriate subspace (orthogonal to all $1$s vector). This semi-norm contraction is sufficient since the formulation of asymptotic variance in (31) is unaffected by constant shifts in the estimation of $Q_{\mu}$.
Remark 5.4.
When the underlying state and action spaces are large, we can approximate $Q_{\mu}$ by its projection onto a lower dimensional linear subspace spanned by a given set of (fixed) vectors. The bounds obtained above can be extended to this setting using the ideas from Section 3. For completeness, we present the algorithm and performance bounds for estimating $\kappa_{\mu}$ with this approximation architecture in Appendix F.
6 Conclusions and Future Work
We proposed the first recursive estimator for the asymptotic variance of a Markov chain that requires $O(1)$ computation at each step, does not require storing historical samples, and enjoys the optimal $O(\frac{1}{n})$ rate of convergence for MSE. We use Stochastic Approximation to design the proposed estimator. Using Lyapunov drift arguments, and Poisson equation to handle the Markov noise in analysis, we established finite sample bounds on the performance of the proposed estimator. We generalized it in several directions, including covariance matrix estimation for vector-valued functions of a Markov chain and estimating the asymptotic variance of a given stationary policy in an average reward RL framework. Our tighter analysis using Poisson equations to handle the Markov noise improves the MSE bound for linear SA in presence of Markov noise, in general, and for policy evaluation problem in RL, in particular. We established the ${O}(\frac{1}{n})$ rate of convergence of the proposed algorithm in the tabular and linear function approximation settings. We also characterized the approximation error in the latter setting. Below, we discuss a few directions for future research.
  • We discuss the dependence of the sample complexity in terms of the underlying Markov chain parameters such as the size of state space $S$ and mixing time parameter $\Delta_{1}$ in Corollary 2.1.1. Deriving lower bounds on the sample complexity in terms of these problem dependent parameters is an important direction that remains open. Tight lower bounds could potentially serve as a guide to designing estimators with improved dependence on these parameters. This will be useful for arriving at efficient estimators in settings where $S$ is large, for example, in modern RL applications.
  • Our estimator with diminishing step size, for example in Algorithm 1 with $\alpha_{k}=O(\frac{1}{k})$, uses multiple SA updates at each time, each with a step size differing from the other by only a constant ($c_{1}$, $c_{2}$, and $c_{3}$). One of the limitations of this so-called single-time scale approach is the constraints on step sizes $c_{1}$, $c_{2}$, and $c_{3}$ in order to achieve $O(\frac{1}{n})$ MSE bound. This is a well known limitation of single-time scale SA which can be overcome by using a multi-time scale SA. In multi-time scale SA, each SA update uses a different step size (in terms of its dependence on $k$ at time $k$). However, deriving finite sample guarantees for multi-time scale SA is a challenging direction. Designing an estimator for $\kappa(f)$ that uses multiple time scales, with finite sample guarantees on MSE is an interesting open direction for future research.
  • As we saw in Section 3, when the state space is large, one approach to estimating $\kappa(f)$ is to approximate the value function $V$ using a given approximation architecture (linear in our setting), and use this approximate value function to arrive at an estimate for an approximation of $\kappa(f)$, which we call $\kappa^{*}$. In general, this approach and our LFA results can extend easily for nice enough infinite state Markov chains. However, a big challenge is to get an exact estimation for $\kappa(f)$ in these settings. Even in finite but large state space settings, getting an exact estimation for $\kappa(f)$ remains open.
  • On the RL front, an immediate future direction is to design an actor-critic style algorithm for identifying a policy that maximizes the long-run average reward subject to an asymptotic variance constraint with provable finite sample guarantees. Our policy-evaluation algorithm (Algorithm 3 or Algorithm 4) can serve as a building block for designing the critic in these algorithms.
Acknowledgements
This work was partially supported by NSF grants EPCN-2144316 and CPS-2240982. The authors thank the International Centre for Theoretical Sciences (ICTS), TIFR, Bangalore. This work is a result of a collaboration that initaited during the meeting – Data Science: Probabilistic and Optimization Methods (code:ICTS/dspom2023/7) at ICTS. S. Agrawal also thanks Shaan Ul Haque for helpful discussions.
Appendix A Finite Sample Bounds for Linear SA in Presence of Markov Noise
In this section, we prove a mean squared error (MSE) bound for iterates of a linear stochastic approximation (SA) update. We begin by describing the setup. For $d>0$ let $\Gamma_{0}=\vec{0}\in\mathbb{R}^{d}$, and $\alpha_{k}$ denote the step size at time $k$ (to be specified later). Let $\mathbb{S}$ be a finite dimensional subspace of $\mathbb{R}^{d}$ ($\mathbb{S}$ can also be $\mathbb{R}^{d}$), and let $\Gamma_{k}\in\mathbb{R}^{d}$ be generated as
[…]
Recall that for $\mathbb{S}=\mathbb{R}^{d}$, Srikant and Ying [2019] develop a finite sample MSE bound for iterates $\Gamma_{k}$ when A is Hurwitz, and Zhang et al. [2021] extend it for the setting where $\mathbb{S}$ is a proper subspace of $\mathbb{R}^{d}$. In this section, we improve on both these results by a multiplicative factor of $\log k$ by using Poisson equation for handling the error terms in our analysis. Use of Poisson equations for analyzing SA algorithms has previously appeared in Benveniste et al. [2012]. More recently, Kaledin et al. [2020] and Haque et al. [2023] use Poisson equation to analyse a two-time scale algorithm in presence of Markov noise.
Theorem A.1.
Consider iterates ${\Gamma}_{k}$ generated by the update in (A) with $\Gamma_{0}=\vec{0}$. Suppose there exists $\gamma_{2}>0$ such that the average matrix A satisfies
$\min\limits_{\Gamma\in\mathbb{S}:\|\Gamma\|_{2}=1}~{}-\Gamma^{T}\textsf{A}~{}\Gamma\geq\gamma_{2}>0.$
[…]
  1. (a)
    Let $\alpha_{i}=\alpha$ for all $i$, such that
    $\alpha<\min\left\{\frac{2}{\gamma_{2}},\frac{1}{28B\left(1+\frac{\textsf{H \!}^{2}}{\gamma_{2}}\right)}\right\}.$
    Then, for all $k\geq 1$,
    $\mathbb{E}\!\left[\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}\right]\leq\psi_{1}\left(1-\frac{\gamma_{2}\alpha}{2}\right)^{k}+\frac{\psi_{2}\alpha\textsf{H \!}^{2}}{\gamma_{2}}+\psi_{2}\alpha.$
  2. (b)
    Let $\alpha_{i}=\frac{\alpha}{i+h}$ for all $i\geq 0$, with $\alpha$ and $h$ chosen so that
    $h\geq\max\left\{1+\frac{\alpha\gamma_{2}}{2},1+28B\left(\frac{\textsf{H \!}^{2}\alpha}{\gamma_{2}}+\alpha+\frac{1}{\gamma_{2}}\right)\right\},\quad\text{and}\quad\alpha>\frac{2}{\gamma_{2}}.$
    Then, for all $k\geq 0$,
    $\mathbb{E}\!\left[\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq\psi_{1}\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}+\frac{5\psi_{2}e^{2}\textsf{H \!}^{2}(1+\gamma_{2})\alpha^{2}}{(k+h)(\alpha\gamma_{2}-2)}+\frac{\psi_{2}\alpha}{k+h}.$
Proof.
[…]
$\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}-\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$ $=\|\Gamma_{k+1}-\Gamma_{k}+\Gamma_{k}-\Gamma^{*}\|^{2}_{2}-\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$
$=\|\Gamma_{k+1}-\Gamma_{k}\|^{2}_{2}+2(\Gamma_{k+1}-\Gamma_{k})^{T}(\Gamma_{k}-\Gamma^{*})$
[…]
$\qquad\qquad\qquad+2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}A(\Gamma_{k}-\Gamma^{*})$
[…]
$\qquad\qquad\qquad-\gamma_{2}\alpha_{k}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}.$
[…]
Then,
[…]
Using this,
$\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}-\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$ $\leq\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}(-\gamma_{2}\alpha_{k}+2\alpha^{2}_{k}\textsf{H \!}^{2})+2\alpha^{2}_{k}\textsf{H \!}^{2}(1+\|\Gamma^{*}\|_{2})^{2}$
[…]
Furthermore, for each $s\in{\mathcal{S}}$, $V_{x}(s):=V_{1}(s)x-V_{2}(s)$ is a vector that solves
[…]
Moreover, there exists a constant $B:=\max\left\{\max_{s}\|V_{1}(s)\|_{2},\max_{s}\|V_{2}(s)\|_{2}\right\}$ such that for each $s$ and $x$,
$\|V_{x}(s)\|_{2}\leq B(\|x\|_{2}+1),$
and $V_{x}(s)$ is a Lipschitz function of $x$, i.e.,
$\|V_{x}(s)-V_{y}(s)\|_{2}=\|V_{1}(s)(x-y)\|_{2}\leq\|V_{1}(s)\|_{2}\|x-y\|_{2}\leq B\|x-y\|_{2}.$
Using these in the cross-term in (40),
[…]
$=2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\left(V_{\Gamma_{k}}(Y_{k})-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k+1})|Y_{k}\right]\right)$
$=2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\big{(}V_{\Gamma_{k}}(Y_{k})-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right]$
$\qquad+\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right]-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k+1})|Y_{k}\right]\big{)}$
$=2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\left(V_{\Gamma_{k}}(Y_{k})-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right]\right)$
$\qquad+2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right]$
$\qquad-2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k+1})|Y_{k}\right].$
Substituting back in (40),
$\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}-\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$ $\leq\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}(-\gamma_{2}\alpha_{k}+2\alpha^{2}_{k}\textsf{H \!}^{2})+2\alpha^{2}_{k}\textsf{H \!}^{2}(1+\|\Gamma^{*}\|_{2})^{2}$
$\quad+\underbrace{2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\left(V_{\Gamma_{k}}(Y_{k})-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right]\right)}_{0-\text{mean term}}$
$\quad+\underbrace{2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\left(\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right]-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k+1})|Y_{k}\right]\right)}_{T_{1}}.$ (41)
Next, we show that $T_{1}$ is a small order. To this end, let
$d_{k}:=(\Gamma_{k}-\Gamma^{*})^{T}\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right].$
Then,
$T_{1}$ $=2\alpha_{k}d_{k}-2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k+1})|Y_{k}\right]-2\alpha_{k}d_{k+1}+2\alpha_{k}d_{k+1}$
$=2\alpha_{k}(d_{k}-d_{k+1})+2\alpha_{k}(\Gamma_{k+1}-\Gamma^{*})^{T}\mathbb{E}\!\left[V_{\Gamma_{k+1}}(Y_{k+1})|Y_{k}\right]$
$\qquad\qquad-2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k+1})|Y_{k}\right]$
$=2\alpha_{k}(d_{k}-d_{k+1})+2\alpha_{k}(\Gamma_{k+1}-\Gamma^{*})^{T}\left(\mathbb{E}\!\left[V_{\Gamma_{k+1}}(Y_{k+1})|Y_{k}\right]-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k+1})|Y_{k}\right]\right)$
$\qquad\qquad+2\alpha_{k}(\Gamma_{k+1}-\Gamma_{k})^{T}\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k+1})|Y_{k}\right]$
$\leq 2\alpha_{k}(d_{k}-d_{k+1})+2\alpha_{k}\|\Gamma_{k+1}-\Gamma^{*}\|_{2}\mathbb{E}\!\left[\|V_{\Gamma_{k+1}}(Y_{k+1})-V_{\Gamma_{k}}(Y_{k+1})\|_{2}|Y_{k}\right]$ (Cauchy-Schwarz)
$\qquad\qquad+2\alpha_{k}\|\Gamma_{k+1}-\Gamma_{k}\|_{2}\mathbb{E}\!\left[\|V_{\Gamma_{k}}(Y_{k+1})\|_{2}|Y_{k}\right]$
$\leq 2\alpha_{k}(d_{k}-d_{k+1})+2B\alpha_{k}\|\Gamma_{k+1}-\Gamma^{*}\|_{2}\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma_{k}\|_{2}|Y_{k}\right]$ (Lipschitzness)
$\qquad\qquad+2B\alpha_{k}\|\Gamma_{k+1}-\Gamma_{k}\|_{2}\mathbb{E}\!\left[(\|{\Gamma_{k}}\|_{2}+1)|Y_{k}\right]$
[…]
Next, observe that
[…]
Using this, as well as the inequality
$2ab\leq a^{2}+b^{2}$
to further bound the cross terms in $T_{1}$, we get
$T_{1}$ $\leq 2\alpha_{k}(d_{k}-d_{k+1})+4B\textsf{H \!}^{2}\alpha^{3}_{k}(1+\|\Gamma^{*}\|_{2})^{2}+4B\textsf{H \!}^{2}\alpha^{3}_{k}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$
$\qquad+2B\alpha^{2}_{k}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}+4B\alpha^{2}_{k}\textsf{H \!}^{2}(1+\|\Gamma^{*}\|_{2})^{2}+4B\alpha^{2}_{k}\textsf{H \!}^{2}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$
$\qquad+B\alpha^{2}_{k}(1+\|\Gamma^{*}\|_{2})^{2}+2B\alpha^{2}_{k}\textsf{H \!}^{2}(1+\|\Gamma^{*}\|_{2})^{2}+2B\alpha^{2}_{k}\textsf{H \!}^{2}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$
$=2\alpha_{k}(d_{k}-d_{k+1})+4B\textsf{H \!}^{2}\alpha^{3}_{k}(1+\|\Gamma^{*}\|_{2})^{2}+4B\textsf{H \!}^{2}\alpha^{3}_{k}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$
$\qquad+2B\alpha^{2}_{k}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}+6B\alpha^{2}_{k}\textsf{H \!}^{2}(1+\|\Gamma^{*}\|_{2})^{2}+6B\alpha^{2}_{k}\textsf{H \!}^{2}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$
$\qquad+B\alpha^{2}_{k}(1+\|\Gamma^{*}\|_{2})^{2}$
$\leq 2\alpha_{k}(d_{k}-d_{k+1})+4B\textsf{H \!}^{2}\alpha^{3}_{k}(1+\|\Gamma^{*}\|_{2})^{2}+4B\textsf{H \!}^{2}\alpha^{3}_{k}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$
[…]
$\leq 2\alpha_{k}(d_{k}-d_{k+1})+12B\textsf{H \!}^{2}\alpha^{2}_{k}(1+\|\Gamma^{*}\|_{2})^{2}+12B\textsf{H \!}^{2}\alpha^{2}_{k}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}.$ (Since $\alpha^{3}_{k}\leq\alpha^{2}_{k}$)
Substituting this bound on $T_{1}$ back in (41), we have
$\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}-\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$
$\leq\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}(-\gamma_{2}\alpha_{k}+2\alpha^{2}_{k}\textsf{H \!}^{2})+2\alpha^{2}_{k}\textsf{H \!}^{2}(1+\|\Gamma^{*}\|_{2})^{2}$
$\qquad+2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\left(V_{\Gamma_{k}}(Y_{k})-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right]\right)$
$\qquad+\underbrace{2\alpha_{k}(d_{k}-d_{k+1})}_{T_{2}}+12B\alpha^{2}_{k}\textsf{H \!}^{2}\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}+12B\alpha^{2}_{k}\textsf{H \!}^{2}(1+\|\Gamma^{*}\|_{2})^{2}$ (Since $B\geq 1$)
$\leq\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}(-\gamma_{2}\alpha_{k}+14B\alpha^{2}_{k}\textsf{H \!}^{2})+14B\alpha^{2}_{k}\textsf{H \!}^{2}(1+\|\Gamma^{*}\|_{2})^{2}$
$\qquad+2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}\left(V_{\Gamma_{k}}(Y_{k})-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right]\right)$
$\qquad+\underbrace{2\alpha_{k}(d_{k}-d_{k+1})}_{T_{2}}.$ (42)
Now, observe that
$|d_{k}|$ $=\left|{(\Gamma_{k}-\Gamma^{*})^{T}\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right]}\right|$
$\leq\|\Gamma_{k}-\Gamma^{*}\|_{2}\mathbb{E}\!\left[\|V_{\Gamma_{k}}(Y_{k})\|_{2}|Y_{k-1}\right]$ (Cauchy-Schwarz and Jensen’s)
$\leq B\|\Gamma_{k}-\Gamma^{*}\|_{2}(\|\Gamma_{k}-\Gamma^{*}\|_{2}+1+\|\Gamma^{*}\|_{2})$ (Lipschitzness)
$=B\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}+B\|\Gamma_{k}-\Gamma^{*}\|_{2}(1+\|\Gamma^{*}\|_{2})$ (43)
$\leq 2B\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}+2B(1+\|\Gamma^{*}\|_{2})^{2},$ (44)
where, for the last inequality, we used $2ab\leq a^{2}+b^{2}$.
Let us now get an upper bound on $T_{2}$ involving terms $2\alpha_{k-1}d_{k}-2\alpha_{k}d_{k+1}$, instead.
$T_{2}$ $=2\alpha_{k}(d_{k}-d_{k+1})$
$=2\alpha_{k-1}d_{k}-2\alpha_{k}d_{k+1}+2(\alpha_{k}-\alpha_{k-1})d_{k}$
$\leq 2\alpha_{k-1}d_{k}-2\alpha_{k}d_{k+1}+2|\alpha_{k}-\alpha_{k-1}||d_{k}|$
$\leq 2\alpha_{k-1}d_{k}-2\alpha_{k}d_{k+1}+\frac{6\alpha^{2}_{k}}{\alpha}|d_{k}|$ (From Lemma G.1)
$=2\alpha_{k-1}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)d_{k}-2\alpha_{k}d_{k+1}+\gamma_{2}\alpha_{k}\alpha_{k-1}d_{k}+\frac{6\alpha^{2}_{k}}{\alpha}|d_{k}|$
$\leq 2\alpha_{k-1}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)d_{k}-2\alpha_{k}d_{k+1}+2\gamma_{2}\alpha^{2}_{k}|d_{k}|+\frac{6\alpha^{2}_{k}}{\alpha}|d_{k}|$ (From Lemma G.2)
$\leq 2\alpha_{k-1}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)d_{k}-2\alpha_{k}d_{k+1}$
$\qquad\qquad+4B\alpha^{2}_{k}\left(\gamma_{2}+\frac{3}{\alpha}\right)\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}+4B\alpha^{2}_{k}\left(\gamma_{2}+\frac{3}{\alpha}\right)(1+\|\Gamma^{*}\|_{2})^{2}.$ (Using (44))
Remark A.1.
Note that a more careful bound for constant step size setting would ensure that the term involving $\frac{3}{\alpha}$ is absent. This corresponds to the bound from Lemma G.1, which is 00 in that setting.
Substituting this back in (42), and noting that $2\alpha_{k}\leq 2$, we have
$\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}-\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}$ $\leq\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}\left(-\gamma_{2}\alpha_{k}+\alpha^{2}_{k}\left(14B\textsf{H \!}^{2}+4B\left(\gamma_{2}+\frac{3}{\alpha}\right)\right)\right)$
$+\alpha^{2}_{k}(1+\|\Gamma^{*}\|_{2})^{2}\left(14\textsf{H \!}^{2}B+4B\left(\gamma_{2}+\frac{3}{\alpha}\right)\right)$
$+2\alpha_{k}(\Gamma_{k}-\Gamma^{*})^{T}(V_{\Gamma_{k}}(Y_{k})-\mathbb{E}\!\left[V_{\Gamma_{k}}(Y_{k})|Y_{k-1}\right])$
$+2\alpha_{k-1}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)d_{k}-2\alpha_{k}d_{k+1}.$
Taking conditional expectation, conditioned on randomness till time $k$,
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}|\Gamma_{k},Y_{k-1}\right]$ $\leq\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}\left(1-\gamma_{2}\alpha_{k}+\alpha^{2}_{k}(14B\textsf{H \!}^{2}+4B\left(\gamma_{2}+\frac{3}{\alpha}\right))\right)$
$+\alpha^{2}_{k}(1+\|\Gamma^{*}\|_{2})^{2}\left(14\textsf{H \!}^{2}B+4B\left(\gamma_{2}+\frac{3}{\alpha}\right)\right)$
$+2\alpha_{k-1}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)d_{k}-2\alpha_{k}d_{k+1},$
which further gives
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq\mathbb{E}\!\left[\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}\right]\left(1-\gamma_{2}\alpha_{k}+\alpha^{2}_{k}\left(14B\textsf{H \!}^{2}+4B\left(\gamma_{2}+\frac{3}{\alpha}\right)\right)\right)$
$\qquad+\alpha^{2}_{k}(1+\|\Gamma^{*}\|_{2})^{2}\left(14\textsf{H \!}^{2}B+4B\left(\gamma_{2}+\frac{3}{\alpha}\right)\right)$
$\qquad+2\alpha_{k-1}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)\mathbb{E}\!\left[d_{k}\right]-2\alpha_{k}\mathbb{E}\!\left[d_{k+1}\right]$
$\leq\mathbb{E}\!\left[\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}\right]\left(1-\gamma_{2}\alpha_{k}+14B\alpha^{2}_{k}\left(\textsf{H \!}^{2}+\gamma_{2}+\frac{3}{\alpha}\right)\right)$
$\qquad+14B\alpha^{2}_{k}(1+\|\Gamma^{*}\|_{2})^{2}\left(\textsf{H\!}^{2}+\gamma_{2}+\frac{3}{\alpha}\right)$
$\qquad+2\alpha_{k-1}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)\mathbb{E}\!\left[d_{k}\right]-2\alpha_{k}\mathbb{E}\!\left[d_{k+1}\right].$
Choose $\alpha_{k}$ small enough so that
$14B\alpha_{k}\left(\textsf{H \!}^{2}+\gamma_{2}+\frac{3}{\alpha}\right)\leq\frac{\gamma_{2}}{2},$
so that
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq\mathbb{E}\!\left[\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}\right]\left(1-\frac{\gamma_{2}\alpha_{k}}{2}\right)$
$\qquad+2\alpha_{k-1}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)\mathbb{E}\!\left[d_{k}\right]-2\alpha_{k}\mathbb{E}\!\left[d_{k+1}\right]$
$\qquad+14B\alpha^{2}_{k}(1+\|\Gamma^{*}\|_{2})^{2}\left(\textsf{H\!}^{2}+\gamma_{2}+\frac{3}{\alpha}\right).$
Let
$B_{1}:=14B(1+\|\Gamma^{*}\|_{2})^{2}\left(\textsf{H \!}^{2}+\gamma_{2}+\frac{3}{\alpha}\right).$
Then, the above recursion becomes
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq\mathbb{E}\!\left[\|\Gamma_{k}-\Gamma^{*}\|^{2}_{2}\right]\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)$
$\qquad\qquad+2\alpha_{k-1}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)\mathbb{E}\!\left[d_{k}\right]-2\alpha_{k}\mathbb{E}\!\left[d_{k+1}\right]+\alpha^{2}_{k}B_{1}.$
On opening the recursion above, we get
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)\left(1-\frac{\alpha_{k-1}\gamma_{2}}{2}\right)\mathbb{E}\!\left[\|\Gamma_{k-1}-\Gamma^{*}\|^{2}_{2}\right]$
$\qquad+B_{1}\left(\alpha^{2}_{k}+\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)\alpha^{2}_{k-1}\right)$
$\qquad+2\alpha_{k-2}\left(1-\frac{\alpha_{k}\gamma_{2}}{2}\right)\left(1-\frac{\alpha_{k-1}\gamma_{2}}{2}\right)\mathbb{E}\!\left[d_{k-1}\right]-2\alpha_{k}\mathbb{E}\!\left[d_{k+1}\right]$
$\leq\mathbb{E}\!\left[\|\Gamma_{0}-\Gamma^{*}\|^{2}_{2}\right]\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)+B_{1}\sum\limits_{i=0}^{k}\alpha^{2}_{i}\prod\limits_{j=i+1}^{k}\left(1-\frac{\alpha_{j}\gamma_{2}}{2}\right)$
$\qquad+2\alpha_{-1}\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)\mathbb{E}\!\left[d_{0}\right]-2\alpha_{k}\mathbb{E}\!\left[d_{k+1}\right].$
Next, upper-bounding the last term,
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq\mathbb{E}\!\left[\|\Gamma_{0}-\Gamma^{*}\|^{2}_{2}\right]\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)+B_{1}\sum\limits_{i=0}^{k}\alpha^{2}_{i}\prod\limits_{j=i+1}^{k}\left(1-\frac{\alpha_{j}\gamma_{2}}{2}\right)$
$\qquad+2\alpha_{-1}\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)\mathbb{E}\!\left[d_{0}\right]+2\alpha_{k}\mathbb{E}\!\left[|d_{k+1}|\right]$
$\leq\mathbb{E}\!\left[\|\Gamma_{0}-\Gamma^{*}\|^{2}_{2}\right]\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)+B_{1}\sum\limits_{i=0}^{k}\alpha^{2}_{i}\prod\limits_{j=i+1}^{k}\left(1-\frac{\alpha_{j}\gamma_{2}}{2}\right)$ (From (44))
$~{}~{}~{}+2\alpha_{-1}\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)\mathbb{E}\!\left[d_{0}\right]+4\alpha_{k}B\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$
$\qquad+4\alpha_{k}B(1+\|\Gamma^{*}\|_{2})^{2}.$
On rearranging, we get
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\left(1-4\alpha_{k}B\right)$
$\leq\mathbb{E}\!\left[\|\Gamma_{0}-\Gamma^{*}\|^{2}_{2}\right]\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)+B_{1}\sum\limits_{i=0}^{k}\alpha^{2}_{i}\prod\limits_{j=i+1}^{k}\left(1-\frac{\alpha_{j}\gamma_{2}}{2}\right)$
$\qquad+2\alpha_{-1}\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)\mathbb{E}\!\left[d_{0}\right]+4B\alpha_{k}(1+\|\Gamma^{*}\|_{2})^{2}$
$\leq\mathbb{E}\!\left[\|\Gamma_{0}-\Gamma^{*}\|^{2}_{2}\right]\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)+B_{1}\sum\limits_{i=0}^{k}\alpha^{2}_{i}\prod\limits_{j=i+1}^{k}\left(1-\frac{\alpha_{j}\gamma_{2}}{2}\right)$
$\qquad+4\alpha_{-1}B(1+\|\Gamma^{*}\|_{2})^{2}\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)+4B\alpha_{k}(1+\|\Gamma^{*}\|_{2})^{2}.$ (Using (43))
Choosing $\alpha_{k}$ small enough so that
$1-4\alpha_{k}B\geq\frac{1}{2},$
we get
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq 2\mathbb{E}\!\left[\|\Gamma_{0}-\Gamma^{*}\|^{2}_{2}\right]\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)+2B_{1}\sum\limits_{j=0}^{k}\alpha^{2}_{j}\prod\limits_{i=j+1}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)$
$\qquad\qquad+8\alpha_{-1}B(1+\|\Gamma^{*}\|_{2})^{2}\prod\limits_{i=0}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)+8B\alpha_{k}(1+\|\Gamma^{*}\|_{2})^{2}.$ (45)
We now make the above bound more explicit in two cases: diminishing step size and constant step size.
Diminishing Step Size
In this subsection, we consider step sizes of the form $\alpha_{k}=\frac{\alpha}{k+h}$, for appropriate choices for $\alpha$ and $h\geq 2$. Let us bound each term in the recursion in (45) separately. First observe that for any $i_{0}\geq 0$,
$\prod\limits_{j=i_{0}}^{k}\left(1-\frac{\gamma_{2}\alpha_{j}}{2}\right)$ $=\prod\limits_{j=i_{0}}^{k}\left(1-\frac{\gamma_{2}\alpha}{2(j+h)}\right)$
$\leq\prod\limits_{j=i_{0}}^{k}e^{-\frac{\gamma_{2}\alpha}{2(j+h)}}$
$=e^{-\frac{\gamma_{2}\alpha}{2}\sum\limits_{j=i_{0}}^{k}\frac{1}{j+h}}$
$\leq e^{-\frac{\gamma_{2}\alpha}{2}\ln\frac{k+h}{i_{0}+h}}$
$=\left(\frac{i_{0}+h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}.$ (46)
Using this in (45) with $i_{0}=0$, we get
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq 2\mathbb{E}\!\left[\|\Gamma_{0}-\Gamma^{*}\|^{2}_{2}\right]\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}+\underbrace{2B_{1}\sum\limits_{j=0}^{k}\alpha^{2}_{j}\prod\limits_{i=j+1}^{k}\left(1-\frac{\alpha_{i}\gamma_{2}}{2}\right)}_{:=T^{(d)}_{1}}$
$\qquad\qquad+8\alpha_{-1}B(1+\|\Gamma^{*}\|_{2})^{2}\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}+8B\alpha_{k}(1+\|\Gamma^{*}\|_{2})^{2}.$
Let us now bound $T^{(d)}_{1}$.
$T^{(d)}_{1}$ $=2B_{1}\sum\limits_{j=0}^{k}\prod_{i=j+1}^{k}\left(1-\frac{\gamma_{2}\alpha_{i}}{2}\right)\alpha^{2}_{j}$
$\leq 2B_{1}\sum\limits_{j=0}^{k}\left(\frac{j+1+h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}\frac{\alpha^{2}}{(j+h)^{2}}$ (From (46))
$=\frac{2B_{1}\alpha^{2}}{(k+h)^{\frac{\alpha\gamma_{2}}{2}}}\sum\limits_{j=0}^{k}\left(\frac{j+1+h}{j+h}\right)^{2}\left(j+1+h\right)^{\frac{\alpha\gamma_{2}}{2}-2}$
$\leq\frac{8B_{1}\alpha^{2}}{(k+h)^{\frac{\alpha\gamma_{2}}{2}}}\sum\limits_{j=0}^{k}\left(j+1+h\right)^{\frac{\alpha\gamma_{2}}{2}-2}.$ (Since $h\geq 2$)
Choose $\alpha$ large enough so that
$\frac{\alpha\gamma_{2}}{2}>1,$ (Condition 3)
and hence,
[…]
$=\frac{1}{\frac{\gamma_{2}\alpha}{2}-1}\left((k+2+h)^{\frac{\gamma_{2}\alpha}{2}-1}-(1+h)^{\frac{\gamma_{2}\alpha}{2}-1}\right)$
$\leq\frac{2(k+2+h)^{\frac{\gamma_{2}\alpha}{2}-1}}{\gamma_{2}\alpha-2}.$
Using this,
$T^{(d)}_{1}$ $\leq\left(\frac{8B_{1}\alpha^{2}}{(k+h)^{\frac{\alpha\gamma_{2}}{2}}}\right)\frac{2(k+2+h)^{\frac{\gamma_{2}\alpha}{2}-1}}{\gamma_{2}\alpha-2}$
$=\frac{16B_{1}\alpha^{2}}{\gamma_{2}\alpha-2}\left(\frac{1}{k+2+h}\right)\left(\frac{k+2+h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}$
$\leq\frac{16B_{1}\alpha^{2}}{\gamma_{2}\alpha-2}\left(\frac{e^{\frac{\alpha\gamma_{2}}{k+h}}}{k+2+h}\right)$ ($1+x\leq e^{x}$)
$=\frac{16B_{1}\alpha^{2}}{\gamma_{2}\alpha-2}\left(\frac{e^{\alpha_{k}\gamma_{2}}}{k+2+h}\right).$
Choosing $h$ large enough to ensure
$\alpha_{k}\gamma_{2}\leq 2,$
we have
$T^{(d)}_{1}\leq\frac{16B_{1}\alpha^{2}e^{2}}{(\gamma_{2}\alpha-2)(k+h+1)}.$
Using this bound, the MSE bound becomes:
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq 2\mathbb{E}\!\left[\|\Gamma_{0}-\Gamma^{*}\|^{2}_{2}\right]\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}+\frac{16B_{1}\alpha^{2}e^{2}}{(\gamma_{2}\alpha-2)(k+h+1)}$
$\qquad\qquad+8\alpha_{-1}B(1+\|\Gamma^{*}\|_{2})^{2}\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}+8B\alpha_{k}(1+\|\Gamma^{*}\|_{2})^{2}$
$\leq 2(1+\|\Gamma^{*}\|_{2})^{2}\left(1+4\alpha_{-1}B\right)\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}$
$\qquad\qquad+\frac{16B_{1}\alpha^{2}e^{2}}{(\gamma_{2}\alpha-2)(k+h)}+8B\alpha_{k}(1+\|\Gamma^{*}\|_{2})^{2}$
$=2(1+\|\Gamma^{*}\|_{2})^{2}\left(1+4\alpha_{-1}B\right)\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}$ (47)
$\qquad\qquad+\frac{2\alpha}{k+h}\left(\frac{8B_{1}\alpha e^{2}}{(\gamma_{2}\alpha-2)}+4B(1+\|\Gamma^{*}\|_{2})^{2}\right).$
Conditions on step size parameters. Recall that $h\geq 2$. In addition,  (A) is satisfied if
$h\geq 1+28B\left(\frac{\textsf{H \!}^{2}\alpha}{\gamma_{2}}+\alpha+\frac{3}{\gamma_{2}}\right),$
which also automatically satisfies (A). (A) is satisfied for
$h\geq 1+\frac{\alpha\gamma_{2}}{2}.$
Moreover,  (Condition 3) requires that
$\alpha>\frac{2}{\gamma_{2}}.$
Final conditions. Thus,
$h\geq\max\left\{2,1+\frac{\alpha\gamma_{2}}{2},1+28B\left(\frac{\textsf{H \!}^{2}\alpha}{\gamma_{2}}+\alpha+\frac{1}{\gamma_{2}}\right)\right\},\quad\text{and}\quad\alpha>\frac{2}{\gamma_{2}}.$
Recall from A,
$B_{1}=14B(1+\|\Gamma^{*}\|_{2})^{2}\left(\textsf{H \!}^{2}+\gamma_{2}+\frac{3}{\alpha}\right)$
Final bound. Using this in (47), we get
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq 2(1+\|\Gamma^{*}\|_{2})^{2}(1+4\alpha_{-1}B)\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}$
$\qquad+\frac{8B\alpha(1+\|\Gamma^{*}\|_{2})^{2}}{k+h}\left(\frac{28\alpha e^{2}}{\gamma_{2}\alpha-2}\left(\textsf{H \!}^{2}+\gamma_{2}+\frac{3}{\alpha}\right)+1\right)$
$\leq 2(1+\|\Gamma^{*}\|_{2})^{2}(1+4\alpha_{-1}B)\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}$
$\qquad+\frac{112B\alpha(1+\|\Gamma^{*}\|_{2})^{2}}{k+h}\left(\frac{2\alpha e^{2}}{\gamma_{2}\alpha-2}\left(\textsf{H \!}^{2}+\gamma_{2}+\frac{3}{\alpha}\right)+1\right).$
Further, from (A), we have for all $k$,
$\alpha_{-1}\leq\frac{1}{8B}.$
Hence, the bound becomes
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq 3(1+\|\Gamma^{*}\|_{2})^{2}\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}$
$\qquad+\frac{112B\alpha(1+\|\Gamma^{*}\|_{2})^{2}}{k+h}\left(\frac{2\alpha e^{2}}{\gamma_{2}\alpha-2}\left(\textsf{H \!}^{2}+\gamma_{2}+\frac{3}{\alpha}\right)+1\right).$
Now, finally, since $\alpha>\frac{2}{\gamma_{2}}$, we have $\frac{1}{\alpha}<\frac{\gamma_{2}}{2}$. Using this in the last term above,
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq 3(1+\|\Gamma^{*}\|_{2})^{2}\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}$
$\qquad+\frac{112B\alpha(1+\|\Gamma^{*}\|_{2})^{2}}{k+h}\left(\frac{2\alpha e^{2}}{\gamma_{2}\alpha-2}\left(\textsf{H \!}^{2}+\frac{5\gamma_{2}}{2}\right)+1\right)$
$\leq 3(1+\|\Gamma^{*}\|_{2})^{2}\left(\frac{h}{k+h}\right)^{\frac{\alpha\gamma_{2}}{2}}$
$\qquad+\frac{112B\alpha(1+\|\Gamma^{*}\|_{2})^{2}}{k+h}\left(\frac{5\alpha e^{2}\textsf{H \!}^{2}}{\gamma_{2}\alpha-2}\left(1+\gamma_{2}\right)+1\right)$
Constant Step Size
In this section, we bound the recursion in (45) for the setting of constant step size, i.e., $\alpha_{k}=\alpha$. In this setting, the recursion in (45) becomes
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq 2\|\Gamma^{*}\|^{2}_{2}\left(1-\frac{\alpha\gamma_{2}}{2}\right)^{k+1}+2B_{1}\alpha^{2}\sum\limits_{j=0}^{k}\left(1-\frac{\alpha\gamma_{2}}{2}\right)^{j}$
$\qquad\qquad+8\alpha_{-1}B(1+\|\Gamma^{*}\|_{2})^{2}\left(1-\frac{\alpha\gamma_{2}}{2}\right)^{k+1}+8B\alpha(1+\|\Gamma^{*}\|_{2})^{2}$
$=2(1+\|\Gamma^{*}\|_{2})^{2}\left(1+4\alpha_{-1}B\right)\left(1-\frac{\alpha\gamma_{2}}{2}\right)^{k+1}$ (49)
$\qquad\qquad+4\alpha\left(\frac{B_{1}}{\gamma_{2}}+2B(1+\|\Gamma^{*}\|_{2})^{2}\right).$
Conditions on step size. Using the observation in Remark A.1, and choosing
$\alpha\leq\frac{1}{28B\left(1+\frac{\textsf{H \!}^{2}}{\gamma_{2}}\right)},$
ensures (A), which also automatically satisfies (A). Moreover, $\alpha$ also needs to satisfy
$\alpha<\frac{2}{\gamma_{2}}.$
Final conditions. Thus,
$\alpha<\min\left\{\frac{2}{\gamma_{2}},\frac{1}{28B\left(1+\frac{\textsf{H \!}^{2}}{\gamma_{2}}\right)}\right\}.$
Under these conditions, and accounting for the observation from Remark A.1, recall from (A) that
$B_{1}:=14B\gamma_{2}(1+\|\Gamma^{*}\|_{2})^{2}\left(1+\frac{\textsf{H \!}^{2}}{\gamma_{2}}\right).$
Final bound. Substituting this in (49), along with (A), which ensures that $\alpha\leq\frac{1}{8B}$, we have
$\mathbb{E}\!\left[\|\Gamma_{k+1}-\Gamma^{*}\|^{2}_{2}\right]$ $\leq 2(1+\|\Gamma^{*}\|_{2})^{2}\left(1+4\alpha B\right)\left(1-\frac{\alpha\gamma_{2}}{2}\right)^{k+1}$
$\qquad\qquad+56B\alpha(1+\|\Gamma^{*}\|_{2})^{2}\left(\frac{\textsf{H \!}^{2}}{\gamma_{2}}+2\right)$
$\leq 3(1+\|\Gamma^{*}\|_{2})^{2}\left(1-\frac{\alpha\gamma_{2}}{2}\right)^{k+1}+112B\alpha(1+\|\Gamma^{*}\|_{2})^{2}\left(\frac{\textsf{H \!}^{2}}{\gamma_{2}}+1\right).$
Appendix B Proofs for Results in Section 1.3
B.1 Proof of Proposition 1.1
From Lemma B.1, we get that the asymptotic variance $\kappa(f)$ is constant independent of the starting state or starting distribution, and is given by the expression in (1.1). Further, from Lemma B.2, we get the first two formulations for $\kappa(f)$ (Equations (3) and (4)). Finally, the third formulation follows by replacing $(f(X)-\bar{f})$ terms in (4) by $V(X)-PV(X)$, which follows from the Poisson equation (1.3).∎
B.2 Variance of Functions of a Discrete Time Markov Chain (DTMC)
Recall that $X=(X_{k}:k\geq 0)$ is an irreducible and aperiodic DTMC on a finite-state space ${\mathcal{S}}$ with transition probability matrix $P$ and a unique stationary distribution $\pi$. Let $f:{\mathcal{S}}\rightarrow\mathbb{R}$ be any function, and let $\bar{f}:=\sum_{x\in{\mathcal{S}}}\pi(x)f(x)$ denote its stationary expectation. Let $V$ denote the solution to the Poisson equation for $f-\bar{f}$, i.e., for each $x\in{\mathcal{S}}$, $V$ satisfies
$V(x)-\sum\nolimits_{x^{\prime}\in{\mathcal{S}}}P(x,x^{\prime})V(x^{\prime})=f(x)-\bar{f}.$
Lemma B.1.
Consider the Markov Chain starting in distribution $\pi_{0}$. The asymptotic variance of $f$, $\kappa(f)$, is a constant independent of the starting distribution of the Markov Chain $\pi_{0}$, i.e.,
$\kappa(f):=\lim\limits_{n\rightarrow\infty}\frac{1}{n}~{}\operatorname{Var}\left[\sum\limits_{k=0}^{n-1}f(X_{k})|X_{0}\sim\pi_{0}\right]=\lim\limits_{n\rightarrow\infty}\frac{1}{n}~{}\operatorname{Var}\left[\sum\limits_{k=0}^{n-1}f(X_{k})|X_{0}\sim\pi\right].$
Proof.
For $k\in\mathbb{N}_{+}$, let $\bar{f}_{k}:=\mathbb{E}_{\pi_{0}\!}\!\left[f(X_{k})\right]=\mathbb{E}\!\left[f(X_{k})|X_{0}\sim\pi_{0}\right]$. Then,
$\kappa(f)$ $=\frac{1}{n}~{}\mathbb{E}_{\pi_{0}\!}\!\left[\left(\sum\limits_{k=0}^{n-1}\left(f(X_{k})-\bar{f}_{k}\right)\right)^{2}\right]$
$=\frac{1}{n}~{}\sum\limits_{k=0}^{n-1}\mathbb{E}_{\pi_{0}\!}\!\left[(f(X_{k})-\bar{f}_{k})^{2}\right]+\frac{2}{n}\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}\mathbb{E}_{\pi_{0}\!}\!\left[\left(f(X_{i})-\bar{f}_{i}\right)\left(f(X_{k})-\bar{f}_{k}\right)\right].$ (50)
Clearly,
$\lim\limits_{n\rightarrow\infty}\frac{1}{n}~{}\sum\limits_{k=0}^{n-1}\mathbb{E}_{\pi_{0}\!}\!\left[(f(X_{k})-\bar{f}_{k})^{2}\right]=\lim\limits_{n\rightarrow\infty}\frac{1}{n}~{}\sum\limits_{t=0}^{n-1}\mathbb{E}_{\pi\!}\!\left[(f(X_{t})-\bar{f})^{2}\right].$
This follows from the following:
$\lim\limits_{n\rightarrow\infty}\frac{1}{n}~{}\sum\limits_{k=0}^{n-1}\mathbb{E}_{\pi_{0}\!}\!\left[(f(X_{k})-\bar{f}_{k})^{2}\right]$
$=\lim\limits_{n\rightarrow\infty}~{}\frac{1}{n}~{}\sum\limits_{k=0}^{n-1}\mathbb{E}_{\pi_{0}\!}\!\left[(f(X_{k})-\bar{f})^{2}\right]-\lim\limits_{n\rightarrow\infty}~{}\frac{1}{n}~{}\sum\limits_{k=0}^{n-1}(\bar{f}-\bar{f}_{k})^{2}$
$=\lim\limits_{n\rightarrow\infty}~{}\frac{1}{n}~{}\sum\limits_{k=0}^{n-1}\mathbb{E}_{\pi\!}\!\left[(f(X_{k})-\bar{f})^{2}\right],$
where we used that
$\lim\limits_{n\rightarrow\infty}\frac{1}{n}\sum\limits_{k=0}^{n-1}(\bar{f}-\bar{f}_{k})^{2}=0,$
which follows from geometric mixing of the underlying Markov chain.
Let us now show that the second term in (50) converges to the right limit. Towards this, we first re-write it as
$\mathcal{B}=\frac{2}{n}~{}\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}\left(\mathbb{E}_{\pi_{0}\!}\!\left[(f(X_{i})-\bar{f})(f(X_{k})-\bar{f})\right]+(\bar{f}_{i}-\bar{f})(\bar{f}-\bar{f}_{k})\right),$
and let the corresponding term under $\pi$ be
$\mathcal{B}^{\prime}=\frac{2}{n}~{}\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}\mathbb{E}_{\pi\!}\!\left[(f(X_{i})-\bar{f})(f(X_{k})-\bar{f})\right].$
Let
$T_{1}:=\mathbb{E}_{\pi_{0}\!}\!\left[(f(X_{i})-\bar{f})(f(X_{k})-\bar{f})\right]\quad\text{ and }\quad T_{2}:=\mathbb{E}_{\pi\!}\!\left[(f(X_{i})-\bar{f})(f(X_{k})-\bar{f})\right].$
Then,
$\left|{\mathcal{B}-\mathcal{B}^{\prime}}\right|$ $=\left|{\frac{2}{n}~{}\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}(T_{1}-T_{2})+\frac{2}{n}~{}\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}(\bar{f}_{i}-\bar{f})(\bar{f}-\bar{f}_{k})}\right|$
$\leq\frac{2}{n}~{}\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}\left|{T_{1}-T_{2}}\right|+\frac{2}{n}~{}\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}\left|{\bar{f}_{i}-\bar{f}}\right|\left|{\bar{f}-\bar{f}_{k}}\right|$
$\overset{(a)}{\leq}\frac{2}{n}~{}\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}\left|{T_{1}-T_{2}}\right|+\frac{2}{n}~{}\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}4f^{2}_{\max}C\alpha^{k},$ (52)
where we used Lemma B.4 to get the bound in $(a)$ above, and recall that $f_{\max}:=\max_{x}f(x)$. Clearly, the second term converges to 00 as $n\rightarrow\infty$. Let us now bound the first term.
$\left|{T_{1}-T_{2}}\right|$ $=\left|{\sum\limits_{x\in{\mathcal{S}}}\sum\limits_{y\in{\mathcal{S}}}\sum\limits_{z\in{\mathcal{S}}}\left(\pi_{0}(x)-\pi(x)\right)P^{i}(x,y)P^{k-i}(y,z)(f(y)-\bar{f})(f(z)-\bar{f})}\right|$
$\leq 4f^{2}_{\max}\sum\limits_{z\in{\mathcal{S}}}\left|{\sum\limits_{x\in{\mathcal{S}}}\sum\limits_{y\in{\mathcal{S}}}\left(\pi_{0}(x)-\pi(x)\right)P^{i}(x,y)P^{k-i}(y,z)}\right|$
$=4f^{2}_{\max}\sum\limits_{z\in{\mathcal{S}}}\left|{\sum\limits_{x\in{\mathcal{S}}}\left(\pi_{0}(x)P^{k}(x,z)-\pi(x)P^{k}(x,z)\right)}\right|$
$=\sum\limits_{z\in{\mathcal{S}}}\left|{\sum\limits_{x\in{\mathcal{S}}}\pi_{0}(x)P^{k}(x,z)-\pi(z)}\right|f^{2}_{\max}$
$\leq\sum\limits_{z\in{\mathcal{S}}}\sum\limits_{x\in{\mathcal{S}}}\pi_{0}(x)\left|{P^{k}(x,z)-\pi(z)}\right|f^{2}_{\max}$
$=2L^{2}\sum\limits_{x\in{\mathcal{S}}}\pi_{0}(x)\|P^{k}(x,\cdot)-\pi\|_{\text{T.V.}}$
$\leq 2f^{2}_{\max}C\alpha^{k}.$
Using the above bound in (52), we get that $\lim\limits_{n\rightarrow\infty}\left|{\mathcal{B}-\mathcal{B}^{\prime}}\right|\leq 0$, proving the desired result. ∎
Lemma B.2.
Consider the Markov Chain starting in distribution $\pi_{0}$. For $j\in\mathbb{N}$ and $\gamma(k):=\mathbb{E}_{\pi\!}\!\left[\left(f(X_{k})-\bar{f}\right)\left(f(X_{0})-\bar{f}\right)\right],$ the asymptotic variance of $f$, $\kappa(f)$, is given by
$\kappa(f)$ $=\gamma(0)+2\lim\limits_{n\rightarrow\infty}\sum\limits_{k=1}^{n-1}\gamma(k)=2\sum\limits_{x\in{\mathcal{S}}}\pi(x)(f(x)-\bar{f})V(x)-\sum\limits_{x\in{\mathcal{S}}}\pi(x)(f(x)-\bar{f})^{2}.$
Proof.
From Lemma B.1, we have
$\kappa(f):=\lim\limits_{n\rightarrow\infty}\frac{1}{n}\operatorname{Var}\left[\sum\limits_{k=0}^{n-1}f(X_{k})|X_{0}\sim\pi_{0}\right]$ $=\lim\limits_{n\rightarrow\infty}\frac{1}{n}\operatorname{Var}\left[\sum\limits_{k=0}^{n-1}f(X_{k})|X_{0}\sim\pi\right].$
We show that the r.h.s. in the above equation is same as the expression in the lemma. Recall that for any $k\geq 1$, $\mathbb{E}\!\left[f(X_{k})|X_{0}\sim\pi\right]=\bar{f}$. For simplicity of notation, we denote $\mathbb{E}_{X_{0}\sim\pi_{0}\!}\!\left[\cdot\right]$ by $\mathbb{E}_{\pi_{0}\!}\!\left[\cdot\right]$. Then,
$\operatorname{Var}\left[\sum\limits_{k=0}^{n-1}f(X_{k})|X_{0}\sim\pi\right]$ $=\mathbb{E}_{\pi\!}\!\left[\left(\sum\limits_{k=0}^{n-1}\left(f(X_{k})-\bar{f}\right)\right)^{2}\right]$
$=\mathbb{E}_{\pi\!}\!\left[\sum\limits_{k=0}^{n-1}(f(X_{k})-\bar{f})^{2}\right]$
$\qquad\qquad+2\sum\limits_{k=0}^{n-1}\sum\limits_{i<k}\mathbb{E}_{\pi\!}\!\left[\left(f(X_{i})-\bar{f}\right)\left(f(X_{k})-\bar{f}\right)\right]$
=$\overset{(a)}{=}\mathbb{E}_{\pi\!}\!\left[\sum\limits_{k=0}^{n-1}(f(X_{k})-\bar{f})^{2}\right]$
$\qquad\qquad+2\sum\limits_{k=1}^{n-1}(n-k)\mathbb{E}_{\pi\!}\!\left[\left(f(X_{k})-\bar{f}\right)\left(f(X_{0})-\bar{f}\right)\right]$
$=n\gamma(0)+2\sum\limits_{k=1}^{n-1}(n-k)\gamma(k),$ (53)
where $(a)$ follows from Markov property. Dividing by $n$ and taking limits, we get
$\lim\limits_{n\rightarrow\infty}\frac{1}{n}\operatorname{Var}\left[\sum\limits_{k=0}^{n-1}f(X_{k})|X_{0}\sim\pi\right]$ $=\gamma(0)+\lim\limits_{n\rightarrow\infty}2\sum\limits_{k=1}^{n-1}\frac{n-k}{n}\gamma(k)$
=$\overset{(a)}{=}\gamma(0)+\lim\limits_{n\rightarrow\infty}2\sum\limits_{k=1}^{n-1}\gamma(k)$ (54)
$=\lim\limits_{n\rightarrow\infty}2\sum\limits_{k=0}^{n-1}\frac{n-k}{n}\gamma(k)-\gamma(0)$
=$\overset{(b)}{=}\lim\limits_{n\rightarrow\infty}2\sum\limits_{k=0}^{n-1}\gamma(k)-\gamma(0),$ (55)
where $(a)$ and $(b)$ follow from Lemma B.3. Expression in (54) corresponds to the first expression in the lemma. Next, we show that the expression in (55) is same as the second expression in the lemma.
Since $V$ is the solution of the Poisson equation, we have the following form for $V$ [Douc et al., 2018, Proposition 21.2.3, Lemma 21.2.2]
$V(x)=\sum\limits_{k=0}^{\infty}\mathbb{E}\!\left[f(X_{k})-\bar{f}|X_{0}=x\right]+c,$
for any constant $c$. Substituting this in the expression in the lemma statement gives:
$2\sum\limits_{x\in{\mathcal{S}}}\pi(x)(f(x)-\bar{f})V(x)-\gamma(0)$
$=2\sum\limits_{x\in{\mathcal{S}}}\pi(x)(f(x)-\bar{f})\left(\sum\limits_{k=0}^{\infty}\mathbb{E}\!\left[f(X_{k})-\bar{f}|X_{0}=x\right]\right)-\gamma(0)$
=$\overset{(a)}{=}2\sum\limits_{k=0}^{\infty}\mathbb{E}_{\pi\!}\!\left[(f(X_{0})-\bar{f})\left(f(X_{k})-\bar{f}\right)\right]-\gamma(0)$
$=2\lim\limits_{n\rightarrow\infty}\sum\limits_{k=1}^{n-1}\gamma(k)-\gamma(0),$
which is same as (55). Here, to change the limits and expectation in $(a)$, we used the bounded convergence theorem since $f$ is bounded (underlying Markov chain is on a finite state space). ∎
Lemma B.3.
For $k\in\mathbb{N}$ and $\gamma(k)$ below,
$\lim\limits_{n\rightarrow\infty}\sum\limits_{k=0}^{n-1}\frac{n-k}{n}\gamma(k)=\lim\limits_{n\rightarrow\infty}\sum\limits_{k=0}^{n-1}\gamma(k),$
where
$\gamma(k):=\mathbb{E}_{\pi\!}\!\left[(f(X_{k})-\bar{f})(f(X_{0})-\bar{f})\right].$
Proof.
The proof of this lemma follows along the lines of the proof of [Mou and Maguluri, 2020, Lemma 3]. However, we give the proof for completeness. Let $f_{\max}:=\max\limits_{s\in{\mathcal{S}}}f(x)$. Since $X$ is an irreducible and aperiodic Markov chain on a finite state space, it mixes geometrically fast Levin and Peres [2017], i.e., there exist constants $C>0$ and $\alpha\in(0,1)$ such that for all $k\in\mathbb{N}_{+}$,
$\sup\limits_{x\in{\mathcal{X}}}~{}d_{TV}(P(X_{k}|X_{0}=x),\pi)\leq C\alpha^{k},$
where $d_{TV}(\cdot,\cdot)$ represents the total variation distance between the two input distributions. Below, we first show that $\gamma(k)$ is bounded.
$\left|{\gamma(k)}\right|$ $=\left|{\mathbb{E}_{\pi\!}\!\left[\left(f(X_{k})-\bar{f}\right)\left(f(X_{0})-\bar{f}\right)\right]}\right|$
$=\left|{\mathbb{E}_{\pi\!}\!\left[\mathbb{E}\!\left[\left(f(X_{k})-\bar{f}\right)\left(f(X_{0})-\bar{f}\right)\right]|X_{0}=x\right]}\right|$
$=\left|{\mathbb{E}_{x\sim\pi\!}\!\left[\mathbb{E}\!\left[\left(f(X_{k})-\bar{f}\right)|X_{0}=x\right]\left(f(x)-\bar{f}\right)\right]}\right|$
$\leq\mathbb{E}_{x\sim\pi\!}\!\left[\left|{\mathbb{E}\!\left[\left(f(X_{k})-\bar{f}\right)|X_{0}=x\right]}\right|\left|{f(x)-\bar{f}}\right|\right]$
$\leq 4f^{2}_{\max}C\alpha^{k},$ (56)
where we used Lemma B.4 to get the last inequality above.
Next, define
$V_{1}:=\lim\limits_{n\rightarrow\infty}\sum\limits_{k=0}^{n-1}\gamma(k),\quad\text{ and }\quad V_{2}:=\lim\limits_{n\rightarrow\infty}\sum\limits_{k=0}^{n-1}\frac{n-k}{n}\gamma(k).$
Consider the following:
$\left|{V_{1}-V_{2}}\right|$ $=\lim\limits_{n\rightarrow\infty}\left|{\sum\limits_{k=0}^{n-1}\frac{k}{n}\gamma(k)}\right|\leq\lim\limits_{n\rightarrow\infty}\sum\limits_{k=0}^{n-1}\frac{k}{n}\left|{\gamma(k)}\right|\overset{(a)}{\leq}4f^{2}_{\max}C\lim\limits_{n\rightarrow\infty}\frac{1}{n}\sum\limits_{k=1}^{n-1}k\alpha^{k}\leq 0,$
where in $(a)$ we used Lemma B.4, and the last inequality follows since since $\alpha<1$. This implies $V_{1}=V_{2}$. ∎
Lemma B.4.
Let $X=(X_{k}:k\geq 0)$ be DTMC on a finite-state space ${\mathcal{S}}$ with a unique stationary distribution $\pi$. Let $f:{\mathcal{S}}\rightarrow\mathbb{R}$ be any function, and let $\bar{f}:=\sum_{x\in{\mathcal{S}}}\pi(x)f(x)$ denote its stationary expectation. Then, for any $k\in\mathbb{N}_{+}$, there exist constants $\alpha\in(0,1)$ and $C>0$, such that for any initial distribution $X_{0}\sim\pi_{0}$, and $k>j$,
$\left|{\mathbb{E}\!\left[f(X_{k})-\bar{f}|X_{0}\sim\pi_{0}\right]}\right|\leq 2f_{\max}C\alpha^{k},$
and
$\left|{\mathbb{E}\!\left[(f(X_{k})-\bar{f})(f(X_{j})-\bar{f})|X_{0}\sim\pi_{0}\right]}\right|\leq 4f^{2}_{\max}C^{2}\alpha^{k-j},$
where $f_{\max}:=\max_{x\in{\mathcal{S}}}f(x)$.
Proof.
This lemma follows from the geometric mixing of the underlying Markov chain. Parts of this lemma are borrowed from Mou and Maguluri [2020].
Proving (B.4). Consider the following inequalities:
$\left|{\mathbb{E}\!\left[f(X_{k})-\bar{f}|X_{0}\sim\pi_{0}\right]}\right|$ $=\left|{\sum\limits_{x^{\prime}\in{\mathcal{S}}}\pi_{0}(x^{\prime})\sum\limits_{x\in{\mathcal{S}}}f(x)\left(\mathbb{P}\left(X_{k}=x|X_{0}=x^{\prime}\right)-\pi(x)\right)}\right|$
$\leq\sum\limits_{x^{\prime}\in{\mathcal{S}}}\pi_{0}(x^{\prime})\sum\limits_{x\in{\mathcal{S}}}\left|{f(x)}\right|\left|{\mathbb{P}\left(X_{k}=x|X_{0}=x^{\prime}\right)-\pi(x)}\right|$
$\leq f_{\max}\sum\limits_{x^{\prime}\in{\mathcal{S}}}\pi_{0}(x^{\prime})\sum\limits_{x\in{\mathcal{S}}}\left|{\mathbb{P}\left(X_{k}=x|X_{0}=x^{\prime}\right)-\pi(x)}\right|$
$\leq f_{\max}~{}\sup\limits_{x^{\prime}\in{\mathcal{S}}}~{}\sum\limits_{x\in{\mathcal{S}}}\left|{\mathbb{P}\left(X_{k}=x|X_{0}=x^{\prime}\right)-\pi(x)}\right|$
$\leq 2f_{\max}C\alpha^{k},$
where the last inequality follows from the definition of total variation distance and geometric mixing of the underlying Markov chain.
Proving (B.4). We now prove the other inequality using the geometric mixing of the Markov chain. Let $\mathbb{E}_{\pi_{0}\!}\!\left[\cdot\right]$ represent $\mathbb{E}\!\left[\cdot|X_{0}\sim\pi_{0}\right]$.
$\left|{\mathbb{E}_{\pi_{0}\!}\!\left[(f(X_{k})-\bar{f})(f(X_{j})-\bar{f})\right]}\right|$ $=\left|{\mathbb{E}_{\pi_{0}\!}\!\left[(f(X_{j})-\bar{f})\mathbb{E}\!\left[(f(X_{k})-\bar{f})|X_{j}\right]\right]}\right|$
$\leq\mathbb{E}_{\pi_{0}\!}\!\left[|f(X_{j})-\bar{f}|\right]\left|{\sup\limits_{x\in{\mathcal{S}}}\mathbb{E}\!\left[(f(X_{k})-\bar{f})|X_{j}=x\right]}\right|.$ (59)
Since (B.4) holds for any initial distribution $\pi_{0}$, we have
$\left|{\sup_{x}\mathbb{E}\!\left[(f(X_{k})-\bar{f})|X_{j}=x\right]}\right|=\sup_{x}\mathbb{E}\!\left[(f(X_{k-j})-\bar{f})|X_{0}=x\right]\leq 2f_{\max}C\alpha^{k-j}.$
Using the above bound in (59),
$\left|{\mathbb{E}_{\pi_{0}\!}\!\left[(f(X_{k})-\bar{f})(f(X_{j})-\bar{f})\right]}\right|$ $\leq 2f_{\max}C\alpha^{k-j}\mathbb{E}_{\pi_{0}\!}\!\left[\left|{(f(X_{j})-\bar{f})}\right|\right]\leq 4f^{2}_{\max}C\alpha^{k-j},$
proving the desired result.
B.3 Discussion on Batched Mean Estimators
A natural first approach to estimate $\kappa(f)$ is to use the Monte Carlo (MC) technique. For this, observe that
$\kappa=\lim\limits_{m\rightarrow\infty}\kappa_{m},\quad\text{where}\quad\kappa_{m}:=\operatorname{Var}[Z_{m}|X_{0}\sim\pi],\quad\text{and}\quad Z_{m}:=\frac{1}{\sqrt{m}}\sum_{k=0}^{m-1}f(X_{k}).$
A vanilla MC estimator, popularly known as Batched Means (BM) estimator, proceeds by generating a single large trajectory of the underlying Markov chain $\mathcal{M}$ starting from any initial distribution $\pi_{0}$ on ${\mathcal{S}}$. It then considers samples in non-overlapping buckets of a fixed length, say $m$. For $i\in\mathbb{N}$, let
$Z^{(i)}_{m}:=\frac{1}{\sqrt{m}}\sum\limits_{k=(i-1)m}^{i(m-1)}f(X_{k}).$
BM method estimates the first and second moments of $Z_{m}$, defined in (B.3), using sample averages of $Z^{(i)}_{m}$ across $b$ buckets, say, and combines them to get an estimate for $\kappa_{m}$.
The asymptotic convergence of BM (as $b\rightarrow\infty$) to $\kappa_{m}$ follows from the Ergodic Theorem for Markov chains [Douc et al., 2018, Theorem 5.2.1]. However, a drawback of BM is that approximating the limit in $\kappa$ by the variance of a truncated random variable $Z_{m}$ introduces a bias, which corresponds to $\left|{\kappa-\kappa_{m}}\right|$. Chien et al. [1997] established that the mean-squared estimation error for estimating $\kappa$ using BM is $O(\frac{1}{m^{2}}+\frac{1}{b})$, where the $O(\frac{1}{m^{2}})$ term corresponds to the bias. Thus, estimating $\kappa$ with mean-squared estimation error at most $\epsilon^{2}$ requires $m=O(\frac{1}{\epsilon})$, and $b=O(\frac{1}{\epsilon^{2}})$. This corresponds to a total of $mb=O(\frac{1}{\epsilon^{3}})$ samples. While this bound can be improved by designing better MC estimators, see Flegal and Jones [2010], MC estimators typically also suffer from being non-adaptive (or non-recursive), i.e., $m$ is a function of $\epsilon$ this necessitates restarting the procedure for different choices of $\epsilon$.
Appendix C Proof of Results in Section 2
C.1 Proof of Theorem 2.1
For $s\in{\mathcal{S}}$, let ${\bf 1}(s)\in\left\{0,1\right\}^{\left|{{\mathcal{S}}}\right|}$ be the indicator for $s$, i.e., it equals $1$ at the $s^{\text{th}}$ coordinate, and 00 otherwise. Furthermore, for $n\in\left\{1,2,\dots\right\}$, define ${\bf 1}_{n}:={\bf 1}(S_{n})$, indicator for $S_{n}$. Then, setting $d=\left|{{\mathcal{S}}}\right|$, basis vectors in the setup in Section 3 as standard basis, and for $k\in\left\{1,2,\dots\right\}$, $\phi_{k}={\bf 1}_{k}$, we recover Algorithm 1 from updates in Section 3.1.
Further, since in this setting, the set of fixed points $V$ of (1.3) lies in the span of the basis vectors, $\kappa^{*}$ from (19) is same as $\kappa(f)$. Hence, guarantees in Theorem 3.1 with these adjustments, reduce to those in Theorem 2.1, proving the result.
C.2 Proof of Corollary 2.1.1
We prove the sample complexity bound separately for the algorithm with a constant step size, and that with the diminishing step size. But first, observe that $c_{1}=O({1}/{\Delta_{1}})$, $c_{2}=O(\Delta_{1})$, and $c_{3}=O(\Delta_{1})$, implying that $\eta=O(1/\Delta_{1})$. Furthermore, $\xi_{1}=O(\|V^{*}\|^{2}_{2})$, and $\xi_{2}=O(\|V^{*}\|^{2}_{2})$.
  1. (a)
    For Algorithm 1 with a constant step size $\alpha$, to estimate $\kappa(f)$ up to a mean estimation error of at most $\epsilon$, we require
    $n\geq O\left(\frac{1}{\Delta_{1}\alpha}\log\frac{\xi_{1}}{\epsilon^{2}}\right)=O\left(\frac{\log\frac{1}{\epsilon^{2}}}{\epsilon^{2}}\right)\tilde{O}\left(\frac{\eta^{2}\|V^{*}\|^{2}_{2}}{\Delta^{2}_{1}}\right)=O\left(\frac{\log\frac{1}{\epsilon^{2}}}{\epsilon^{2}}\right)\tilde{O}\left(\frac{\|V^{*}\|^{2}_{2}}{\Delta^{4}_{1}}\right).$
    This follows from Theorem 2.1(a), where setting the second term to at most $\epsilon^{2}$ gives
    $\alpha=O\left(\epsilon^{2}\right)O\left(\frac{\Delta_{1}}{\|V^{*}\|^{2}_{2}\eta^{2}}\right).$
    Choosing $\alpha$ to satisfy this condition, and setting the first term in the error bound to $\epsilon^{2}$, we get the required sample complexity bound.
  2. (b)
    Next, for Algorithm 1 with diminishing step size $\alpha_{i}=\frac{\alpha}{i+h}$ for all $i\geq 1$, for estimating $\kappa$ up to a mean estimation error at most $\epsilon$, we require
    $n+h=O\left(\frac{\alpha^{2}\xi_{2}\eta^{2}}{\epsilon^{2}}\right)=O\left(\frac{1}{\epsilon^{2}}\right)O\left(\frac{\eta^{2}\|V^{*}\|^{2}_{2}}{\Delta^{2}_{1}}\right)=O\left(\frac{1}{\epsilon^{2}}\right)O\left(\frac{\|V^{*}\|^{2}_{2}}{\Delta^{4}_{1}}\right).$
    This follows from the mean-square estimation error bound in Theorem 2.1(b) by setting the rate-determining second term to at most $\epsilon^{2}$, after optimizing the term
    $\left(\frac{\alpha^{2}}{\Delta_{1}\alpha-40}\right)$
    over $\alpha$. Note that $\alpha=O(1/\Delta_{1})$ is the optimal choice. Further, with this choice of $\alpha$, since $h=O(1/\Delta^{4})$, we get
    $n=O\left(\frac{1}{\epsilon^{2}}\right)O\left(\frac{\|V^{*}\|^{2}_{2}}{\Delta^{4}_{1}}\right).$
C.3 MSE Bound for Stationary Variance Estimation from Section 2.3
In this section, we prove Theorem 2.2 in $2$ steps, that we discuss below.
C.3.1 Step 1: Algorithm 2 as a Linear SA
Algorithm 2 for estimating the stationary variance of $f$, $v(f)$, can be formulated as a linear SA update, as discussed in this section. Consider the following matrices:
$A^{S}_{k}:=A^{S}(X_{k})=\begin{bmatrix}-1&0\\ -cf(X_{k})&-c\end{bmatrix},\quad\text{and}\quad b^{S}_{k}:=b^{S}(X_{k})=\begin{bmatrix}f(X_{k})\\ cf^{2}(X_{k})\end{bmatrix}.$
For $(\Theta^{S}_{k})^{T}:=[\bar{f}_{k}~{}v_{k}]$, Algorithm 2 corresponds to the following update
$\Theta^{S}_{k+1}=\Theta^{S}_{k}+\alpha_{k}(A^{S}_{k}\Theta^{S}_{k}+b^{S}_{k}).$
Define $A^{S}:=\mathbb{E}_{\pi\!}\!\left[A_{k}\right]$ and $b^{S}:=\mathbb{E}_{\pi\!}\!\left[b_{k}\right]$. Then,
$A^{S}=\begin{bmatrix}-1&0\\ -c\bar{f}&-c\end{bmatrix},\quad\text{and}\quad b^{S}=\begin{bmatrix}\bar{f}\\ c\mathbb{E}_{\pi\!}\!\left[f^{2}\right]\end{bmatrix}.$
We now prove a contraction property for the average matrix $A^{S}$.
Lemma C.1.
For any $\gamma>1$ and $c>0$ such that
$c{\bar{f}}^{2}\leq 2\left(-(\gamma-1)+\sqrt{(\gamma-1)\left(\gamma-1+\gamma{\bar{f}}^{2}\right)}\right),$
matrix $A^{S}$ satisfies
$\min\limits_{\Theta\in\mathbb{R}^{2},\|\Theta\|_{2}=1}-\Theta^{T}A^{S}\Theta>\gamma>0.$
Proof.
Consider $\Theta^{T}=[\theta_{1}~{}\theta_{2}]$. Then,
$-\Theta^{T}A^{S}\Theta=\theta^{2}_{1}+c\bar{f}\theta_{1}\theta_{2}+c\theta^{2}_{2}.$
Now, consider the following:
$\min\limits_{\begin{subarray}{c}\theta_{1}\in\mathbb{R},\theta_{2}\in\mathbb{R}\\ \theta^{2}_{1}+\theta^{2}_{2}=1\end{subarray}}-\Theta^{T}A^{S}\Theta$ $\geq\min\limits_{\begin{subarray}{c}\theta_{1}\in\mathbb{R},\theta_{2}\in\mathbb{R},\\ \theta^{2}_{1}+\theta^{2}_{2}=1\end{subarray}}~{}\theta^{2}_{1}-c|\bar{f}||\theta_{1}||\theta_{2}|+c\theta^{2}_{2}$
$=\min\limits_{\begin{subarray}{c}|\theta_{1}|\leq 1\end{subarray}}~{}\theta^{2}_{1}-c\left|{\bar{f}}\right||\theta_{1}|\sqrt{1-\theta^{2}_{1}}+c(1-\theta^{2}_{1})$
$=\min\limits_{\begin{subarray}{c}|\theta_{1}|\leq 1\end{subarray}}~{}(1-c)\theta^{2}_{1}-c\left|{\bar{f}}\right||\theta_{1}|\sqrt{1-\theta^{2}_{1}}+c$
$=\frac{1+c}{2}-\frac{1}{2}\sqrt{(1-c)^{2}+c^{2}(\bar{f})^{2}}.$ (Lemma G.3)
This lower bound greater than $\gamma$ is equivalent to
$\frac{c^{2}\bar{f}^{2}}{4}-c(1-\gamma)+\gamma-\gamma^{2}\leq 0,$
which has a solution (for $c$) if $\gamma>1$ or $\gamma<\frac{1}{1+\bar{f}^{2}}$. For $\gamma>1$, the upper bound on $c$ in the condition in the statement of the lemma is the positive root of the above equation. ∎
To make the above condition explicit for specific choices for $\gamma$,
$\gamma=1+\bar{f}^{2}\implies c\leq 2\left[-1+\sqrt{2+\bar{f}^{2}}\right];$
$\gamma=2\implies c{\bar{f}^{2}}\leq 2\left[-1+\sqrt{1+2\bar{f}^{2}}\right].$
C.3.2 Step 2: Proving the MSE Bound
Let $\Theta^{s}:=[\bar{f}~{}v(f)]$ and $\Theta_{k}=[\bar{f}_{k}~{}v_{k}]$. Define
$\eta=\max\left\{1,\|A^{S}\|_{2},\|b^{S}\|_{2}\right\}.$
The MSE bound in Theorem 2.2 follows from Theorem A.1. To see that the conditions of Theorem A.1 are satisfied, first observe that the iterates $\Theta^{S}$ are generated by a linear SA, as formulated in the previous section (Section C.3.1). Furthermore, for $d=2$ and subspace $\mathbb{S}=\mathbb{R}^{2}$, Lemma C.1 establishes the required contraction property for the averaged matrix $A^{S}$, and $\eta$ in this section corresponds to the constant H . Then, the iterates $\Theta^{S}_{k}$ satisfy the bounds in Theorem A.1 with $\Theta^{*}=[\bar{f}~{}v(f)]$, and $\gamma_{2}=\gamma$.
C.4 Estimating Variance: IID Setting
Consider $X_{1},X_{2},\dots$, independent samples from a distribution with mean 00 and an unknown variance $\sigma^{2}$. Further assume that the sampling distribution has bounded $4^{th}$ moment. The goal is to estimate $\sigma^{2}$ using these samples.
Since, in this setting, $\operatorname{Var}\left[X\right]=\mathbb{E}\!\left[X^{2}\right]$, let
$\hat{\sigma}^{2}_{n}:=\frac{1}{n}\sum\limits_{j=1}^{n}X^{2}_{j}$
denote the estimate for $\sigma^{2}$ using $n$ samples. Clearly, $\mathbb{E}\!\left[\hat{\sigma}^{2}_{n}\right]=\sigma^{2}$, i.e., $\hat{\sigma}^{2}_{n}$ is an unbiased estimator.
Next, consider the following
$\mathbb{E}\!\left[(\hat{\sigma}^{2}_{n}-\sigma^{2})^{2}\right]$ $=\mathbb{E}\!\left[(\hat{\sigma}^{2}_{n})^{2}+\sigma^{4}-2\sigma^{2}\hat{\sigma}^{2}_{n}\right]$
$=\mathbb{E}\!\left[(\hat{\sigma}^{2}_{n})^{2}\right]-\sigma^{4}$
$=\frac{1}{n^{2}}\mathbb{E}\!\left[\sum\limits_{j=1}^{n}X^{4}_{j}+\sum\limits_{i\neq j}X^{2}_{i}X^{2}_{j}\right]-\sigma^{4}$
$=\frac{\mathbb{E}\!\left[X^{4}\right]}{n}+\frac{n(n-1)\sigma^{4}}{n^{2}}-\sigma^{4}$
$=\frac{1}{n}\left(\mathbb{E}\!\left[X^{4}\right]-\sigma^{4}\right)$
$=\frac{c_{1}}{n},$
where $c_{1}=\mathbb{E}\!\left[X^{4}\right]-(\mathbb{E}\!\left[X^{2}\right])^{2}$. Thus, we have that the mean-squared estimation error in this setting is exactly $O(\frac{1}{n})$. In fact, Cramér-Rao lower bound Nielsen [2013] for the mean-squared error in estimating $\sigma^{2}$ using an unbiased estimators is $O(\frac{1}{n})$, establishing that this rate cannot be improved in certain settings.
For $\sigma>0$, let $\mathcal{N}(0,\sigma^{2})$ denote the Gaussian distribution with variance $\sigma^{2}$. If $X\sim\mathcal{N}(0,\sigma^{2})$, $c_{1}$ equals $3\sigma^{2}$, and in this case we have that
$\mathbb{E}\!\left[\left(\hat{\sigma}^{2}_{n}-\sigma^{2}\right)^{2}\right]=\frac{3\sigma^{2}}{n},\quad\text{ for $c>0$}.$
Appendix D Details of Proof for MSE Bound in Theorem 3.1
In this section, we give the missing details and proofs of results from Section 3.4 (proof of Theorem 3.1). As discussed in the main text, the proof of Theorem 3.1 proceeds in two steps. We give details of Step 1 in Appendix D.1, and that for the Step 2 in Appendix D.2.
D.1 Details of Step 1: Algorithm as a Linear SA
In this section, we add details to the discussion in Section 3.4.1 about formulating the proposed algorithm as a linear SA update. Below, we first present the exact form of the average matrices (Section D.1.1), followed by a discussion on the limit point (if convergence happens) of the proposed updates (Section D.1.2). We then conclude with a proof for Lemma 3.3 in Section D.1.3.
D.1.1 Stationary-Average Matrices
The average matrices $A$ and $b$ are given by
$A=\begin{bmatrix}-c_{1}&\vec{\bf 0}^{T}&0&0\\ -\Pi_{2,E}\Phi^{T}D_{\pi}\vec{\bf 1}&{\Pi_{2,E}\Phi^{T}D_{\pi}(P-I)\Phi}&\vec{\bf 0}&\vec{\bf 0}\\ 0&c_{2}\vec{\bf 1}^{T}D_{\pi}\Phi&-c_{2}&0\\ c_{3}\bar{f}&2c_{3}\mathcal{F}^{T}D_{\pi}\Phi&-2c_{3}\bar{f}&-c_{3}\end{bmatrix},\quad\text{and}\quad b=\begin{bmatrix}c_{1}\bar{f}\\ \Pi_{2,E}\Phi^{T}D_{\pi}\mathcal{F}\\ 0\\ -{c_{3}}\mathbb{E}_{\pi\!}\!\left[f^{2}(X)\right]\end{bmatrix},$
where $P$ is the transition matrix for the Markov chain $\mathcal{M}$, and $I$ is the identity matrix. This follows from linearity of the expectation (projection is a linear operator) and the observation that the stationary expectation of each entry of matrix $A(\cdot)$ and $b(\cdot)$ is given by the corresponding entry of $A$ and $b$, respectively.
D.1.2 Algorithm’s Limit: A Discussion
Let $\Theta^{*T}:=[\bar{f}~{}\theta^{*T}~{}\widetilde{V}~{}\kappa^{*}]$, where recall that
$\kappa^{*}=\mathbb{E}_{\pi\!}\!\left[2f(X)[\Phi\theta^{*}](X)-2f(X)\widetilde{V}-f^{2}(X)+f(X)\bar{f}\right],$
$\Phi\theta^{*}\in\mathbb{R}^{S}$, and $[\Phi\theta^{*}](i)$ represents its $i^{th}$ component, $\widetilde{V}=\mathbb{E}_{\pi\!}\!\left[[\Phi\theta^{*}](X)\right]$, and $\theta^{*}$ is the unique vector in $E$ that is also a solution for
$\Phi\theta=\Pi_{D_{\pi},W_{\Phi}}T\Phi\theta,$
where $\Pi_{D_{\pi},W_{\Phi}}$ is the projection matrix onto $W_{\Phi}:=\left\{\Phi\theta|\theta\in\mathbb{R}^{d}\right\}$ with respect to $D_{\pi}$ norm. Here, $T$ is an operator that for a vector $V\in\mathbb{R}^{S}$, satisfies $TV=\mathcal{F}-\bar{f}{\bf e}+PV$.
Let $\Theta_{\infty}:=[{\bar{f}}_{\infty}~{}\theta_{\infty}~{}\widetilde{V}_{\infty}~{}\kappa_{\infty}]$, where $\theta_{\infty}\in E$. We now show that $\Theta_{\infty}=\Theta^{*}$ is the unique solution to $A\Theta_{\infty}+b=0$. This equation corresponds to the following:
$-c_{1}\bar{f}_{\infty}+c_{1}\bar{f}$ $=0,$ (61)
$-\Pi_{2,E}\Phi^{T}D_{\pi}{\bf 1}\bar{f}_{\infty}+\Pi_{2,E}\Phi^{T}D_{\pi}(P-I)\Phi\theta_{\infty}+\Pi_{2,E}\Phi^{T}D_{\pi}\mathcal{F}$ $=0,$ (62)
$c_{2}\vec{\bf 1}^{T}D_{\pi}\Phi\theta_{\infty}-c_{2}\widetilde{V}_{\infty}$ $=0,$ (63)
$c_{3}\bar{f}\bar{f}_{\infty}+2c_{3}\mathcal{F}^{T}D_{\pi}\Phi\theta_{\infty}-2c_{3}\bar{f}\widetilde{V}_{\infty}-c_{3}\kappa_{\infty}-c_{3}\mathbb{E}_{\pi\!}\!\left[f^{2}\right]$ $=0.$ (64)
Clearly, (61) implies that $\bar{f}_{\infty}=\bar{f}$. Moreover, if $\theta_{\infty}=\theta^{*}+a\theta_{e}$, for $a\in\mathbb{R}$, where $\theta_{e}$ is the unique vector in $\mathbb{R}^{d}$ such that $\Phi\theta_{e}={\bf e}$, then (63) implies that $\widetilde{V}_{\infty}=\mathbb{E}_{\pi\!}\!\left[\Phi(\theta^{*}+a\theta_{e})\right]=\widetilde{V}+a{\bf e}$ and then, (64) implies $\kappa_{\infty}=\kappa^{*}$. It remains to show that $\theta_{\infty}=\theta^{*}+a\theta_{e}$, for some $a\in\mathbb{R}$. In fact, since $\theta_{\infty}\in E$, we show below that $\theta_{\infty}=\theta^{*}$.
Let’s use $\bar{f}_{\infty}=\bar{f}$, and re-write the LHS of (62) as below:
$\Pi_{2,E}\Phi^{T}D_{\pi}\left((\mathcal{F}-\bar{f}{\bf e})+(P-I)\Phi\theta_{\infty}\right).$
Clearly, it equals 00 for $\theta_{\infty}=\theta^{*}$. This follows from the definition of $\theta^{*}$. Hence, $\theta^{*}$ is a solution for (62). Furthermore, $\theta=\theta^{*}+c\theta_{e}$ for any $c\in\mathbb{R}$ cannot be a solution, since $\theta^{*}+c\theta_{e}\notin E$ and $\theta_{\infty}\in E$. We now show that there does not exist any $\theta\in E$ different from $\theta^{*}$ that satisfies (62). To this end, suppose such a $\theta^{\prime}\in E$ exists. Then, (62) evaluated at $\theta^{\prime}$ re-writes as
$\Pi_{2,E}\Phi^{T}D_{\pi}\left(\underbrace{-\bar{f}{\bf e}+(P-I)\Phi\theta^{*}+\mathcal{F}}_{=0}\right)+\Pi_{2,E}\Phi^{T}D_{\pi}(P-I)\Phi(\theta^{\prime}-\theta^{*}),$
where the first term equals 00 from the definition of $\theta^{*}$. Now, recall that $\Pi_{2,E}=I-\theta_{e}\theta^{T}_{e}$, with the convention that $\theta_{e}=0$ if ${\bf e}\notin W_{\Phi}$, the column space of $\Phi$. Further, $\theta^{T}_{e}\Phi^{T}={\bf e}^{T}\in\mathbb{R}^{S}$, which implies
$\theta^{T}_{e}\Phi^{T}D_{\pi}(P-I)=0.$
Using these, the second term above equals
$\Phi^{T}D_{\pi}(P-I)\Phi(\theta^{\prime}-\theta^{*})-\theta_{e}\underbrace{\theta^{T}_{e}\Phi^{T}D_{\pi}(P-I)\Phi(\theta^{\prime}-\theta^{*})}_{=0}.$
Now, the first term above is non-zero since $\theta^{\prime}-\theta^{*}\in E$, and hence, $\Phi(\theta^{\prime}-\theta^{*})$ is non-constant vector that does not belong to the null space of $P-I$. Thus, $\nexists\theta^{\prime}\in E$ different from $\theta^{*}$ that also satisfies (62).
D.1.3 Proof of Lemma 3.3
In this appendix, we will show the following:
$\min\limits_{\begin{subarray}{c}\Theta\in\mathbb{R}\times E\times\mathbb{R}\times\mathbb{R},\\ \|\Theta\|^{2}_{2}=1\end{subarray}}~{}-\Theta^{T}A\Theta>\frac{\Delta_{2}}{20}>0,$
where
$\Delta_{2}:=\min\limits_{\|\theta\|_{2}=1,\theta\in E}~{}\theta^{T}\Phi^{T}D_{\pi}(I-P)\Phi\theta.$
Recall that for ${\theta}\in E$, ${\Phi\theta}$ is a non-constant vector in $\mathbb{R}^{S}$. Thus, $\theta^{T}\Phi^{T}D_{\pi}(I-P)\Phi{\theta}>0,$ for ${\theta}\in E$ [Tsitsiklis and Van Roy, 1997, Lemma 7]. Since the set $\left\{\theta\in E:\|\theta\|^{2}_{2}=1\right\}$ is non-empty and compact, by extreme value theorem, we have $\Delta_{2}>0$.
Next, we re-write the above minimization problem as (set $\gamma=1$ in the following)
$\min\limits_{\begin{subarray}{c}\|\Theta\|^{2}_{2}=1,\\ \vec{\theta}_{2}\in E\end{subarray}}~{}-\Theta^{T}A\Theta$
$=\min\limits_{\begin{subarray}{c}\|\Theta\|^{2}_{2}=1,\\ \vec{\theta}_{2}\in E\end{subarray}}~{}~{}c_{1}\theta^{2}_{1}+\gamma\vec{\theta}^{T}_{2}\Pi_{2,E}\Phi^{T}D_{\pi}\vec{1}\theta_{1}+\gamma\vec{\theta}^{T}_{2}\Pi_{2,E}\Phi^{T}D_{\pi}(I-P)\Phi\vec{\theta}_{2}-c_{2}\vec{1}^{T}D_{\pi}\Phi\vec{\theta}_{2}\theta_{3}$
$\qquad\qquad\qquad+c_{2}\theta^{2}_{3}-c_{3}\bar{f}\theta_{1}\theta_{4}-2c_{3}\mathcal{F}^{T}D_{\pi}\Phi\vec{\theta}_{2}\theta_{4}+2c_{3}\bar{f}\theta_{3}\theta_{4}+c_{3}\theta^{2}_{4}$
=$\overset{(a)}{=}\min\limits_{\begin{subarray}{c}\|\Theta\|^{2}_{2}=1,\\ \vec{\theta}_{2}\in E\end{subarray}}~{}~{}c_{1}\theta^{2}_{1}+\gamma\vec{\theta}^{T}_{2}\Phi^{T}D_{\pi}\vec{1}\theta_{1}+\gamma\vec{\theta}^{T}_{2}\Phi^{T}D_{\pi}(I-P)\Phi\vec{\theta}_{2}-c_{2}\vec{1}^{T}D_{\pi}\Phi\vec{\theta}_{2}\theta_{3}$
$\qquad\qquad\qquad+c_{2}\theta^{2}_{3}-c_{3}\bar{f}\theta_{1}\theta_{4}-2c_{3}\mathcal{F}^{T}D_{\pi}\Phi\vec{\theta}_{2}\theta_{4}+2c_{3}\bar{f}\theta_{3}\theta_{4}+c_{3}\theta^{2}_{4}$
$\overset{(b)}{\geq}\min\limits_{\begin{subarray}{c}\|\Theta\|^{2}_{2}=1,\\ \vec{\theta}_{2}\in E\end{subarray}}~{}~{}c_{1}\theta^{2}_{1}+\gamma\vec{\theta}^{T}_{2}\Phi^{T}D_{\pi}\vec{1}\theta_{1}+\gamma\Delta_{2}\|\vec{\theta}_{2}\|^{2}_{2}-c_{2}\vec{1}^{T}D_{\pi}\Phi\vec{\theta}_{2}\theta_{3}$
$\qquad\qquad\qquad+c_{2}\theta^{2}_{3}-c_{3}\bar{f}\theta_{1}\theta_{4}-2c_{3}\mathcal{F}^{T}D_{\pi}\Phi\vec{\theta}_{2}\theta_{4}+2c_{3}\bar{f}\theta_{3}\theta_{4}+c_{3}\theta^{2}_{4}$
$=\min\limits_{\begin{subarray}{c}\|\Theta\|^{2}_{2}=1,\\ \vec{\theta}_{2}\in E\end{subarray}}~{}~{}c_{1}\theta^{2}_{1}+\gamma\Delta_{2}\|\vec{\theta}_{2}\|^{2}_{2}+c_{2}\theta^{2}_{3}+c_{3}\theta^{2}_{4}$
$\qquad\qquad\qquad+\vec{\theta}^{T}_{2}\Phi^{T}D_{\pi}\left(\gamma\vec{1}\theta_{1}-c_{2}\vec{1}\theta_{3}-2c_{3}\theta_{4}\mathcal{F}\right)+2c_{3}\bar{f}\theta_{3}\theta_{4}-c_{3}\bar{f}\theta_{1}\theta_{4}.$
where, $(a)$ follows since $\theta_{2}\in E$, and $(b)$ follows from the definition of $\Delta_{2}$. Now,
$\left|{\vec{\theta}^{T}_{2}\Phi^{T}D_{\pi}\left(\gamma\vec{1}\theta_{1}-c_{2}\vec{1}\theta_{3}-2c_{3}\theta_{4}\mathcal{F}\right)}\right|$
$\leq\left|{\gamma\theta_{1}\vec{\theta}^{T}_{2}\Phi^{T}D_{\pi}\vec{1}}\right|+\left|{c_{2}\theta_{3}\vec{\theta}^{T}_{2}\Phi^{T}D_{\pi}\vec{1}}\right|+\left|{2c_{3}\theta_{4}\vec{\theta}^{T}_{2}\Phi^{T}D_{\pi}\mathcal{F}}\right|$
$\leq\gamma\left|{\theta_{1}}\right|\|\Phi\vec{\theta}_{2}\|_{\infty}\|D_{\pi}\vec{1}\|_{1}+c_{2}\left|{\theta_{3}}\right|\|\Phi\vec{\theta}_{2}\|_{\infty}\|D_{\pi}\vec{1}\|_{1}+2c_{3}\left|{\theta_{4}}\right|\|\Phi\vec{\theta}_{2}\|_{\infty}\|D_{\pi}\mathcal{F}\|_{1}$
$\leq\gamma\left|{\theta_{1}}\right|\|\Phi\vec{\theta}_{2}\|_{\infty}+c_{2}\left|{\theta_{3}}\right|\|\Phi\vec{\theta}_{2}\|_{\infty}+2c_{3}\left|{\theta_{4}}\right|\|\Phi\vec{\theta}_{2}\|_{\infty}f_{\max}$
$\overset{(a)}{\leq}\gamma\left|{\theta_{1}}\right|\|\vec{\theta}_{2}\|_{2}+c_{2}\left|{\theta_{3}}\right|\|\vec{\theta}_{2}\|_{2}+2c_{3}\left|{\theta_{4}}\right|\|\vec{\theta}_{2}\|_{2}f_{\max},$
where $f_{\max}:=\max_{s}f(s)$, and $(a)$ follows since
$\|\Phi\vec{\theta}_{2}\|_{\infty}\leq\max\limits_{s\in{\mathcal{S}}}\|\phi(s)\|_{2}\|\vec{\theta}_{2}\|_{2}\leq\|\vec{\theta}_{2}\|_{2},$
where the first inequality follows from the definition of $\|\cdot\|_{\infty}$ and Holder’s inequality, and the second from Assumption 1.
Next, using the previous bound,
$\min\limits_{\begin{subarray}{c}\|\Theta\|^{2}_{2}=1,\\ \vec{\theta}_{2}\in E\end{subarray}}~{}-\Theta^{T}A\Theta$
$\geq$ $\min\limits_{\begin{subarray}{c}\|\Theta\|^{2}_{2}=1,\\ \vec{\theta}_{2}\in E\end{subarray}}~{}~{}c_{1}\theta^{2}_{1}+\gamma\Delta_{2}\|\vec{\theta}_{2}\|^{2}_{2}+c_{2}\theta^{2}_{3}+c_{3}\theta^{2}_{4}+2c_{3}\bar{f}\theta_{3}\theta_{4}-c_{3}\bar{f}\theta_{1}\theta_{4}$
$\qquad\qquad\qquad-\gamma\left|{\theta_{1}}\right|\|\vec{\theta}_{2}\|_{2}-c_{2}\left|{\theta_{3}}\right|\|\vec{\theta}_{2}\|_{2}-2c_{3}\left|{\theta_{4}}\right|\|\vec{\theta}_{2}\|_{2}f_{\max}$
$=$ $\min\limits_{\begin{subarray}{c}\theta^{2}_{3}+\theta^{2}_{4}\in[0,1]\end{subarray}}~{}~{}\min\limits_{\begin{subarray}{c}\theta^{2}_{1}+\|\vec{\theta}_{2}\|^{2}_{2}=1-\theta^{2}_{3}-\theta^{2}_{4}\\ \vec{\theta}_{2}\in E,\theta_{1}\in\mathbb{R}\end{subarray}}~{}~{}c_{1}\theta^{2}_{1}+\gamma\Delta_{2}\|\vec{\theta}_{2}\|^{2}_{2}+c_{2}\theta^{2}_{3}+c_{3}\theta^{2}_{4}+2c_{3}\bar{f}\theta_{3}\theta_{4}-c_{3}\bar{f}\theta_{1}\theta_{4}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\gamma\left|{\theta_{1}}\right|\|\vec{\theta}_{2}\|_{2}-c_{2}\left|{\theta_{3}}\right|\|\vec{\theta}_{2}\|_{2}-2c_{3}\left|{\theta_{4}}\right|\|\vec{\theta}_{2}\|_{2}f_{\max}$
$=$ $\min\limits_{\begin{subarray}{c}\theta^{2}_{3}+\theta^{2}_{4}\in[0,1]\end{subarray}}~{}c_{2}\theta^{2}_{3}+c_{3}\theta^{2}_{4}+2c_{3}\bar{f}\theta_{3}\theta_{4}+(P1)$
$\geq$ $\min\limits_{\begin{subarray}{c}\theta^{2}_{3}+\theta^{2}_{4}\in[0,1]\end{subarray}}~{}c_{2}\theta^{2}_{3}+c_{3}\theta^{2}_{4}-2c_{3}f_{\max}\left|{\theta_{3}}\right|\left|{\theta_{4}}\right|+(P1),$ (66)
where,
$(P1):=\min\limits_{\begin{subarray}{c}\theta^{2}_{1}+\|\vec{\theta}_{2}\|^{2}_{2}=1-\theta^{2}_{3}-\theta^{2}_{4}\\ \vec{\theta}_{2}\in E,\theta_{1}\in\mathbb{R}\end{subarray}}$$~{}~{}\big{\{}c_{1}\theta^{2}_{1}+\gamma\Delta_{2}\|\vec{\theta}_{2}\|^{2}_{2}-c_{3}\bar{f}\theta_{1}\theta_{4}-\gamma\left|{\theta_{1}}\right|\|\vec{\theta}_{2}\|_{2}$
$\qquad\qquad\qquad\qquad-c_{2}\left|{\theta_{3}}\right|\|\vec{\theta}_{2}\|_{2}-2c_{3}\left|{\theta_{4}}\right|\|\vec{\theta}_{2}\|_{2}f_{\max}\big{\}}.$
Suppose for $y_{1}\in\mathbb{R}$, $y_{2}\in\mathbb{R}$, and $y_{3}\in\mathbb{R}$,
$c_{1}-\gamma\Delta_{2}\geq y_{1},$ $\qquad c_{2}-c_{3}\geq y_{2},$
$c_{3}-\gamma\Delta_{2}-\frac{{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}}{2}$ $+\frac{y_{2}-\sqrt{y^{2}_{2}+4c^{2}_{3}f^{2}_{\max}}}{2}\geq y_{3}..$ (67)
Let us first simplify $(P1)$ as below:
$(P1)$ $=\min\limits_{\begin{subarray}{c}\theta^{2}_{1}\in[0,1-\theta^{2}_{3}-\theta^{2}_{4}]\\ \theta_{1}\in\mathbb{R}\end{subarray}}\left\{c_{1}\theta^{2}_{1}+\gamma\Delta_{2}(1-\theta^{2}_{1}-\theta^{2}_{3}-\theta^{2}_{4})-c_{3}\bar{f}\theta_{1}\theta_{4}\right.$
$\qquad\qquad\qquad\qquad\qquad\left.-\left(\gamma\left|{\theta_{1}}\right|+c_{2}\left|{\theta_{3}}\right|+2c_{3}\left|{\theta_{4}}\right|f_{\max}\right)\sqrt{1-\theta^{2}_{1}-\theta^{2}_{3}-\theta^{2}_{4}}\right\}$
$\geq\min\limits_{\theta_{1}\in\left[0,\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}\right]}\left\{c_{1}\theta^{2}_{1}+\gamma\Delta_{2}(1-\theta^{2}_{1}-\theta^{2}_{3}-\theta^{2}_{4})-c_{3}{f}_{\max}\left|{\theta_{1}}\right|\left|{\theta_{4}}\right|\right.$
$\qquad\qquad\qquad\qquad\qquad\left.-\left(\gamma\left|{\theta_{1}}\right|+c_{2}\left|{\theta_{3}}\right|+2c_{3}\left|{\theta_{4}}\right|f_{\max}\right)\sqrt{1-\theta^{2}_{1}-\theta^{2}_{3}-\theta^{2}_{4}}\right\}$
$\geq\gamma\Delta_{2}(1-\theta^{2}_{3}-\theta^{2}_{4})-c_{3}{f}_{\max}\left|{\theta_{4}}\right|\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}$
$\qquad\qquad+\min\limits_{\theta_{1}\in\left[0,\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}\right]}\bigg{\{}(c_{1}-\gamma\Delta_{2})\theta^{2}_{1}$
$\qquad\qquad\qquad\qquad\qquad\quad\quad-\left(\gamma\left|{\theta_{1}}\right|+c_{2}\left|{\theta_{3}}\right|+2c_{3}\left|{\theta_{4}}\right|f_{\max}\right)\sqrt{1-\theta^{2}_{1}-\theta^{2}_{3}-\theta^{2}_{4}}\bigg{\}}$
$\geq\gamma\Delta_{2}(1-\theta^{2}_{3}-\theta^{2}_{4})-c_{3}{f}_{\max}\left|{\theta_{4}}\right|\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}$
$\qquad\qquad\qquad-\left(c_{2}\left|{\theta_{3}}\right|+2c_{3}\left|{\theta_{4}}\right|f_{\max}\right)\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}$
$\qquad\qquad\qquad+\min\limits_{\theta_{1}\in\left[0,\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}\right]}\left\{(c_{1}-\gamma\Delta_{2})\theta^{2}_{1}-\gamma\sqrt{\theta^{2}_{1}\left(1-\theta^{2}_{1}-\theta^{2}_{3}\right)-\theta^{4}_{4}}\right\}$
$\geq\gamma\Delta_{2}(1-\theta^{2}_{3}-\theta^{2}_{4})-c_{3}{f}_{\max}\left|{\theta_{4}}\right|\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}$
$\qquad\qquad\qquad-\left(c_{2}\left|{\theta_{3}}\right|+2c_{3}\left|{\theta_{4}}\right|f_{\max}\right)\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}$
$\qquad\qquad\qquad+\min\limits_{\theta_{1}\in\left[0,\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}\right]}\left\{y_{1}\theta^{2}_{1}-\gamma\sqrt{\theta^{2}_{1}\left(1-\theta^{2}_{1}-\theta^{2}_{3}\right)-\theta^{4}_{4}}\right\}$
=$\overset{(a)}{=}\gamma\Delta_{2}(1-\theta^{2}_{3}-\theta^{2}_{4})-c_{3}{f}_{\max}\left|{\theta_{4}}\right|\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}$
$\qquad\qquad\qquad-\left(c_{2}\left|{\theta_{3}}\right|+2c_{3}\left|{\theta_{4}}\right|f_{\max}\right)\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}$
$\qquad\qquad\qquad+\frac{1-\theta^{2}_{3}-\theta^{2}_{4}}{2}\left(y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}\right),$
where $(a)$ follows from Lemma G.3. Now, substituting $(P1)$ in (66), we get
$\min\limits_{\begin{subarray}{c}\|\Theta\|^{2}_{2}=1,\\ \vec{\theta}_{2}\in E\end{subarray}}~{}-\Theta^{T}A\Theta$
$\geq$ $\min\limits_{\begin{subarray}{c}\theta^{2}_{3}+\theta^{2}_{4}\in[0,1]\end{subarray}}~{}c_{2}\theta^{2}_{3}+c_{3}\theta^{2}_{4}-2c_{3}f_{\max}\left|{\theta_{3}}\right|\left|{\theta_{4}}\right|+\gamma\Delta_{2}(1-\theta^{2}_{3}-\theta^{2}_{4})$
$\qquad\qquad\qquad-c_{3}{f}_{\max}\left|{\theta_{4}}\right|\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}-\left(c_{2}\left|{\theta_{3}}\right|+2c_{3}\left|{\theta_{4}}\right|f_{\max}\right)\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}$
$\qquad\qquad\qquad+\frac{1-\theta^{2}_{3}-\theta^{2}_{4}}{2}\left(y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}\right)$
$=$ $\min\limits_{x^{2}\in[0,1]}~{}\min\limits_{\begin{subarray}{c}\theta^{2}_{3}+\theta^{2}_{4}=x^{2}\end{subarray}}~{}c_{2}\theta^{2}_{3}+c_{3}\theta^{2}_{4}-2c_{3}f_{\max}\left|{\theta_{3}}\right|\left|{\theta_{4}}\right|+\gamma\Delta_{2}(1-\theta^{2}_{3}-\theta^{2}_{4})$
$\qquad\qquad\qquad-\left(c_{2}\left|{\theta_{3}}\right|+3c_{3}\left|{\theta_{4}}\right|f_{\max}\right)\sqrt{1-\theta^{2}_{3}-\theta^{2}_{4}}$
$\qquad\qquad\qquad+\frac{1-\theta^{2}_{3}-\theta^{2}_{4}}{2}\left(y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}\right)$
$=$ $\gamma\Delta_{2}+\frac{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}{2}+~{}\min\limits_{x^{2}\in[0,1]}~{}\left(c_{3}-\gamma\Delta_{2}-\frac{{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}}{2}\right)x^{2}$
$\qquad\qquad\qquad+~{}\min\limits_{\begin{subarray}{c}\theta^{2}_{3}\in[0,x^{2}]\end{subarray}}~{}~{}~{}\bigg{\{}(c_{2}-c_{3})\theta^{2}_{3}$
$\qquad\qquad\qquad\qquad-2c_{3}f_{\max}\sqrt{\theta^{2}_{3}x^{2}-\theta^{4}_{3}}-\left(c_{2}\left|{\theta_{3}}\right|+3c_{3}f_{\max}\sqrt{x^{2}-\theta^{2}_{3}}\right)\sqrt{1-x^{2}}\bigg{\}}$
$\geq$ $\gamma\Delta_{2}+\frac{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}{2}+~{}\min\limits_{x\in[0,1]}~{}\left\{c_{3}-\gamma\Delta_{2}-\frac{{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}}{2}\right\}x^{2}$
$\qquad\qquad\qquad-\left(c_{2}+3c_{3}f_{\max}\right)\sqrt{x^{2}-x^{4}}$
$\qquad\qquad\qquad~{}+~{}\min\limits_{\begin{subarray}{c}\theta_{3}\in[0,x]\end{subarray}}~{}~{}~{}\left\{(c_{2}-c_{3})\theta^{2}_{3}-2c_{3}f_{\max}\sqrt{\theta^{2}_{3}x^{2}-\theta^{4}_{3}}\right\}$
$\geq$ $\gamma\Delta_{2}+\frac{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}{2}+~{}\min\limits_{x\in[0,1]}~{}\left\{c_{3}-\gamma\Delta_{2}-\frac{{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}}{2}\right\}x^{2}$
$\qquad\qquad\qquad-\left(c_{2}+3c_{3}f_{\max}\right)\sqrt{x^{2}-x^{4}}+~{}\min\limits_{\begin{subarray}{c}\theta_{3}\in[0,x]\end{subarray}}~{}~{}\left\{y_{2}\theta^{2}_{3}-2c_{3}f_{\max}\sqrt{\theta^{2}_{3}x^{2}-\theta^{4}_{3}}\right\}$
=$\overset{(a)}{=}$ $\gamma\Delta_{2}+\frac{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}{2}-\left(c_{2}+3c_{3}f_{\max}\right)\sqrt{x^{2}-x^{4}}$
$\qquad\qquad\qquad+~{}\min\limits_{x^{2}\in[0,1]}~{}\left\{c_{3}-\gamma\Delta_{2}-\frac{{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}}{2}+\frac{y_{2}-\sqrt{y^{2}_{2}+4c^{2}_{3}f^{2}_{\max}}}{2}\right\}x^{2}$
$\geq$ $\gamma\Delta_{2}+\frac{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}{2}+~{}\min\limits_{x^{2}\in[0,1]}~{}y_{3}x^{2}-\left(c_{2}+3c_{3}f_{\max}\right)\sqrt{x^{2}-x^{4}}$
=$\overset{(b)}{=}$ $\gamma\Delta_{2}+\frac{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}{2}+\frac{y_{3}-\sqrt{y^{2}_{3}+(c_{2}+3c_{3}f_{\max})^{2}}}{2},$
where, $(a)$ and $(b)$ follow from Lemma G.3.
Conditions on $c_{1}$, $c_{2}$, and $c_{3}$ to ensure a strictly positive lower bound. For $d_{1}\in[-1,0]$, $d_{2}<0$, and $d_{3}\in[-1,0]$, let
$\frac{y_{1}-\sqrt{y^{2}_{1}+c^{2}_{2}}}{2}=d_{1}\Delta_{2},$
$\frac{y_{2}-\sqrt{y^{2}_{2}+4c^{2}_{3}f^{2}_{\max}}}{2}=d_{2}\Delta_{2},$
and
$\frac{y_{3}-\sqrt{y^{2}_{3}+(c_{2}+3c_{3}f_{\max})^{2}}}{2}=d_{3}\Delta_{2}.$
Then,
$y_{1}=-\frac{c^{2}_{2}}{4d_{1}\Delta_{2}}+d_{1}\Delta_{2},\quad y_{2}=-\frac{c^{2}_{3}f^{2}_{\max}}{d_{2}\Delta_{2}}+d_{2}\Delta_{2},\quad\text{ and }\quad y_{3}=-\frac{(c_{2}+3c_{3}f_{\max})^{2}}{4d_{3}\Delta_{2}}+d_{3}\Delta_{2}.$
On choosing $\gamma=1$, the lower bound becomes
$(1+d_{1}+d_{3})\Delta_{2},$
and for $c>0$, setting
$d_{2}=-c(1+d_{1}+d_{3}),$
the conditions in (67) reduce to
$c_{1}\geq-\frac{1}{4d_{1}\Delta_{2}}+\Delta_{2}(1+d_{1}),$
$c_{2}-c_{3}-\frac{c^{2}_{3}f^{2}_{\max}}{c(1+d_{1}+d_{3})\Delta_{2}}+c(1+d_{1}+d_{3})\Delta_{2}\geq 0$
$c_{3}+\frac{(c_{2}+3c_{3}f_{\max})^{2}}{4d_{3}\Delta_{2}}-(1+c)(1+d_{1}+d_{3})\Delta_{2}\geq 0,$
which for $f_{\max}=1$, reduce to
$c_{1}\geq-\frac{1}{4d_{1}\Delta_{2}}+\Delta_{2}(1+d_{1}),$ (68)
$c_{2}-c_{3}-\frac{c^{2}_{3}}{c(1+d_{1}+d_{3})\Delta_{2}}+c(1+d_{1}+d_{3})\Delta_{2}\geq 0$ (69)
$c_{3}+\frac{(c_{2}+3c_{3})^{2}}{4d_{3}\Delta_{2}}-(1+c)(1+d_{1}+d_{3})\Delta_{2}\geq 0.$ (70)
Specific choices for the parameters. The conditions (69) and (70) imply
$d_{3}\leq-\frac{9(1+c)(1+d_{1})}{1+9(1+c)},$
$\max\left\{(1+c)(1+d_{1}+d_{3})\Delta_{2},-\frac{2\Delta_{2}}{9}\left(d_{3}+\sqrt{d^{2}_{3}+9d_{3}(1+c)(1+d_{1}+d_{3})}\right)\right\}\leq c_{3},$
$c_{3}\leq\frac{2\Delta_{2}}{9}\left(-d_{3}+\sqrt{d^{2}_{3}+9d_{3}(1+c)(1+d_{1}+d_{3})}\right),$
$c_{3}+\frac{c^{2}_{3}}{c(1+d_{1}+d_{3})\Delta_{2}}-c(1+d_{1}+d_{3})\Delta_{2}\leq c_{2},$
and
$c_{2}\leq-3c_{3}+2\sqrt{d_{3}\Delta_{2}(-c_{3}+(1+c)(1+d_{1}+d_{3})\Delta_{2})}.$
Choose $c$, $d_{1}$, $d_{3}$, and $c_{1}$ as below:
$c=-\frac{7}{24},\quad d_{1}=-\frac{1}{2},\quad d_{3}=-\frac{75}{166},\quad c_{1}\geq\frac{1}{2\Delta_{2}}+\frac{\Delta_{2}}{2}.$
Further, let $c_{2}$ and $c_{3}$ satisfy
$\frac{5}{249}(5-2\sqrt{2})\Delta_{2}\leq c_{3}\leq\frac{5}{249}(5+2\sqrt{2})\Delta_{2},$
$c_{3}-\frac{498c^{2}_{3}}{7\Delta_{2}}+\frac{7\Delta_{2}}{498}\leq c_{2}\leq-3c_{3}+\frac{5}{83}\sqrt{498c_{3}\Delta_{2}-17\Delta^{2}_{2}}.$
We remark that the range of $c_{2}$ above is non-trivial for the feasible choices of $c_{3}$, and with these choices,
$\min\limits_{\begin{subarray}{c}\|\Theta\|^{2}_{2}=1,\\ \vec{\theta}_{2}\in E\end{subarray}}~{}-\Theta^{T}A\Theta\geq(1+d_{1}+d_{3})\Delta_{2}\geq\frac{4\Delta_{2}}{83}>\frac{\Delta_{2}}{20}>0.$
This proves the required contraction property of the update matrix $A$.
D.2 Details of Step 2: MSE Bound
In this section, we add details to the discussion in Section 3.4.2. In this appendix, we verify that the conditions in Theorem A.1 are met by the iterates and update matrices of our estimator, which will enable us to use Theorem A.1 to arrive at the desired bounds.
Consider the sequence of iterates $\Theta_{k}:=[\bar{f}_{k}~{}\theta^{T}_{k}~{}\widetilde{V}_{k}~{}\kappa_{k}]$ generated by the proposed algorithm in Section 3.1 with $\Theta_{0}={\vec{\bf 0}}$ and $\theta_{k}\in E$. Recall from Section 3.4 that these iterates can be rewritten as a linear stochastic-approximation update given below:
$\Theta_{k+1}=\Theta_{k}+\alpha_{k}\left(A(Y_{k})\Theta_{k}+b(Y_{k})\right).$
Moreover, in our setting, recall that $\mathbb{S}=E$, and $\Theta_{k}\in E$, $\Theta^{*}\in E$. Lemma 3.3 implies that the condition A.1 on update matrix A in Theorem A.1 is also satisfied in our setting with $\gamma_{2}=\frac{\Delta_{2}}{20}$. It only remains to verify the bound H  on update matrices. In what follows, we will show that H  in Theorem A.1 corresponds to $\eta$ in our setting. Theorem 3.1 then corresponds to Theorem A.1 with these choices.
Bounds on ${A}(Y_{t})$ and ${b}(Y_{t})$. Define $\tilde{A}(Y_{k})$ and $\tilde{b}(Y_{k})$ so that $A(Y_{k})=\Pi\tilde{A}(Y_{k})$ and $b(Y_{k})=\Pi\tilde{b}(Y_{k})$, where
$\Pi:=\begin{bmatrix}1&\vec{\bf 0}^{T}&0&0\\ \vec{\bf 0}&\Pi_{2,E}&\vec{\bf 0}&\vec{\bf 0}\\ 0&\vec{\bf 0}^{T}&1&0\\ 0&\vec{\bf 0}^{T}&0&1\end{bmatrix}.$
Similarly, define $\tilde{A}$ and $\tilde{b}$ so that $A=\Pi\tilde{A}$ and $b=\Pi\tilde{b}$, i.e.,
$\tilde{A}=\begin{bmatrix}-c_{1}&\vec{\bf 0}^{T}&0&0\\ -\Phi^{T}D_{\pi}\vec{\bf 1}&{\Phi^{T}D_{\pi}(P-I)\Phi}&\vec{\bf 0}&\vec{\bf 0}\\ 0&c_{2}\vec{\bf 1}^{T}D_{\pi}\Phi&-c_{2}&0\\ c_{3}\bar{f}&2c_{3}\mathcal{F}^{T}D_{\pi}\Phi&-2c_{3}\bar{f}&-c_{3}\end{bmatrix},\quad\text{and}\quad\tilde{b}=\begin{bmatrix}c_{1}\bar{f}\\ \Phi^{T}D_{\pi}\mathcal{F}\\ 0\\ -{c_{3}}\mathbb{E}_{\pi\!}\!\left[f^{2}(X)\right]\end{bmatrix},$
For a matrix $M$, let $\|M\|_{2}$ and $\|M\|_{F}$ denote its induced $2$-norm and Frobenius norm, respectively. Then, recall that $\|M\|_{2}\leq\|M\|_{F}$.
$\|{A}(Y_{t})\|_{2}$ $=\|\Pi\tilde{A}(Y_{t})\|_{2}\leq\|\Pi\|_{2}\|\tilde{A(Y_{t})}\|_{2},$
where the above inequality follows from sub-multiplicativity of the induced $2$-norm. Now, since $\Pi$ is an orthogonal projection matrix, $\|\Pi\|_{2}=1$. Thus, we have
$\|A(Y_{t})\|_{2}$ $\leq\|\tilde{A}(Y_{t})\|_{2}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\|\tilde{A}(Y_{t})\|_{F}$
$\leq\sqrt{c^{2}_{1}+1+\|\phi_{k}(\phi^{T}_{k+1}-\phi^{T}_{k})\|^{2}_{F}+2c^{2}_{2}+c^{2}_{3}f^{2}_{\max}+8c^{2}_{3}f^{2}_{\max}+c^{2}_{3}}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\sqrt{c^{2}_{1}+1+(\|\phi_{k}\phi^{T}_{k+1}\|_{F}+\|\phi_{k}\phi^{T}_{k}\|_{F})^{2}+2c^{2}_{2}+10c^{2}_{3}}$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sqrt{c^{2}_{1}+1+(\|\phi_{k}\|_{2}\|\phi_{k+1}\|_{2}+\|\phi_{k}\|_{2}\|\phi_{k}\|_{2})^{2}+2c^{2}_{2}+10c^{2}_{3}}$
$\stackrel{{\scriptstyle(d)}}{{\leq}}\sqrt{c^{2}_{1}+5+2c^{2}_{2}+10c^{2}_{3}},$
where $(a)$ follows since for a matrix $M$, $\|M\|_{2}\leq\|M\|_{F}$, $(b)$ uses triangle inequality for $\|\cdot\|_{F}$ and that $f_{\max}=1$, and $(d)$ uses Assumption 1. To conclude $(c)$, let $x_{1}\in\mathbb{R}^{d}$ and $x_{2}\in\mathbb{R}^{d}$. Observe the following equalities for $\|x_{1}x^{T}_{2}\|_{F}$.
$\|x_{1}x^{T}_{2}\|_{F}=\sqrt{\operatorname{Trace}(x_{1}x^{T}_{2}x_{2}x^{T}_{1})}=\sqrt{\operatorname{Trace}(x_{1}x^{T}_{1}\|x_{2}\|^{2}_{2})}=\|x_{2}\|_{2}(x^{T}_{1}x_{1})=\|x_{1}\|_{2}\|x_{2}\|_{2}.$
Next, to bound ${b}(Y_{t})$,
$\|{b}(Y_{t})\|_{2}=\|\Pi~{}\tilde{b}(Y_{t})\|_{2}\leq\|\tilde{b}(Y_{t})\|_{2}=\sqrt{c^{2}_{1}+1+{c^{2}_{3}}},$
where we again used that $f_{\max}=1$. Thus,
$\eta:=\sqrt{c^{2}_{1}+5+2c^{2}_{2}+10c^{2}_{3}}$
is the required upper bound H .
Appendix E Approximation Error due to LFA: Proof of Proposition 3.2
Approximation error in $V$ estimation. Since any element from $S_{f}$ is a valid estimator for $V$ function, and since from Theorem 3.1, the proposed algorithm’s estimate for $V$ converges to $\Phi\theta^{*}$, where $\theta^{*}$ is the unique vector in the subspace $E$ that satisfies (3.2), the approximation error incurred in estimating $V$ using the proposed TD-like algorithm from Section 3.1 is the minimum weighted distance between the estimate using limit of the algorithm, i.e. $\Phi\theta^{*}$, and the set $S_{f}$, given by
${\mathcal{E}}_{D_{\pi},S_{f}}\left[\Phi\theta^{*}\right]:=\inf\limits_{V\in S_{f}}\|\Phi\theta^{*}-V\|^{2}_{D_{\pi}}.$
Following arguments similar to those in Tsitsiklis and Van Roy [1999], one can relate this error to the minimum possible approximation error (using any algorithm) incurred due to the chosen architecture,
$\mathcal{E}:=\inf\limits_{\theta\in\mathbb{R}^{d}}\inf\limits_{V\in S_{f}}\|\Phi\theta-V\|_{D_{\pi}}.$
In particular, it can be shown that for a constant $\lambda\in(0,1)$,
${\mathcal{E}}_{D_{\pi},S_{f}}\left[\Phi\theta^{*}\right]\leq\frac{\mathcal{E}}{\sqrt{1-\lambda^{2}}}.$
Note that the bound above is a blow-up of the minimum error possible due to the chosen architecture for approximation.
Approximation error in variance estimation. Using this, we now bound the approximation error for the variance estimation using the proposed algorithm. Recall from (4) that for any $V\in S_{f}$,
$\kappa(f)=2\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})V(X)\right]-\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})^{2}\right].$
Moreover, from (19), one can see that
$\kappa^{*}=2\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})[\Phi\theta^{*}](X)\right]-\mathbb{E}_{\pi\!}\!\left[(f(X)-\bar{f})^{2}\right].$
Thus, for any $c\in\mathbb{R}$,
$\left|{\kappa(f)-\kappa^{*}}\right|$ $=2\left|{\mathbb{E}_{\pi\!}\!\left[\left(f(X)\!-\!\bar{f}\right)\left(V^{*}(X)+c\!-\!\left[\Phi\theta^{*}\right](X)\right)\right]}\right|$
$=2\left|{\left<\mathcal{F}-\bar{f}{\bf e},V^{*}+c{\bf e}-\Phi\theta^{*}\right>_{D_{\pi}}}\right|$
$=2\left|{\left<{D^{\frac{1}{2}}_{\pi}}\left(\mathcal{F}-\bar{f}{\bf e}\right),{D^{\frac{1}{2}}_{\pi}}\left(V^{*}+c{\bf e}-\Phi\theta^{*}\right)\right>}\right|$
$\leq 2\|D^{\frac{1}{2}}_{\pi}\left(\mathcal{F}-\bar{f}{\bf e}\right)\|_{2}~{}\|D^{\frac{1}{2}}_{\pi}\left(V^{*}+c{\bf e}-\Phi\theta^{*}\right)\|_{2}$
$\leq 4f_{\max}~{}\|V^{*}+c{\bf e}-\Phi\theta^{*}\|_{D_{\pi}}$
$=4~{}\|V^{*}+c{\bf e}-\Phi\theta^{*}\|_{D_{\pi}}.$
Since the above inequality is true for all $c\in\mathbb{R}$, the approximation error for $\kappa(f)$ is bounded as below:
$(\kappa(f)-\kappa^{*})^{2}$ $\leq 16~{}\inf\limits_{c\in\mathbb{R}}~{}\|\Phi\theta^{*}-V^{*}-c{\bf e}\|^{2}_{D_{\pi}}$
$=16~{}\inf\limits_{V\in S_{f}}~{}\|\Phi\theta^{*}-V\|^{2}_{D_{\pi}}$
$=16~{}\left({\mathcal{E}}_{D_{\pi},S_{f}}\left[\Phi\theta^{*}\right]\right)^{2}$
$\leq\frac{16~{}{\mathcal{E}}^{2}}{1-\lambda^{2}}.\qed$
Appendix F Policy Evaluation with Linear Function Approximation in RL
When the state and action spaces of the underlying MDP are large, estimating the $Q_{\mu}$ function for each state-action pair (as in Algorithm 3) requires a lot of memory and may be intractable. To address this, we consider estimating an approximation of $\kappa_{\mu}$ by extending the setup of Section 3 to RL, by approximating $Q_{\mu}$ with its projection on to a linear subspace spanned by a given fixed set of $d$ vectors. We discuss this framework in the current section.
F.1 Setup
Let us again denote the given set of $d$ vectors by $\{\tilde{\phi}_{1},\dots,\tilde{\phi}_{d}\}$, where $\tilde{\phi}_{i}\in\mathbb{R}^{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}}\right|}$ for $i\in[d]$. In particular, for $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$, and $\theta\in\mathbb{R}^{d}$, we consider a linear function approximation
$Q_{\theta}(s,a)=\phi(s,a)^{T}\theta$
of $Q_{\mu}(s,a)$, where
$\phi^{T}(s,a):=[\tilde{\phi}_{1}(s,a)\dots\tilde{\phi}_{d}(s,a)]$
is the feature vector for state-action pair $(s,a)$ and $\phi(s,a)\in\mathbb{R}^{d}$. With this notation, let $\Phi$ be a $\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}}\right|\times d$ matrix with $\tilde{\phi}_{i}$ being the $i^{\text{th}}$ column. Let $W_{\Phi}=\{\Phi\theta:\theta\in\mathbb{R}^{d}\}$ denote the column space of $\Phi$. Then, $Q_{\theta}=\Phi\theta$, where $Q_{\theta}\in\mathbb{R}^{\left|{{\mathcal{S}}}\right|\left|{{\mathcal{A}}}\right|\times 1}$ is an approximation for $Q_{\mu}$ using $\theta$. We make the following assumption on the feature matrix $\Phi$, which can be achieved by feature normalization, and is standard in literature; see Tsitsiklis and Van Roy [1999], Bertsekas and Tsitsiklis [1996].
Assumption 3.
The matrix $\Phi$ is full rank, i.e., the set of feature vectors $\{\tilde{\phi}_{1},\dots,\tilde{\phi}_{d}\}$ are linearly independent. Additionally, $\|\phi(s,a)\|_{2}\leq 1$, for each $(s,a)\in{\mathcal{S}}\times{\mathcal{A}}$.
Define subspace $S_{\Phi,e}$ of $\mathbb{R}^{d}$ as $S_{\Phi,e}:=\text{span}\left(\left\{\theta|\Phi\theta={\bf e}\right\}\right)$. It equals $\left\{c\theta_{e}|c\in\mathbb{R}\right\}$ if ${\bf e}\in W_{\Phi},$ and $\theta_{e}$ is the unique vector (if it exists) such that $\Phi\theta_{e}={\bf e}$. Otherwise, $S_{\Phi,e}=\{0\}$. Let $E$ be the subspace of $\mathbb{R}^{d}$ that is orthogonal complement (in $\ell_{2}$-norm) of $S_{\Phi,e}$, i.e., $E=\{\theta\in\mathbb{R}^{d}:~{}\theta^{T}\theta_{e}=0\}$ if ${\bf e}\in W_{\Phi}$. It equals $\mathbb{R}^{d}$, otherwise. Additionally, let $\Pi_{2,E}$ denote the orthogonal projection of vectors in $\mathbb{R}^{d}$ (in $2$-norm) on the subspace $E$.
Observe that for $\theta\in E$, $\Phi\theta$ is a non-constant vector. This follows since there does not exist $\theta\in E$ such that $\Phi\theta={\bf e}$. In particular, if ${\bf e}\in W_{\Phi}$, then the unique vector $\theta_{e}$ such that $\Phi\theta_{e}={\bf e}$ does not belong to $E$. Similarly, if ${\bf e}\notin W_{\Phi}$, then there does not exist a vector $\theta\in\mathbb{R}^{d}$, hence in $E$, such that $\Phi\theta={\bf e}$.
In this setup, let $\theta^{*}_{\mu}$ denote the unique vector in $E$ that is a solution for
$\Phi\theta^{*}_{\mu}=\Pi_{D_{\mu},W_{\Phi}}T\Phi\theta^{*}_{\mu},$
where recall that $D_{\mu}$ is the diagonal matrix of stationary distribution of the underlying Markov chain (defined around (26)), $\Pi_{D_{\mu},W_{\Phi}}$ is the projection matrix onto the subspace $W_{\Phi}$ in $D_{\mu}$ norm, and $T$ is the operator for which $Q_{\mu}$ is a fixed point, i.e., it acts on vectors $Q\in\mathbb{R}^{|{\mathcal{S}}||{\mathcal{A}}|}$, and satisfies
$TQ=\mathcal{R}-J_{\mu}{\bf e}+P_{2}Q,$
where $P_{2}$ is the transition matrix for the underlying Markov chain, defined in (26).
Algorithm 4 in the next section, estimates $\theta^{*}_{\mu}$ using stochastic approximation for (F.1), and estimates
$\kappa^{*}_{\mu}=\mathbb{E}_{d_{\mu}\!}\!\left[2r(S,A)[\Phi\theta^{*}_{\mu}](S,A)-2r(S,A)\widetilde{Q}_{\mu}-r^{2}(S,A)+r(S,A)J_{\mu}\right],$
where $\widetilde{Q}_{\mu}:=\mathbb{E}_{d_{\mu}\!}\!\left[[\Phi\theta^{*}_{\mu}](S,A)\right]$ as a proxy for $\kappa_{\mu}$. Observe that $\kappa^{*}_{\mu}$ differs from $\kappa_{\mu}$ in that $Q_{\mu}$ and $\bar{Q}_{\mu}$ in formulation (31) are replaced by the estimate in the subspace spanned by $\Phi$ and $\widetilde{Q}$, respectively.
F.2 Algorithm
We now adapt the Algorithm from Section 3.1 to RL for estimating the asymptotic variance of a given stationary policy $\mu$ with linear function approximation. Here, we estimate a vector $\theta\in\mathbb{R}^{d}$ at each step; call this estimate at step $k$ as $\theta_{k}$. The corresponding estimate for $Q_{\mu}$ is then $\Phi\theta_{k}$.
Input: Time horizon $n>0$, constants $c_{1}>0$, $c_{2}>0$, $c_{3}>0$ and step-size sequence $\left\{\alpha_{k}\right\}$.
Initialization: $\bar{r}_{0}=0$, $\theta_{0}=\vec{\bf 0}$, $\widetilde{Q}_{0}=0$, and $\kappa_{0}=0$.
for $k\leftarrow 1$ to $n$ do
      Observe ($S_{k},A_{k},r(S_{k},X_{k}),S_{k+1},A_{k+1}$).
      // Average reward $J_{\mu}$ estimation
$\bar{r}_{k+1}=\bar{r}_{k}+c_{1}\alpha_{k}(r(S_{k},A_{k})-\bar{r}_{k}).$
      // $\theta^{*}_{\mu}$ estimation
$\delta_{k}:=r(S_{k},A_{k})-\bar{r}_{k}+(\phi(S_{k+1},A_{k+1})-\phi(S_{k},A_{k}))^{T}\theta_{k},$
$\theta_{k+1}=\theta_{k}+\alpha_{k}\Pi_{2,E}\phi(S_{k},A_{k})\delta_{k}.$
       // $\widetilde{Q}_{\mu}$ estimation
$\widetilde{Q}_{k+1}=(1-c_{2}\alpha_{k})\widetilde{Q}_{k}+c_{2}\alpha_{k}\phi^{T}(S_{k},A_{k})\theta_{k}.$
      // Asymptotic Variance $\kappa_{\mu}$ estimation
$\kappa_{k+1}=(1-c_{3}\alpha_{k})\kappa_{k}$ $+c_{3}\alpha_{k}\big{(}2r(S_{k},A_{k})\phi^{T}(S_{k},A_{k})\theta_{k}$
$\qquad\qquad\qquad-2r(S_{k},A_{k})\widetilde{Q}_{k}-r^{2}(S_{k},A_{k})+r(S_{k},A_{k})\bar{r}_{k}\big{)}.$
end for
Output : Estimates $\kappa_{n}$, $\widetilde{Q}_{n}$, $\theta_{n}$, $\bar{r}_{n}$
Algorithm 4 Estimating $\kappa_{\mu}$: Linear Function Approximation Setting
Note that Algorithm 3 for the tabular setting is a special case with $d=\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|$ and $\Phi=I$, the identity matrix in $\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|$ dimensions. We conclude this section with a finite-sample bound on the estimation error of the proposed algorithm with linear function approximation.
F.3 Convergence Rates
Let $\Theta^{T}_{k}:=[r_{k}~{}\theta^{T}_{k}~{}\widetilde{Q}_{k}~{}\kappa_{k}]$ denote the vector of estimates of Algortithm 4 at time $k$. Recall the Markov chain $\mathcal{M}_{3}$ introduced in Section 5.5. Define
$\tilde{\Delta}_{2}:=\min\limits_{\|\theta\|_{2}=1,\theta\in E}~{}\theta^{T}\Phi^{T}D_{\mu}(I-P_{2})\Phi\theta.$
As earlier, this can be shown to be strictly positive since for $\theta\in E$, $\Phi\theta$ is a non-constant vector. Finally, let $\eta:=(c^{2}_{1}+5+2c^{2}_{2}+10c^{2}_{3})^{\frac{1}{2}}$, and let $\Theta^{*T}_{\mu}:=[J_{\mu}~{}\theta^{*T}_{\mu}~{}\widetilde{Q}_{\mu}~{}\kappa^{*}_{\mu}]$, where each of these terms is defined around (F.1). The following theorem bounds the mean-squared distance between $\Theta_{k}$, the vector of estimates of Algorthm 4 at time $k$, and its limit point $\Theta^{*}_{\mu}$, under constant as well as diminishing step sizes.
Theorem F.1.
Consider estimates ${\Theta}_{k}$ generated by Algorithm 4 with $c_{1}$, $c_{3}$, and $c_{2}$ satisfying:
$c_{1}\geq\frac{1}{2\tilde{\Delta}_{2}}+\frac{\tilde{\Delta}_{2}}{2},\qquad\frac{5}{249}(5-2\sqrt{2})\tilde{\Delta}_{2}\leq c_{3}\leq\frac{5}{249}(5+2\sqrt{2})\tilde{\Delta}_{2},$
and
$c_{3}-\frac{498c^{2}_{3}}{7\tilde{\Delta}_{2}}+\frac{7\tilde{\Delta}_{2}}{498}\leq c_{2}\leq-3c_{3}+\frac{5}{83}\sqrt{498c_{3}\tilde{\Delta}_{2}-17\tilde{\Delta}^{2}_{2}}.$
For a constant $B>1$, let $\xi_{1}:=3\left(1+\|\Theta^{*}_{\mu}\|_{2}\right)^{2}$, $\xi_{2}:=112B(1+\|\Theta^{*}_{\mu}\|_{2})^{2}$, where recall that $\eta$ and $\Theta^{*}_{\mu}$ were defined below (F.3).
  1. (a)
    Let $\alpha_{i}=\alpha$ for all $i$, such that
    $\alpha<\min\left\{\frac{40}{\tilde{\Delta}_{2}},\frac{1}{28B\left(1+\frac{20\eta^{2}}{\tilde{\Delta}_{2}}\right)}\right\}.$
    Then, for all $k\geq 1$,
    $\mathbb{E}\!\left[\|\Theta_{k}-\Theta^{*}_{\mu}\|^{2}_{2}\right]\leq\xi_{1}\left(1-\frac{\tilde{\Delta}_{2}\alpha}{40}\right)^{k}+\frac{20\xi_{2}\alpha\eta^{2}}{\tilde{\Delta}_{2}}+\xi_{2}\alpha.$
  2. (b)
    Let $\alpha_{i}=\frac{\alpha}{i+h}$ for all $i\geq 1$, with $\alpha$ and $h$ chosen so that
    $h\geq\max\left\{1+\frac{\alpha\tilde{\Delta}_{2}}{40},1+28B\left(\frac{20\eta^{2}\alpha}{\tilde{\Delta}_{2}}+\alpha+\frac{20}{\tilde{\Delta}_{2}}\right)\right\},\quad\text{and}\quad\alpha>\frac{40}{\tilde{\Delta}_{2}}.$
    Then, for all $k\geq 0$,
    $\mathbb{E}\!\left[\|\Theta_{k}-\Theta^{*}_{\mu}\|^{2}_{2}\right]$ $\leq\xi_{1}\left(\frac{h}{k+h}\right)^{\frac{\alpha\tilde{\Delta}_{2}}{40}}+\frac{5\xi_{2}e^{2}\eta^{2}(20+\tilde{\Delta}_{2})\alpha^{2}}{(k+h)(\alpha\tilde{\Delta}_{2}-40)}+\frac{\xi_{2}\alpha}{k+h}.$
Proof sketch. The bounds in this theorem follow immediately from those in Theorem 3.1 by observing that the setting considered here corresponds to that of Theorem 3.1 with $r(\cdot,\cdot)$, $d_{\mu}$, $\theta^{*}_{\mu}$, $\widetilde{Q}_{\mu}$, and $\kappa^{*}_{\mu}$ replacing $f(\cdot)$, $\pi$, $\theta^{*}$, $\widetilde{V}$, and $\kappa^{*}$, respectively. ∎
Remark F.1.
As in Corollary 5.1.1, Theorem A.1 $(b)$ gives a MSE bound of $O(\frac{1}{n})$ for policy evaluation problem in average reward RL setting with LFA, improving the result of Zhang et al. [2021] by a multiplicative factor of $\log n$.
Appendix G Auxiliary Technical Lemmas
Lemma G.1.
For any $\alpha>0$ and $h\geq 2$, for all $k\geq 1$, and $\alpha_{k}=\frac{\alpha}{k+h}$ or $\alpha_{k}=\alpha$, we have
$\alpha_{k-1}-\alpha_{k}\leq\frac{3\alpha^{2}_{k}}{\alpha}.$
Proof.
The result is trivially true for $\alpha_{k}$ being a constant ($\alpha_{k}=\alpha$). Hence, we only provide a proof for the setting of a diminishing step-size ($\alpha_{k}=\frac{\alpha}{k+h}$). To this end, consider the following:
$\alpha_{k-1}-\alpha_{k}$ $=\frac{\alpha}{k+h-1}-\frac{\alpha}{k+h}$
$=\frac{\alpha}{(k+h)(k+h-1)}$
$=\frac{\alpha^{2}_{k}}{\alpha\left(1-\frac{1}{k+h}\right)}$
$\leq\frac{\alpha^{2}_{k}}{\alpha}\left(1+\frac{2}{k+h}\right)$
$\leq\frac{3\alpha^{2}_{k}}{\alpha},$
where we used $\frac{1}{1-x}\leq 1+2x$ for $x\leq\frac{1}{2}$ in the first inequality above. ∎
Lemma G.2.
For any $\alpha>0$, and either $\alpha_{k}=\alpha$ for all $k\geq 0$, or for $h\geq 2$, $\alpha_{k}=\frac{\alpha}{k+h}$, for all $k\geq 0$, we have
$\alpha_{k}\alpha_{k-1}\leq 2\alpha^{2}_{k}.$
Proof.
This is trivially true for $\alpha_{k}=\alpha$. Below, we show for diminishing step sizes of the form $\alpha_{k}=\frac{\alpha}{k+h}$, for $h\geq 2$. Consider the following:
$\alpha_{k}\alpha_{k-1}$ $=\frac{\alpha^{2}_{k}}{1-\frac{1}{k+h}}\leq\frac{\alpha^{2}_{k}}{1-\frac{1}{h}}\leq 2\alpha^{2}_{k},$
proving the lemma. ∎
Lemma G.3.
For $b\geq 0$, and $c\geq 0$,
$\min\limits_{x\in[0,c]}\left\{ax-b\sqrt{cx-x^{2}}\right\}=\frac{c}{2}\left(a-\sqrt{a^{2}+b^{2}}\right).$
Proof.
Let $f(x):=ax-b\sqrt{cx-x^{2}}$. Let $f^{\prime}(x)$ denote the derivative of $f(x)$ and $f^{{}^{\prime\prime}}(x)$ denote the corresponding second derivative, both evaluated at $x$. Then,
$f^{\prime}(x)=a-\frac{b(c-2x)}{2\sqrt{cx-x^{2}}},$
and $f^{{}^{\prime\prime}}(x)>0$ for $b\geq 0$. Solving for x such that $f^{\prime}(x)=0$, we get $x_{1}$ defined below as the optimizer.
$x_{1}:=\frac{c}{2}\left(1-\frac{a}{\sqrt{a^{2}+b^{2}}}\right),$
and $f(x_{1})$ is the desired optimal value. ∎